### v2

# Xing Huang a), Feng-Yu Wang a)

Abstract
The log-Harnack inequality and Bismut formula are established for McKean-Vlasov SDEs with singularities in all (time, space, distribution) variables, where the drift satisfies a standard integrability condition in time-space, and may be discontinuous in the distance induced by any Dini function. The main results considerably improve the existing ones for the case where the drift is $L$-differentiable and Lipschitz continuous in distribution with respect to the quadratic Wasserstein distance.
AMS subject Classification:  60H10, 60B05.
Keywords: McKean-Vlasov SDEs, Log-Harnack inequality, Bismut formula, Dini function, Wasserstein distance.
1 Introduction
Let $\mathscr{P}$ be the set of all probability measures on $\mathbb{R}^{d}$ equipped with the weak topology, and let $W_{t}$ be an $m$-dimensional Brownian motion on a complete filtration probability space $(\Omega,\{\mathscr{F}_{t}\}_{t\geq 0},\mathscr{F},\mathbb{P})$. Consider the following McKean-Vlasov SDE on $\mathbb{R}^{d}$:
(1.1) $\text{\rm{d}}X_{t}=b_{t}(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}t+\sigma_{t}(X_{t})\text{\rm{d}}W_{t},\ \ t\in[0,T],$
where $T>0$ is a fixed time, $\mathscr{L}_{X_{t}}$ is the distribution of $X_{t}$, and
$b:[0,T]\times\mathbb{R}^{d}\times\mathscr{P}\rightarrow\mathbb{R}^{d},\ \ \sigma:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{m}$
are measurable. Because of its wide applications, this type SDE has been intensively investigated, see for instance [4, 5, 8, 9, 16, 17, 22] and the survey [11].
In this paper, we study the regularity of (1.1) for distributions in
$\mathscr{P}_{k}:=\big{\{}\mu\in\mathscr{P}:\|\mu\|_{k}:=\mu(|\cdot|^{k})^{\frac{1}{k}}<\infty\big{\}},\ \ k\in(1,\infty).$
Note that $\mathscr{P}_{k}$ is a Polish space under the Wasserstein distance
$\mathbb{W}_{k}(\mu,\nu)=\inf_{\pi\in\mathscr{C}(\mu,\nu)}\bigg{(}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{k}\pi(\text{\rm{d}}x,\text{\rm{d}}y)\bigg{)}^{\frac{1}{k}},$
where $\mathscr{C}(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$. The SDE (1.1) is called well-posed for distributions in $\mathscr{P}_{k}$, if for any initial value $X_{0}$ with $\mathscr{L}_{X_{0}}\in\mathscr{P}_{k}$ (respectively, any initial distribution $\gamma\in\mathscr{P}_{k}$), it has a unique solution (respectively, a unique weak solution) $X=(X_{t})_{t\in[0,T]}$ such that $\mathscr{L}_{X_{\cdot}}:=(\mathscr{L}_{X_{t}})_{t\in[0,T]}\in C([0,T];\mathscr{P}_{k}).$ In this case, for any $\gamma\in\mathscr{P}_{k}$, let $P_{t}^{*}\gamma=\mathscr{L}_{X_{t}^{\gamma}}$ for the solution $X_{t}^{\gamma}$ with $\mathscr{L}_{X_{0}^{\gamma}}=\gamma$. We study the regularity of the map
$\mathscr{P}_{k}\ni\gamma\mapsto P_{t}f(\gamma):=\mathbb{E}[f(X_{t}^{\gamma})]=\int_{\mathbb{R}^{d}}f\text{\rm{d}}\{P_{t}^{*}\gamma\}$
for $t\in(0,T]$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d}),$ where $\mathscr{B}_{b}(\mathbb{R}^{d})$ is the space of bounded measurable functions on $\mathbb{R}^{d}$.
As powerful tools characterizing the regularity in distribution for stochastic systems, the dimension-free Harnack inequality due to [24], the log-Harnack inequality introduced in [25], and the Bismut (also called Bismut-Elworthy-Li) formula developed from [6, 10], have been intensively investigated. See for instance the monograph [26] for an account of related study on SPDEs.
In recent years, the log-Harnack inequality and Bismut type formula have also been established for McKean-Vlasov SDEs with coefficients regular in the distribution variable. Below we present a brief summary.
Write $b_{t}(x,\mu)=b_{t}^{(0)}(x)+b_{t}^{(1)}(x,\mu)$. According to [28], if $b^{(0)}$ satisfies some integrability condition on $(t,x)$, and there exists a constant $K_{b}\geq 0$ such that
$|b_{t}^{(1)}(x,\mu)-b_{t}^{(1)}(y,\nu)|\leq K_{b}(|x-y|^{2}+\mathbb{W}_{2}(\mu,\nu)^{2}),\ \ (x,\mu),(y,\nu)\in\mathbb{R}^{d}\times\mathscr{P}_{2},t\in[0,T],$
then there exists a constant $c>0$ such that the log-Harnack inequality
$P_{t}\log f(\tilde{\gamma})\leq\log P_{t}f(\gamma)+\frac{c}{t}\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ t\in(0,T],\ f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d}),\ \gamma,\tilde{\gamma}\in\mathscr{P}_{2}$
holds, where $\mathscr{B}_{b}^{+}(\mathbb{R}^{d})$ is the space of positive elements in $\mathscr{B}_{b}(\mathbb{R}^{d})$. This inequality is equivalent to the entropy-cost inequality
${\rm Ent}(P_{t}^{*}\gamma|P_{t}^{*}\tilde{\gamma})\leq\frac{c}{t}\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ t\in(0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{2},$
where ${\rm Ent}$ is the relative entropy, i.e. for any $\mu,\nu\in\mathscr{P}$, ${\rm Ent}(\nu|\mu):=\infty$ if $\nu$ is not absolutely continuous with respect to $\mu$, while
${\rm Ent}(\nu|\mu):=\mu(\rho\log\rho)=\int_{\mathbb{R}^{d}}(\rho\log\rho)\text{\rm{d}}\mu,\ \ \text{if}\ \rho:=\frac{\text{\rm{d}}\nu}{\text{\rm{d}}\mu}\ \text{exists. }$.
See also [14, 20, 27] for log-Harnack inequalities with more regular $b^{(0)}$, and see [18] for the dimension-free Harnack inequality with power.
If furthermore $b_{t}^{(1)}(x,\mu)$ is $L$-differentiable in $\mu\in\mathscr{P}_{k}$ (see Definition 3.1 below), the following Bismut type formula has been established in [29] for the intrinsic derivative $D^{I}_{\phi}$ (see Definition 3.1 below):
$D_{\phi}^{I}P_{t}f(\mu)=\mathbb{E}\big{[}f(X_{t}^{\mu})M_{t}^{\mu,\phi}\big{]},\ \ t\in(0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\mu\in\mathscr{P}_{k},\phi\in L^{k}(\mathbb{R}^{d}\rightarrow\mathbb{R}^{d};\mu),$
where $M_{t}^{\mu,\phi}$ is an explicit martingale. See [3, 5, 12, 19] for earlier results with more regular $b^{(0)}$. See [2] for the case where $\mu=\delta_{x}$ is the Dirac measure at $x\in\mathbb{R}^{d}$, and see [21, 23] for a less explicit Bismut formula involving in the inverse of the Malliavin matrix of the solution.
We emphasize that existing results on log-Harnack inequality and Bismut formula for McKean-Vlasov SDEs only apply to the case with coefficients regular in the distribution variable, i.e. either $\mathbb{W}_{2}$-Lipschitz continuous or $L$-differentiable. The reason is that the Zvonkin transform technique [33] used in these references only kills singularities in the time-spatial variables $(t,x)$, but not the distribution variable.
On the other hand, a derivative estimate has been presented in [7] for the heat kernel when the drift is of type $b_{t}(x,\mu(V))$, where $V$ is a Hölder continuous function, and $\mu(V):=\int_{\mathbb{R}^{d}}V\text{\rm{d}}\mu$. In this case, the drift is only Lipschitz continuous in distribution with respect to
$\mathbb{W}_{\varepsilon}(\mu,\nu):=\sup\big{\{}\big{|}\mu(f)-\nu(f)\big{|}:|f(x)-f(y)|\leq|x-y|^{\varepsilon}\big{\}}$
rather than $\mathbb{W}_{1}$, and hence also has certain singularity in the distribution variable. This result encourages us to establish the log-Harnack inequality and Bismut formula for McKean-Vlasov SDEs with coefficients singular in all time-spatial-distribution variables.
Indeed, we will establish the log-Harnack inequality and Bismut formula for McKean-Vlasov SDEs with stronger singularity in the distribution variable: the drift is only Lipschitz continuous with respect to
$\mathbb{W}_{\alpha}(\mu,\nu):=\sup\big{\{}\big{|}\mu(f)-\nu(f)\big{|}:|f(x)-f(y)|\leq\alpha(|x-y|)\big{\}},$
where $\alpha$ is the square root of a Dini function, i.e. $\alpha$ is concave with $\alpha(0)=0,\alpha^{\prime}(s)>0$ for $s>0$, and $\int_{0}^{1}\frac{\alpha(s)^{2}}{s}\text{\rm{d}}s<\infty$ (Dini condition for $\alpha^{2}$). Thus, the continuity in the distribution variable is even weaker than the Dini continuity, so that the existing study in the literature is considerably improved.
To make this improvement, we develop a new technique by establishing a-priori derivative estimates and Bismut formula for the decoupled singular SDEs. The key point is to find a constant $c>0$ such that
(1.2) $\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{2}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \gamma,\tilde{\gamma}\in\mathscr{P}_{2},$
and for the drift of type $b_{t}(x,\mu(V))$ with $[V]_{\alpha}:=\sup_{x\neq y}\frac{|V(x)-V(y)|}{\alpha(|x-y|)}\leq 1$,
(1.3) $\|D^{I}P_{t}V(\mu)\|_{L^{\frac{k}{k-1}}(\mu)}\leq\frac{c\,\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},$
where $D^{I}$ is the intrinsic derivative (see Definition 3.1 below).
2 Log-Harnack Inequality
Since $\mathbb{W}_{2}$ is involved in the log-Harnack inequality, in this section we mainly consider (1.1) for distributions in $\mathscr{P}_{2}$, but the drift may be discontinuous in $\mathbb{W}_{k}$ for any $k>0$. We first state the concrete assumption and the main result on the log-Harnack inequality, then present a complete proof in a separate subsection.
2.1 Assumption and main result
We will allow $b_{t}(x,\cdot)$ to be merely Lipschitz continuous in the sum of $\mathbb{W}_{2}$ and the Wasserstein distance induced by the square root of a Dini function.
Let $\alpha$ be in the following class
$\mathscr{A}:=\bigg{\{}\alpha:[0,\infty)\rightarrow[0,\infty)\ \text{is\ increasing\ and\ concave,\ }\alpha(0)=0,\ \int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r\in(0,\infty)\bigg{\}},$
where $\int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r<\infty$ is the Dini condition for $\alpha^{2}$. For a (real or Banach valued) function $f$, let
$[f]_{\alpha}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{\alpha(|x-y|)}$
be its continuity modulus in $\alpha$. Define the Wasserstein distance induced by $\alpha$:
$\mathbb{W}_{\alpha}(\mu,\nu):=\sup_{[f]_{\alpha}\leq 1}|\mu(f)-\nu(f)|,\ \ \mu,\nu\in\mathscr{P}_{\alpha}:=\big{\{}\mu\in\mathscr{P}:\mu(\alpha(|\cdot|))<\infty\big{\}},$
where $f$ are real functions and $\mu(f)=\int_{\mathbb{R}^{d}}f\text{\rm{d}}\mu$.
By the concavity of $\alpha$, $\mathbb{W}_{\alpha}$ is a complete distance on $\mathscr{P}_{\alpha}$, $\mathscr{P}_{k}\subset\mathscr{P}_{\alpha}$ for $k\geq 1$, and
(2.1) $\alpha(s+t)\leq\alpha(s)+\alpha(t),\ \ \alpha(rt)\leq r\alpha(t),\ \ \\ \ s,t>0,r\geq 1.$
These inequalities follow from $\alpha(0)=0$ and the decreasing monotonicity of $\alpha^{\prime}$ such that
$\alpha^{\prime}(s+t)\leq\alpha^{\prime}(s),\ \ \frac{\text{\rm{d}}}{\text{\rm{d}}t}\alpha(rt)=r\alpha^{\prime}(rt)\leq r\alpha^{\prime}(t),\ \ s,t\geq 0,r\geq 1.$
Since $\alpha$ is increasing with $\int_{0}^{1}\frac{\alpha(s)^{2}}{s}\text{\rm{d}}s>0,$ we have $\alpha(1)>0$ so that the second estimate in (2.1) with $r=t^{-1}$ yields
(2.2) $\alpha(t)\geq\alpha(1)t>0,\ \ \ t\in(0,1].$
To measure the singularity in $(t,x)\in[0,T]\times\mathbb{R}^{d}$, we recall locally integrable functional spaces presented in [30]. For any $t>s\geq 0$ and $p,q\in(1,\infty]$, we write $f\in\tilde{L}_{p}^{q}([s,t])$ if $f:[s,t]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is measurable with
$\|f\|_{\tilde{L}_{p}^{q}([s,t])}:=\sup_{y\in\mathbb{R}^{d}}\bigg{\{}\int_{s}^{t}\bigg{(}\int_{B(y,1)}|f(r,x)|^{p}\text{\rm{d}}x\bigg{)}^{\frac{q}{p}}\text{\rm{d}}r\bigg{\}}^{\frac{1}{q}}<\infty,$
where $B(y,1):=\{x\in\mathbb{R}^{d}:|x-y|\leq 1\}$ is the unit ball centered at the point $y$. When $s=0$, we simply denote
$\tilde{L}_{p}^{q}(t)=\tilde{L}_{p}^{q}([0,t]),\ \ \|f\|_{\tilde{L}_{p}^{q}(t)}=\|f\|_{\tilde{L}_{p}^{q}([0,t])}.$
We take $(p,q)$ from the space
$\mathscr{K}:=\Big{\{}(p,q)\in(2,\infty]^{2},\ \frac{d}{p}+\frac{2}{q}<1\Big{\}},$
and make the following assumption where $\nabla$ is the gradient in $x\in\mathbb{R}^{d}$.
  1. (A)
    There exist $\alpha\in\mathscr{A}$, $k\in(1,\infty),K\in(0,\infty),l\in\mathbb{N},$ and
[…]
    such that the following conditions hold.
  2. $(A_{1})$
    $(\sigma_{t}\sigma^{*}_{t})(x)$ is invertible and $\sigma_{t}(x)$ is weakly differentiable in $x$ such that
    $\|\sigma\sigma^{*}\|_{\infty}+\|(\sigma\sigma^{*})^{-1}\|_{\infty}<\infty,\ \ |\nabla\sigma|\leq\sum_{i=1}^{l}f_{i},$
    $\lim_{\varepsilon\downarrow 0}\sup_{t\in[0,T],|x-x^{\prime}|\leq\varepsilon}\|(\sigma_{t}\sigma_{t}^{\ast})(x)-(\sigma_{t}\sigma_{t}^{\ast})(x^{\prime})\|=0.$
  3. $(A_{2})$
    $b_{t}(x,\mu)=b_{t}^{(0)}(x)+b_{t}^{(1)}(x,\mu),$ where for any $t\in[0,T],x,y\in\mathbb{R}^{d},\mu,\nu\in\mathscr{P}_{k},$
    $|b^{(0)}_{t}(x)|\leq f_{0}(t,x),\ \ |b_{t}^{(1)}(0,\delta_{0})|\leq K,$
    $|b_{t}^{(1)}(x,\mu)-b_{t}^{(1)}(y,\nu)|\leq K\big{\{}|x-y|+\mathbb{W}_{\alpha}(\mu,\nu)+\mathbb{W}_{k}(\mu,\nu)\big{\}}.$
We first observe that (A) implies the well-posedness of (1.1) for distributions in $\mathscr{P}_{k}$. Since $\alpha(0)=0$ and $\alpha$ is concave, there exists a constant $c>0$ such that
$\sup_{[f]_{\alpha}\leq 1}|f(x)-f(0)|\leq\alpha(|x|)\leq c+c|x|^{k},\ \ x\in\mathbb{R}^{d}.$
Thus,
(2.3) $\frac{1}{c}\mathbb{W}_{\alpha}(\mu,\nu)\leq\mathbb{W}_{k,var}(\mu,\nu):=\sup_{|f|\leq 1+|\cdot|^{k}}|\mu(f)-\nu(f)|.$
So, by [28, Theorem 3.1(1)] for $D=\mathbb{R}^{d}$, under assumption (A), (1.1) is well-posed for distributions in $\mathscr{P}_{k}$, and for any $n\geq 1$ there exists a constant $c_{n}>0$ such that
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}|^{n}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{n}(1+|X_{0}|^{n}).$
Consequently,
(2.4) $\sup_{t\in[0,T]}\|P_{t}^{*}\gamma\|_{n}^{n}\leq\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\gamma}|^{n}\Big{]}\leq c_{n}(1+\|\gamma\|_{n}^{n}),\ \ n\geq 1.$
Theorem 2.1.
Assume (A) with $k=2$. Then there exists a constant $c>0$ such that
(2.5) ${\rm Ent}(P_{t}^{\ast}\gamma|P_{t}^{\ast}\tilde{\gamma})\leq\frac{c}{t}\,\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ t\in(0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{2}.$
Example 2.2.
Let $h:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfy
$|h(x_{1},y_{1})-h(x_{2},y_{2})|\leq K_{h}|x_{1}-x_{2}|+\alpha(|y_{1}-y_{2}|),\ \ x_{1},x_{2},y_{1},y_{2}\in\mathbb{R}^{d}$
for some $\alpha\in\mathscr{A}$ and $K_{h}\geq 0$. Then $b_{t}^{(1)}(x,\mu)=\int_{\mathbb{R}^{d}}h(x,y)\mu(\text{\rm{d}}y)$ satisfies ($A_{2}$).
2.2 Proof of Theorem 2.1
Although in Theorem 2.1 we assume (A) for $k=2$, for later use we will also consider general $k\in(1,\infty)$. For any $\gamma\in\mathscr{P}_{k}$, consider the decoupled SDE of (1.1):
(2.6) $\text{\rm{d}}X^{x,\gamma}_{t}=b_{t}(X^{x,\gamma}_{t},P_{t}^{*}\gamma)\text{\rm{d}}t+\sigma_{t}(X^{x,\gamma}_{t})\text{\rm{d}}W_{t},\ \ X^{x,\gamma}_{0}=x.$
To characterize properties of (2.6), consider the following PDE for $u:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ :
(2.7) $\frac{\partial}{\partial t}u_{t}(x)+(\mathscr{L}_{t}^{\gamma}u_{t})(x)+b_{t}^{(0)}(x)=\lambda u_{t}(x),\ \ u_{T}=0,$
where $\lambda>0$ is a constant, and
(2.8) $\mathscr{L}_{t}^{\gamma}:=\frac{1}{2}\mathrm{tr}\big{\{}(\sigma_{t}\sigma_{t}^{\ast})\nabla^{2}\big{\}}+\nabla_{b_{t}(\cdot,P_{t}^{*}\gamma)}.$
By [32, Theorem 2.1] and (A), for large enough constants $\lambda,c>0$ independent of $\gamma$, (2.7) has a unique solution ${u}^{\lambda,\gamma}$ satisfying
(2.9) $\|u^{\lambda,\gamma}\|_{\infty}+\|\nabla{u}^{\lambda,\gamma}\|_{\infty}\leq\frac{1}{5},\ \ \|\nabla^{2}{u}^{\lambda,\gamma}\|_{\tilde{L}_{p_{0}}^{q_{0}}(T)}\leq c.$
So, for any $t\in[0,T]$,
(2.10) $x\mapsto\Theta^{\lambda,\gamma}_{t}(x):=x+{u}^{\lambda,\gamma}_{t}(x),\ \ x\in\mathbb{R}^{d}$
is a homeomorphism on $\mathbb{R}^{d}$.
Moreover, for any $\gamma\in\mathscr{P}_{k}$, consider
(2.11) $\text{\rm{d}}\theta^{\lambda,\gamma}_{t}(x)=b^{(1)}_{t}((\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma}_{t}(x)),P_{t}^{\ast}\gamma)\text{\rm{d}}t,\ \ t\in[0,T],\theta^{\lambda,\gamma}_{0}(x)=\Theta^{\lambda,\gamma}_{0}(x),x\in\mathbb{R}^{d},$
and let
(2.12) $\tilde{\theta}^{\lambda,\gamma}_{t}(x)=(\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma}_{t}(x)),\ \ t\in[0,T],x\in\mathbb{R}^{d}.$
Then we have
(2.13) $\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(\tilde{\theta}^{\lambda,\gamma}_{t}(x))=b^{(1)}_{t}(\tilde{\theta}^{\lambda,\gamma}_{t}(x),P_{t}^{\ast}\gamma)\text{\rm{d}}t,\ \ t\in[0,T],\tilde{\theta}^{\lambda,\gamma}_{0}(x)=x\in\mathbb{R}^{d}.$
Lemma 2.3.
Let $\sigma$ and $b$ satisfy (A). The the following assertions hold.
(1) The SDE (2.6) is well-posed and for any $v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}$, the limit
$\nabla_{v}X_{t}^{x,\gamma}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{x+\varepsilon v,\gamma}-X_{t}^{x,\gamma}}{\varepsilon},\ \ \ \ t\in[0,T]$
exists in $L^{p}(\Omega\rightarrow C([0,T];\mathbb{R}^{d});\mathbb{P})$ for any $p\geq 1$, and there exists a constant $c_{p}>0$ such that
(2.14) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{v}X_{t}^{x,\gamma}|^{p}\Big{]}\leq c_{p}|v|^{p},\ \ v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}.$
Moreover, letting $\zeta_{s}:=\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1}$ and $P_{t}^{\gamma}$ be the associated Markov semigroup, i.e.
$P_{t}^{\gamma}f(x):=\mathbb{E}[f(X_{t}^{x,\gamma})],\ \ t\in[0,T],x\in\mathbb{R}^{d},f\in\mathscr{B}_{b}(\mathbb{R}^{d}),$
then the Bismut formula for $P^{\gamma}_{t}$ holds:
$\nabla_{v}P^{\gamma}_{t}f(x)=\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right],\ \ x,v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},t\in(0,T].$
(2) For any $p\geq 1$, there exists a constant $c_{p}>0$ such that
(2.15) $\mathbb{E}[|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|^{p}]\leq c_{p}t^{\frac{p}{2}},\ \ t\in[0,T],x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},$
Proof.
By [29, Theorem 2.1] for $\beta_{s}=\frac{s}{t}$, (A) implies (1). It remains to prove (2.15). To this end, we use Zvonkin’s transform.
By Itô’s formula (see [32, Lemma 3.3]), we derive
(2.16) $\begin{split}\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})&=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})+b_{t}^{(1)}(X^{x,\gamma}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{x,\gamma}_{t})\,\text{\rm{d}}W_{t}.\end{split}$
By (A), (2.9), there exists a constant $C>1$ such that
$C^{-1}|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|\leq\big{|}\Theta^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(\tilde{\theta}^{\lambda,\gamma}_{t}(x))\big{|}\leq C|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|,$
$\big{|}b_{t}^{(1)}(X^{x,\gamma}_{t})-b_{t}^{(1)}(\tilde{\theta}^{\lambda,\gamma}_{t}(x))\big{|}\leq C|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|,$
$\big{|}\lambda u^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})\big{|}+\big{\|}\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{x,\gamma}_{t})\big{\|}\leq C,\ \ (t,x,\gamma)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}.$
This together with (2.13), (2.16) and Gronwall’s inequality implies (2.15). ∎
To verify (2.5), in the following Lemma 2.4 and Lemma 2.6 we will prove
(2.17) $\int_{0}^{t}\{\mathbb{W}_{\alpha}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})+\mathbb{W}_{2}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})\}^{2}\text{\rm{d}}s\leq c\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ t\in[0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{2}$
for some constant $c>0$.
Lemma 2.4.
Assume (A). Then there exists a constant $c>0$ such that
(2.18) $\mathbb{W}_{k}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{k}(\gamma,\tilde{\gamma})+c\int_{0}^{t}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{\ast}\tilde{\gamma})\text{\rm{d}}s,\ \ t\in[0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
Proof.
We take $\mathscr{F}_{0}$-measurable random variables $X_{0}^{\gamma},X_{0}^{\tilde{\gamma}}$ such that
(2.19) $\mathscr{L}_{X_{0}^{\gamma}}=\gamma,\ \ \mathscr{L}_{X_{0}^{\tilde{\gamma}}}=\tilde{\gamma},\ \ \mathbb{W}_{k}(\gamma,\tilde{\gamma})^{k}=\mathbb{E}[|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k}].$
Recall that $\Theta^{\lambda,\gamma}_{t}$ is defined in (2.10). By (2.7), (2.8) and Itô’s formula, we derive
(2.20) $\begin{split}\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{\gamma}_{t})&=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{\gamma}_{t})+b_{t}^{(1)}(X^{\gamma}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{\gamma}_{t})\,\text{\rm{d}}W_{t},\end{split}$
and
(2.21) $\begin{split}&\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})+b_{t}^{(1)}(X^{\tilde{\gamma}}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t\\ &\qquad+\nabla_{b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\tilde{\gamma})-b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\gamma)}\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{\tilde{\gamma}}_{t})\,\text{\rm{d}}W_{t}.\end{split}$
Combining this with (2.20) and (A), we prove the desired estimate by using the maximal functional inequality, Khasminskii’s estimate and stochastic Gronwall’s inequality, see for instance the proof of [15, Lemma 2.1] for details. Below we simply outline the procedure.
By $(A_{2})$ we have
$|b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\tilde{\gamma})-b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\gamma)|+|b_{t}^{(1)}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\gamma)-b_{t}^{(1)}(X^{\gamma}_{t},P_{t}^{\ast}\gamma)|$
$\leq K\big{\{}|X^{\gamma}_{t}-X^{\tilde{\gamma}}_{t}|+\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{t}^{*}\gamma,P_{t}^{*}\tilde{\gamma})\big{\}}.$
Combining this with (2.20), (2.21), $(A_{1})$, the maximal functional inequality and Khasminskii’s estimate (see [30, Lemma 2.1 and Lemma 4.1]), we derive
$\text{\rm{d}}\big{|}\Theta^{\lambda,\gamma}_{t}(X^{\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\big{|}^{k+1}\leq\text{\rm{d}}M_{t}+|X^{\gamma}_{t}-X^{\tilde{\gamma}}_{t}|^{k+1}\text{\rm{d}}\mathscr{L}_{t}$
$\qquad+c_{1}\big{\{}\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{t}^{*}\gamma,P_{t}^{*}\tilde{\gamma})\big{\}}|\Theta^{\lambda,\gamma}_{t}(X^{\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\big{|}^{k}\text{\rm{d}}t,$
where $c_{1}>0$ is a constant, $\mathscr{L}_{t}$ is an adapted increasing process with $\mathbb{E}[\text{\rm{e}}^{\delta\mathscr{L}_{T}}]<\infty$ for any $\delta>0$, and $M_{t}$ is a local martingale. Since (2.9) implies
$\frac{1}{2}|X^{\tilde{\gamma}}_{t}-X^{\gamma}_{t}|\leq|\Theta^{\lambda,\gamma}_{t}(X^{\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\big{|}\leq 2|X^{\tilde{\gamma}}_{t}-X^{\gamma}_{t}|,$
by the stochastic Gronwall inequality (see [31, Lemma 3.7]), we find a constant $c_{2}>1$ such that
$\bigg{\{}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{\}}^{1+k^{-1}}-c_{2}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k+1}$
$\leq c_{2}\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})\big{\}}\mathbb{E}\Big{[}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\text{\rm{d}}s,\ \ t\in[0,T].$
So, there exists a constant $c_{3}>0$ such that for any $t\in[0,T]$,
$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}-c_{2}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k}$
$\leq c_{2}\bigg{(}\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})\big{\}}\mathbb{E}\Big{[}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\text{\rm{d}}s\bigg{)}^{\frac{k}{k+1}}$
$\leq\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}+c_{3}\bigg{(}\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})\big{\}}\text{\rm{d}}s\bigg{)}^{k}.$
This together with (2.19) yields
$\mathbb{W}_{k}(P_{t}^{*}\gamma,P_{t}^{*}\tilde{\gamma})\leq\sup_{s\in[0,t]}\big{(}\mathbb{E}[|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}]\big{)}^{\frac{1}{k}}$
$\leq(2c_{2})^{\frac{1}{k}}\mathbb{W}_{k}(\gamma,\tilde{\gamma})+(2c_{3})^{\frac{1}{k}}\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})\big{\}}\text{\rm{d}}s,\ \ t\in[0,T].$
By Gronwall’s inequality, this implies the desired estimate for some constant $c>0$. ∎
To build the estimate (1.2), we present the following Hölder inequality for concave functions.
Lemma 2.5.
Let $\alpha:[0,\infty)\rightarrow[0,\infty)$ be concave. For any non-negative random variables $\xi$ and $\eta$,
$\mathbb{E}[\alpha(\xi)\eta]\leq\|\eta\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\xi\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1.$
Consequently, for any random variable $\xi$ on $\mathbb{R}^{d}$, $f\in C(\mathbb{R}^{d};\mathbb{B})$ for a Banach space $(\mathbb{B},\|\cdot\|_{\mathbb{B}})$ with $[f]_{\alpha}<\infty$, and any real random variable $\eta$ with $\mathbb{E}[\eta]=0$,
(2.22) $\big{\|}\mathbb{E}[f(\xi)\eta]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\|\eta\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\xi-y\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1,y\in\mathbb{R}^{d}.$
Proof.
Since the assertion holds trivially for $p=\infty$, we only prove for $p<\infty$. It suffices to prove for $\mathbb{E}[|\eta|^{p}]\in(0,\infty)$. Let $\mathbb{Q}:=\frac{\eta}{\mathbb{E}[\eta]}\mathbb{P}.$ By Jensen’s and Hölder’s inequalities, and using (2.1), we obtain
$\mathbb{E}[\alpha(\xi)\eta]=\mathbb{E}[|\eta|]\mathbb{E}_{\mathbb{Q}}[\alpha(\xi)]\leq\mathbb{E}[|\eta|]\alpha(\mathbb{E}_{\mathbb{Q}}[\xi])\leq\mathbb{E}[|\eta|]\alpha\bigg{(}\frac{(\mathbb{E}[|\eta|^{p}])^{\frac{1}{p}}}{\mathbb{E}[|\eta|]}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\bigg{)}$
$\leq\mathbb{E}[|\eta|]\bigg{\{}\frac{(\mathbb{E}[|\eta|^{p}])^{\frac{1}{p}}}{\mathbb{E}[|\eta|]}\ \alpha\Big{(}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}\bigg{\}}=\big{(}\mathbb{E}[|\eta|^{p}]\big{)}^{\frac{1}{p}}\alpha\Big{(}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}.$
Then the second inequality follows by noting that $\mathbb{E}[\eta]=0$ implies
$\big{\|}\mathbb{E}[f(\xi)\eta]\big{\|}_{\mathbb{B}}=\big{\|}\mathbb{E}[\{f(\xi)-f(y)\}\eta]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\mathbb{E}[\alpha(|\xi-y|)|\eta|].$
Noting that $X_{t}^{x,\gamma}$ solves (1.1) if the initial value $x$ is random with distribution $\gamma$, by the standard Markov property of $X_{t}^{x,\gamma}$, we have
(2.23) $P_{t}f(\gamma):=\int_{\mathbb{R}^{d}}f(x)(P_{t}^{*}\gamma)(\text{\rm{d}}x)=\int_{\mathbb{R}^{d}}P^{\gamma}_{t}f(x)\gamma(\text{\rm{d}}x),\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
The following lemma provides an extension to (1.2), which combined with Lemma 2.4 implies the desired (2.17).
Lemma 2.6.
Assume (A). Then there exists a constant $c>0$ such that
$\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
Consequently, there exists a constant $c>0$ such that for any $\gamma,\tilde{\gamma}\in\mathscr{P}_{k},$
(2.24) $\sup_{t\in[0,T]}\mathbb{W}_{k}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{k}(\gamma,\tilde{\gamma}).$
Proof.
It suffices to prove the first estimate, since it implies (2.24) according to Lemma 2.4.
Let $X_{0}^{\gamma}$ and $X_{0}^{\tilde{\gamma}}$ be in (2.19). For any $\varepsilon\in[0,2]$, let
$X_{0}^{\gamma^{\varepsilon}}:=X_{0}^{\gamma}+\varepsilon(X_{0}^{\tilde{\gamma}}-X_{0}^{\gamma}),\ \ \ \gamma^{\varepsilon}:=\mathscr{L}_{X_{0}^{\gamma^{\varepsilon}}},$
and let $X_{t}^{\gamma^{\varepsilon}}$ solve (1.1) with initial value $X_{0}^{\gamma^{\varepsilon}}.$ Then
(2.25) $\gamma^{\varepsilon}(|\cdot|)\leq 2\|\gamma\|_{k}+2\|\tilde{\gamma}\|_{k},\ \ \varepsilon\in[0,2],$
(2.26) $\mathbb{W}_{k}(\gamma^{\varepsilon},\gamma^{\varepsilon+r})^{k}\leq\mathbb{E}\big{[}|X_{0}^{\gamma^{\varepsilon}}-X_{0}^{\gamma^{\varepsilon+r}}|^{k}\big{]}=r^{k}\mathbb{W}_{k}(\gamma,\tilde{\gamma})^{k},\ \ \varepsilon,r\in[0,1].$
For any $\varepsilon\geq 0$, consider the SDE
(2.27) $\text{\rm{d}}X_{t}^{x,\gamma^{\varepsilon}}=b_{t}(X_{t}^{x,\gamma^{\varepsilon}},P_{t}^{*}\gamma^{\varepsilon})\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\gamma^{\varepsilon}})\text{\rm{d}}W_{t},\ \ X_{0}^{x,\gamma^{\varepsilon}}=x,t\in[0,T].$
For any $r\in(0,1)$, let
$\eta_{t}^{\varepsilon,r}=[\sigma_{t}^{\ast}(\sigma_{t}\sigma_{t}^{\ast})^{-1}](X^{x,\gamma^{\varepsilon}}_{t})[b_{t}(X^{x,\gamma^{\varepsilon}}_{t},P_{t}^{*}\gamma^{\varepsilon+r})-b_{t}(X^{x,\gamma^{\varepsilon}}_{t},P_{t}^{*}\gamma^{\varepsilon})],\ \ t\in[0,T].$
By (A), there exists a constant $c_{1}>0$ such that
(2.28) $\sup_{t\in[0,T]}|\eta^{\varepsilon,r}_{t}|\leq c_{1}\big{\{}\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma^{\varepsilon},P_{t}^{\ast}\gamma^{\varepsilon+r})+\mathbb{W}_{k}(P_{t}^{\ast}\gamma^{\varepsilon},P_{t}^{\ast}\gamma^{\varepsilon+r})\big{\}},\ \ r,\varepsilon\in[0,1].$
By Girsanov’s theorem,
$R^{\varepsilon,r}_{t}:=\exp\left\{\int_{0}^{t}\langle\eta^{\varepsilon,r}_{s},\text{\rm{d}}W_{s}\rangle-\frac{1}{2}\int_{0}^{t}|\eta^{\varepsilon,r}_{s}|^{2}\text{\rm{d}}s\right\},\ \ t\in[0,T]$
is a martingale, and
$W_{t}^{\varepsilon,r}=W_{t}-\int_{0}^{t}\eta_{s}^{\varepsilon,r}\text{\rm{d}}s,\ \ \ t\in[0,T]$
is a Brownian motion under the probability measure $\mathbb{Q}^{\varepsilon,r}:=R^{\varepsilon,r}_{T}\mathbb{P}.$ Rewrite (2.27) as
$\text{\rm{d}}X^{x,\gamma^{\varepsilon}}_{t}=b_{t}(X^{x,\gamma^{\varepsilon}}_{t},P_{t}^{*}\gamma^{\varepsilon+r})\text{\rm{d}}t+\sigma_{t}(X^{x,\gamma^{\varepsilon}}_{t})\text{\rm{d}}W^{\varepsilon,r}_{t},\ \ X^{x,\gamma^{\varepsilon}}_{0}=x,\ \ t\in[0,T].$
By the weak uniqueness we obtain
$\mathscr{L}_{\{X^{x,\gamma^{\varepsilon}}_{t}\}_{t\in[0,T]}|\mathbb{Q}^{\varepsilon,r}}=\mathscr{L}_{\{X^{x,\gamma^{\varepsilon+r}}_{t}\}_{t\in[0,T]}},$
where $\mathscr{L}_{\cdot|\mathbb{Q}^{\varepsilon,r}}$ is the law under $\mathbb{Q}^{\varepsilon,r}$, so that
$P_{t}^{\gamma^{\varepsilon+r}}f(x)-P^{\gamma^{\varepsilon}}_{t}f(x)=\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right],\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\varepsilon,r\in(0,1].$
Hence, by (2.23), we have
$P_{t}f(\gamma^{\varepsilon+r})-P_{t}f(\gamma^{\varepsilon})=\gamma^{\varepsilon+r}(P_{t}^{\gamma^{\varepsilon+r}}f)-\gamma^{\varepsilon}(P^{\gamma^{\varepsilon}}_{t}f)$
$=\gamma^{\varepsilon+r}(P_{t}^{\gamma^{\varepsilon+r}}f-P^{\gamma^{\varepsilon}}_{t}f)+\gamma^{\varepsilon+r}(P_{t}^{\gamma^{\varepsilon}}f)-\gamma^{\varepsilon}(P^{\gamma^{\varepsilon}}_{t}f)$
$=\int_{\mathbb{R}^{d}}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\gamma^{\varepsilon+r}(\text{\rm{d}}x)+\mathbb{E}\big{[}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+r}})-P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon}})\big{]},$
so that
(2.29) $\begin{split}&\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon+r},P_{t}^{*}\gamma^{\varepsilon})^{2}=\sup_{[f]_{\alpha}\leq 1}\big{|}P_{t}f(\gamma^{\varepsilon+r})-P_{t}f(\gamma^{\varepsilon})\big{|}^{2}\leq I_{1}+I_{2},\\ &I_{1}:=2\sup_{[f]_{\alpha}\leq 1}\bigg{|}\int_{\mathbb{R}^{d}}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{|}^{2},\\ &I_{2}:=2\bigg{|}\mathbb{E}\int_{0}^{r}\frac{\text{\rm{d}}}{\text{\rm{d}}\theta}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})\text{\rm{d}}\theta\bigg{|}^{2}=2\bigg{|}\mathbb{E}\int_{0}^{r}\Big{\{}\nabla_{X_{0}^{\tilde{\gamma}}-X_{0}^{\gamma}}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})\Big{\}}\text{\rm{d}}\theta\bigg{|}^{2}.\end{split}$
Below we estimate $I_{1}$ and $I_{2}$ respectively.
By (2.28), we obtain
(2.30) $\begin{split}&\mathbb{E}|R_{t}^{\varepsilon,r}-1|^{2}=\mathbb{E}\big{[}(R_{t}^{\varepsilon,r})^{2}-1\big{]}\leq\mathrm{esssup}_{\Omega}(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}-1)\\ &\leq\mathrm{esssup}_{\Omega}\left(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s\right)\\ &\leq\psi(\varepsilon,r)\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\big{\}}\text{\rm{d}}s,\end{split}$
where for $c_{2}:=2c_{1}^{2}$,
(2.31) $\psi(\varepsilon,r):=c_{2}\text{\rm{e}}^{c_{2}\int_{0}^{T}\{\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\}\text{\rm{d}}s}.$
By (2.3) and (2.4) for $n=k$, we have
(2.32) $\bar{\psi}:=\sup_{\varepsilon,r\in[0,1]}\psi(\varepsilon,\gamma)<\infty.$
Combining this with (2.1), (2.15), (2.30) and (2.22) with $y=\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$, where $\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$ is defined in (2.12) with $\gamma^{\varepsilon}$ replacing $\gamma$, we can find constants $k_{1},k_{2}>1$ such that
$\bigg{(}\int_{\mathbb{R}^{d}}\sup_{[f]_{\alpha}\leq 1}\Big{|}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\Big{|}\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{)}^{2}$
$\leq\bigg{(}\int_{\mathbb{R}^{d}}\alpha\big{(}k_{1}t^{\frac{1}{2}}\big{)}\sup_{x}\big{(}\mathbb{E}[|R_{t}^{\varepsilon,r}-1|^{2}]\big{)}^{\frac{1}{2}}\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{)}^{2}$
$\leq\alpha\big{(}k_{1}t^{\frac{1}{2}}\big{)}^{2}\sup_{x}\mathbb{E}[|R_{t}^{\varepsilon,r}-1|^{2}]$
$\leq k_{2}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}\psi(\varepsilon,r)\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\big{\}}\text{\rm{d}}s,\ \ t\in[0,T].$
Combining this with (2.1), (2.18), (2.26), (2.25), and letting
$\Gamma_{t}(\varepsilon,r):=\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r})^{2}+\int_{0}^{t}\mathbb{W}_{\alpha}(P_{s}^{*}\gamma^{\varepsilon},P_{s}^{*}\gamma^{\varepsilon+r})^{2}\text{\rm{d}}s,$
we find a constant $c_{4}>0$ such that
(2.33) $I_{1}\leq c_{4}\alpha\big{(}\sqrt{T}\big{)}^{2}\psi(\varepsilon,r)\bigg{(}r^{2}\mathbb{W}_{k}(\gamma,\tilde{\gamma})^{2}+\int_{0}^{t}\Gamma_{s}(\varepsilon,r)\text{\rm{d}}s\bigg{)},\ \ t\in[0,T].$
Combining this with Lemma 2.3(1), (2.15), (A) and (2.22) with $y=\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$, we find constants $c_{5}>0$ such that
$\begin{split}\sup_{[f]_{\alpha}\leq 1}\big{|}\nabla P^{\gamma^{\varepsilon}}_{t}f\big{|}(x)&\leq\frac{c_{5}}{\sqrt{t}}\alpha\big{(}t^{\frac{1}{2}}\big{)}.\end{split}$
Combining this with (2.1), we find a constant $c_{6}>0$ such that
(2.34) $\begin{split}I_{2}&\leq 2\bigg{(}\mathbb{E}\bigg{[}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|\int_{0}^{r}|\nabla P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})|\text{\rm{d}}\theta\bigg{]}\bigg{)}^{2}\\ &\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}}{t}r^{2}(\mathbb{E}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|)^{2}\\ &\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}r^{2}}{t}\big{(}\mathbb{E}[|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k}]\big{)}^{\frac{2}{k}}.\end{split}$
Let
(2.35) $\tilde{\alpha}(r):=\bigg{(}\int_{0}^{r}\frac{\alpha(t)^{2}}{t}\text{\rm{d}}t\bigg{)}^{\frac{1}{2}},\ \ \ r\geq 0.$
By (2.1), we find some constant $c^{\prime}>0$ such that
(2.36) $\begin{split}&\int_{0}^{T}\frac{\alpha(rt^{\frac{1}{2}})^{2}}{t}\text{\rm{d}}t=2\int_{0}^{rT^{\frac{1}{2}}}\frac{\alpha(s)^{2}}{s}\text{\rm{d}}s\leq c^{\prime}\tilde{\alpha}(r)^{2}<\infty,\ \ r\geq 1.\end{split}$
So, (2.34) together with (2.29) and (2.33) yields that for some constant $c_{7}>0$,
(2.37) $\begin{split}&\Gamma_{t}(\varepsilon,r)\leq c_{7}r^{2}\mathbb{W}_{k}(\gamma,\tilde{\gamma})^{2}H_{t}(\varepsilon,r)+c_{7}\psi(\varepsilon,r)\int_{0}^{t}\Gamma_{s}(\varepsilon,r)\text{\rm{d}}s,\\ &H_{t}(\varepsilon,r):=\psi(\varepsilon,r)+\tilde{\alpha}\big{(}1\big{)}^{2}+\frac{\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}}{t},\ \ \ \varepsilon,r\in[0,1],t\in[0,T].\end{split}$
By Gronwall’s inequality and (2.37), for any $\varepsilon,r\in[0,1]$ we have
$\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r})^{2}\leq\Gamma_{t}(\varepsilon,r)$
$\leq c_{7}r^{2}\mathbb{W}_{k}(\gamma,\tilde{\gamma})^{2}\bigg{\{}H_{t}(\varepsilon,r)+c_{7}\psi(\varepsilon,r)\text{\rm{e}}^{c_{7}\psi(\varepsilon,r)T}\int_{0}^{t}H_{s}(\varepsilon,r)\text{\rm{d}}s\bigg{\}},\ \ t\in[0,T].$
This together with (2.18), (2.36) and (2.31)-(2.32) implies that $\psi(\varepsilon,r)$ is bounded in $(\varepsilon,r)\in[0,1]^{2}$ with $\psi(\varepsilon,r)\rightarrow c_{2}$ as $r\rightarrow 0$, so that by the dominated convergence theorem we find a constant $c>0$ such that
(2.38) $\begin{split}&\limsup_{r\downarrow 0}\frac{\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r})}{r}\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\bigg{\{}\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}}+1\bigg{\}}.\end{split}$
By the triangle inequality,
$\big{|}\mathbb{W}_{\alpha}(P_{t}^{*}\gamma,P_{t}^{*}\gamma^{\varepsilon})-\mathbb{W}_{\alpha}(P_{t}^{*}\gamma,P_{t}^{*}\gamma^{\varepsilon+r})\big{|}\leq\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r}),\ \ \varepsilon,r\in[0,1],$
so that (2.38) implies that $\mathbb{W}_{\alpha}(P_{t}^{*}\gamma,P_{t}^{*}\gamma^{\varepsilon})$ is Lipschitz continuous (hence a.e. differentiable) in $\varepsilon\in[0,1]$ for any $t\in(0,T]$, and
$\Big{|}\frac{\text{\rm{d}}}{\text{\rm{d}}\varepsilon}\mathbb{W}_{\alpha}(P_{t}^{*}\gamma,P_{t}^{*}\gamma^{\varepsilon})\Big{|}\leq\limsup_{r\downarrow 0}\frac{\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r})}{r}\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\bigg{\{}\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}}+1\bigg{\}},\ \ \varepsilon\in[0,1].$
Noting that $\gamma^{1}=\tilde{\gamma}$, this implies the desired estimate. ∎
Proof of Theorem 2.1.
Let $k=2$. According to [28, Theorem 2.5] for $D=\mathbb{R}^{d}$, see also [32, Theorem 4.1], (A) implies the following log-Harnack inequality for some constant $c_{0}>0$ and any $\gamma\in\mathscr{P}_{2}$:
$P_{t}^{\gamma}\log f(x)\leq\log P_{t}^{\gamma}f(y)+\frac{c_{0}}{t}|x-y|^{2},\ \ x,y\in\mathbb{R}^{d},t\in(0,T],f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d}).$
Then by [28, (4.13)], see also [13, Theorem 2.1], it suffices to find a constant $c>0$ such that
(2.39) $\begin{split}&\sup_{t\in(0,T]}\log\mathbb{E}[|R_{t}^{\gamma,\tilde{\gamma}}|^{2}]\leq c\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ \gamma,\tilde{\gamma}\in\mathscr{P}_{2},\end{split}$
where
$R_{t}^{\gamma,\tilde{\gamma}}:=\text{\rm{e}}^{\int_{0}^{t}\langle\eta_{s}^{\gamma,\tilde{\gamma}},\text{\rm{d}}W_{s}\rangle-\frac{1}{2}\int_{0}^{t}|\eta_{s}^{\gamma,\tilde{\gamma}}|^{2}\text{\rm{d}}s},$
$\eta_{s}^{\gamma,\tilde{\gamma}}:=\big{\{}\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1}\big{\}}(X_{s}^{\gamma})\big{\{}b_{s}(X_{s}^{\gamma},P_{s}^{*}\tilde{\gamma})-b_{s}(X_{s}^{\gamma},P_{s}^{*}\gamma)\big{\}},\ \ s\leq t\leq T.$
Noting that (A) implies
$|\eta_{s}^{\gamma,\tilde{\gamma}}|^{2}\leq c_{1}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}+\mathbb{W}_{2}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}\big{\}},\ \ s\in[0,T]$
for some constant $c_{1}>0$, we have
$\mathbb{E}[|R_{t}^{\gamma,\tilde{\gamma}}|^{2}]\leq\text{\rm{e}}^{c_{1}\int_{0}^{t}\{\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}+\mathbb{W}_{2}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}\}\text{\rm{d}}s}.$
Moreover, by (2.36) and Lemma 2.6, there exists a constant $c>0$ such that
$\sup_{t\in[0,T]}\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}+\mathbb{W}_{2}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}\big{\}}\text{\rm{d}}s\leq c\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2}.$
Therefore, (2.39) holds for some constant $c>0$.
3 Bismut Formula
Let $k\in(1,\infty)$ and denote $k^{*}:=\frac{k}{k-1}$. In this part, we establish the Bismut formula for the intrinsic derivative of $P_{t}f(\gamma)$ for $\gamma\in\mathscr{P}_{k}.$ To this end, we assume
(3.1) $b_{t}(x,\mu)=b_{t}^{(0)}(x)+B_{t}(x,\mu,\mu(V)),\ \ t\in[0,T],x\in\mathbb{R}^{d},\mu\in\mathscr{P},$
where for a Banach space $(\mathbb{B},\|\cdot\|_{\mathbb{B}})$,
$V:\mathbb{R}^{d}\rightarrow\mathbb{B},\ \ \ \ B:[0,T]\times\mathbb{R}^{d}\times\mathscr{P}\times\mathbb{B}\rightarrow\mathbb{R}^{d}$
are measurable such that $[V]_{\alpha}\leq 1$ for some $\alpha\in\mathscr{A}$, i.e. $V$ is only square root Dini continuous and hence $B_{t}(x,\mu,\mu(V))$ may be not intrinsically differentiable in $\mu$, which produces essential difficulty. The following definition is taken from [3] using the idea of [1].
Definition 3.1.
Let $f\in C(\mathscr{P}_{k};\mathbb{B})$ for a Banach space $\mathbb{B}$. The function $f$ is called intrinsically differentiable at a point $\mu\in\mathscr{P}_{k}$, if
$T_{\mu,k}:=L^{k}(\mathbb{R}^{d}\rightarrow\mathbb{R}^{d};\mu)\ni\phi\mapsto D_{\phi}^{I}f(\mu):=\lim_{\varepsilon\downarrow 0}\frac{f(\mu\circ(id+\varepsilon\phi)^{-1})-f(\mu)}{\varepsilon}\in\mathbb{B}$
is a well defined bounded linear operator. In this case, the norm of the intrinsic derivative $D^{I}f(\mu)$ is given by
$\|D^{I}f(\mu)\|_{L^{k^{*}}(\mu)}:=\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|D^{I}_{\phi}f(\mu)\|_{\mathbb{B}}.$
If moreover
$\lim_{\|\phi\|_{T_{\mu,k}}\downarrow 0}\frac{\|f(\mu\circ(id+\phi)^{-1})-f(\mu)-D_{\phi}^{I}f(\mu)\|_{\mathbb{B}}}{\|\phi\|_{T_{\mu,k}}}=0,$
then $f$ is called $L$-differentiable at $\mu$. The function $f$ is called intrinsically differentiable or $L$-differentiable on $\mathscr{P}_{k}$, if it is so at any $\mu\in\mathscr{P}_{k}$.
Recall that a real function $f$ on a Banach space $\mathbb{B}$ is called Gateaux differentiable, if for any $z\in\mathbb{B}$,
$v\mapsto\nabla_{v}^{\mathbb{B}}f(z):=\lim_{\varepsilon\downarrow 0}\frac{f(z+\varepsilon v)-f(z)}{\varepsilon}$
is a well-defined bounded linear functional. In this case we denote
$\|\nabla^{\mathbb{B}}f(z)\|_{\mathbb{B}^{*}}:=\sup_{\|v\|_{\mathbb{B}}\leq 1}|\nabla_{v}^{\mathbb{B}}f(z)|.$
Moreover, $f$ is called Fréchet differentiable if it is Gateaux differentiable and
$\lim_{\|v\|_{\mathbb{B}}\downarrow 0}\frac{|f(z+v)-f(z)-\nabla_{v}^{\mathbb{B}}f(z)|}{\|v\|_{\mathbb{B}}}=0,\ \ z\in\mathbb{B}.$
It is well-known that a Gateaux differentiable function $f$ is Fréchet differentiable provided $\nabla_{v}^{\mathbb{B}}f(z)$ is continuous in $(v,z)\in\mathbb{B}\times\mathbb{B}$.
When $\mathbb{B}=\mathbb{R}^{l}$ for some $l\geq 1$, we simply denote $\nabla^{\mathbb{B}}=\nabla$.
3.1 Main result
We will establish a Bismut formula for the intrinsic derivative of $P_{t}f$ under the following assumption.
  1. (B)
    Let $k\in(1,\infty)$ and let $b$ in (3.1).
  2. $(B_{1})$
    $b^{(0)}$ and $\sigma$ satisfy the corresponding conditions in (A).
  3. $(B_{2})$
    There exists $\alpha\in\mathscr{A}$ and $K_{V}\geq 0$ such that
    $[V]_{\alpha}:=\sup_{x\neq y}\frac{\|V(x)-V(y)\|_{\mathbb{B}}}{\alpha(|x-y|)}\leq 1,\ \ \|V(x)\|_{\mathbb{B}}\leq K_{V}(1+|x|^{k-1}),\ \ x\in\mathbb{R}^{d}.$
  4. $(B_{3})$
    For any $t\in[0,T]$, we have $B_{t}\in C(\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B})$. $B_{t}(x,\mu,z)$ is differentiable in $x$, $L$-differentiable in $\mu\in\mathscr{P}_{k}$, and Fréchet differentiable in $z\in\mathbb{B}$ such that $\nabla_{v}^{\mathbb{B}}B_{t}(x,\mu,z)$ is continuous in $(v,z)\in\mathbb{B}\times\mathbb{B}$. Next, there exists a constant $K>0$ such that
    $|B_{t}(0,\delta_{0},0)|\leq K,$
    $|\nabla B_{t}(\cdot,\mu,z)(x)|+\|D^{L}B_{t}(x,\cdot,z)(\mu)\|_{L^{k^{*}}(\mu)}+\|\nabla^{\mathbb{B}}B_{t}(x,\mu,\cdot)(z)\|_{\mathbb{B}^{*}}\leq K,$
    $\ \ (t,x,\mu,z)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B}.$
    Moreover, for any $(t,x,\mu,z)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B}$, there exists a constant $c(t,x,\mu,z)>0$ such that
    |$|\{D^{L}B_{t}(x,\cdot,z)(\mu)(y)|\leq c(t,x,\mu,z)(1+|y|^{k-1}),\ \ y\in\mathbb{R}^{d}.$
Example 3.1.
Let $l$ be a positive integer and let $\alpha(s)=s^{\varepsilon}$ for some $\varepsilon\in(0,1\wedge(k-1))$. Then $V:\mathbb{R}^{d}\rightarrow\mathbb{R}^{l}$ with
$|V(x)-V(y)|\leq\alpha(|x-y|),\ \ x,y\in\mathbb{R}^{d}$
satisfies ($B_{2}$). Moreover, let $g\in C^{1}(\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{l};\mathbb{R}^{d})$ with bounded gradient. Then $B_{t}(x,\mu,z)=\int_{\mathbb{R}^{d}}g(x,y,z)\mu(\text{\rm{d}}y)$ satisfies ($B_{3}$).
Since (B) implies (A), as explained before that under this assumption (1.1) is well-posed for distributions in $\mathscr{P}_{k}.$
For $\mu\in\mathscr{P}_{k}$, consider the decoupled SDE
(3.2) $\begin{split}&\text{\rm{d}}X_{t}^{x,\mu}=\big{\{}b_{t}^{(0)}(X_{t}^{x,\mu})+B_{t}(X_{t}^{x,\mu},P_{t}^{*}\mu,P_{t}V(\mu))\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\mu})\text{\rm{d}}W_{t},\\ &\qquad\qquad\qquad X_{0}^{x,\mu}=x,\ \ t\in[0,T].\end{split}$
To state the Bismut formula for $P_{t}f$, we introduce the quantity $I_{t}^{f}$: for fixed $t\in(0,T]$, let
(3.3) $\begin{split}I_{t}^{f}(\mu,\phi)&:=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),\\ \zeta_{s}&:=\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1},\ \ \ s\in[0,t],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}.\end{split}$
By $(B_{1})$ and (2.14), we find a constant $c>0$ such that
(3.4) $|I_{t}^{f}(\mu,\phi)|\leq\frac{c}{\sqrt{t}}\big{(}P_{t}|f|^{k^{*}}(\mu)\big{)}^{\frac{1}{k^{*}}}\|\phi\|_{L^{k}(\mu)},\ \ \mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}.$
Next, let $X_{0}^{\mu}$ be $\mathscr{F}_{0}$-measurable such that $\mathscr{L}_{X_{0}^{\mu}}=\mu$, and let $X_{t}^{\mu}$ solve (1.1) with initial value $X_{0}^{\mu}$. For any $\varepsilon\geq 0$, denote
$\mu_{\varepsilon}:=\mu\circ(id+\varepsilon\phi)^{-1},\ \ X_{0}^{\mu_{\varepsilon}}:=X_{0}^{\mu}+\varepsilon\phi(X_{0}^{\mu}).$
Let $X_{t}^{\mu_{\varepsilon}}$ solve (1.1) with initial value $X_{0}^{\mu_{\varepsilon}}$. So,
$X_{t}^{\mu}=X_{t}^{\mu_{0}},\ \ P_{t}^{*}\mu_{\varepsilon}=\mathscr{L}_{X_{t}^{\mu_{\varepsilon}}},\ \ t\in[0,T],\varepsilon\geq 0.$
Theorem 3.2.
[…]
  1. $(1)$
[…]
  2. $(2)$
    The limit
    (3.5) $\nabla_{\phi}X_{t}^{\mu}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{\mu_{\varepsilon}}-X_{t}^{\mu}}{\varepsilon},\ \ \ \ t\in[0,T]$
    exists in $L^{k}(\Omega\rightarrow C([0,T],\mathbb{R}^{d}),\mathbb{P})$, and there exists a constant $c>0$ such that
    $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{\phi}X_{t}^{\mu}|^{k}\Big{]}\leq c\|\phi\|_{L^{k}(\mu)}^{k},\ \ \mu\in\mathscr{P}_{k}.$
  3. $(3)$
    For any $t\in(0,T]$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, $P_{t}f$ is intrinsically differentiable on $\mathscr{P}_{k}$. Moreover, for any $\mu\in\mathscr{P}_{k}$ and $\phi\in T_{\mu,k}$,
    (3.6) $\begin{split}&D_{\phi}^{I}P_{t}f(\mu)=I_{t}^{f}(\mu,\phi)+\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\\ &N_{s}:=\big{\{}\nabla^{\mathbb{B}}_{D_{\phi}^{I}P_{s}V(\mu)}B_{s}(X_{s}^{\mu},P_{s}^{\ast}\mu,\cdot)\big{\}}(P_{s}V(\mu)),\\ &\tilde{N}_{s}:=\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{\ast}\mu)(X_{s}^{\mu}),\nabla_{\phi}X_{s}^{\mu}\big{\rangle}\big{]}|_{y=X_{s}^{\mu}},\end{split}$
[…]
By (3.4)-(3.6) and ($B_{3}$), we find a constant $c>0$ such that
$\|D^{I}P_{t}f(\mu)\|_{L^{k^{*}}(\mu)}\leq c\frac{\{P_{t}|f|^{k^{*}}(\mu)\}^{\frac{1}{k^{*}}}}{\sqrt{t}},\ \ t\in(0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\mu\in\mathscr{P}_{k}.$
Remark 3.3.
Note that if $\nabla V$ is continuous and bounded, then the chain rule (see for instance [19, Proposition 3.1]) implies
(3.7) $D^{I}_{\phi}P_{t}V(\mu)=\mathbb{E}\nabla_{\nabla_{\phi}X_{t}^{\mu}}V(X_{t}^{\mu}).$
However, in the present case, $V$ only satisfies $(B_{2})$ so that (3.7) does not hold and there is essential difficulty in the study of $D^{I}_{\phi}P_{t}V$.
To explain the main idea in the proof of Theorem 3.2, we first figure out a sketch. By the definition of the intrinsic derivative, we intend to calculate
(3.8) $D_{\phi}^{I}P_{t}f(\mu):=\lim_{\varepsilon\downarrow 0}\frac{P_{t}f(\mu_{\varepsilon})-P_{t}f(\mu)}{\varepsilon}=\lim_{\varepsilon\downarrow 0}\frac{\mathbb{E}[f(X_{t}^{\mu_{\varepsilon}})-f(X_{t}^{\mu})]}{\varepsilon}.$
To this end, for any $\gamma\in\mathscr{P}_{k},$ consider the decoupled SDE (3.2) for $\gamma$ replacing $\mu$, i.e.
(3.9) $\begin{split}\text{\rm{d}}X_{t}^{x,\gamma}=\big{\{}b_{t}^{(0)}(X_{t}^{x,\gamma})+&B_{t}\big{(}X_{t}^{x,\gamma},P_{t}^{\ast}\gamma,P_{t}V(\gamma)\big{)}\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\gamma})\text{\rm{d}}W_{t},\\ &t\in[0,T],X_{0}^{x,\gamma}=x\in\mathbb{R}^{d}.\end{split}$
Let
$P_{t}^{\gamma}f(x):=\mathbb{E}[f(X_{t}^{x,\gamma})],\ \ \ x\in\mathbb{R}^{d},$
$P_{t}^{\gamma}f(\tilde{\gamma}):=\int_{\mathbb{R}^{d}}P_{t}^{\gamma}f\text{\rm{d}}\tilde{\gamma},\ \ t\geq 0,f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
For $\varepsilon\geq 0$, let $X_{t}^{\mu_{\varepsilon},\gamma}$ be the solution of (3.9) with initial value $X_{0}^{\varepsilon}$, i.e,
$\begin{split}\text{\rm{d}}X_{t}^{\mu_{\varepsilon},\gamma}=\big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\gamma})+&B_{t}\big{(}X_{t}^{\mu_{\varepsilon},\gamma},P_{t}^{\ast}\gamma,P_{t}V(\gamma)\big{)}\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\gamma})\text{\rm{d}}W_{t},\\ &t\in[0,T],\ X_{0}^{\mu_{\varepsilon},\gamma}=X_{0}^{\varepsilon}.\end{split}$
Then $X_{t}^{\mu_{\varepsilon},\mu_{\varepsilon}}$ solves (1.1) with initial value $X_{0}^{\varepsilon}$, so that
$P_{t}f(\mu_{\varepsilon})=P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})=\mathbb{E}[f(X_{t}^{\mu_{\varepsilon},\mu_{\varepsilon}})],\ \ \varepsilon\geq 0,t\in[0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
Noting that $\mu_{0}=\mu$, (3.8) reduces to
(3.10) $\begin{split}&D_{\phi}^{I}P_{t}f(\mu)=\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu)}{\varepsilon}\\ &=\lim_{\varepsilon\downarrow 0}\Big{\{}\frac{P_{t}^{\mu}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu)}{\varepsilon}+\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}\Big{\}}.\end{split}$
So, to calculate $D_{\phi}^{I}P_{t}f(\mu),$ we only need to study the limits of
$J_{1}f(t,\varepsilon):=\frac{P_{t}^{\mu}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu)}{\varepsilon},\ \ J_{2}f(t,\varepsilon):=\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}.$
By Lemma 2.3(1), for any $t\in(0,T],\varepsilon\geq 0$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, we have
(3.11) $\begin{split}&\frac{\text{\rm{d}}}{\text{\rm{d}}\varepsilon}P_{t}^{\mu}f(\mu_{\varepsilon}):=\lim_{r\downarrow 0}\frac{P_{t}^{\mu}f(\mu_{\varepsilon+r})-P_{t}^{\mu}f(\mu_{\varepsilon})}{r}\\ &=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x+\varepsilon\phi(x),\mu})\frac{1}{t}\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+\varepsilon\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x)\\ &=:I_{t}^{f}(\varepsilon,\mu,\phi),\ \ t\in(0,T].\end{split}$
In particular,
(3.12) $\lim_{\varepsilon\downarrow 0}J_{1}f(t,\varepsilon)=\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu)}{\varepsilon}=I_{t}^{f}(\mu,\phi),\ \ t\in(0,T].$
Consequently, it remains to prove
$\lim_{\varepsilon\rightarrow 0}J_{2}f(t,\varepsilon)=\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}$
for $N_{s},\tilde{N}_{s}$ defined in (3.6), which involves in $D^{I}P_{t}V(\mu)$ and $\nabla_{\phi}X_{t}^{\mu}$ defined in (3.5). Therefore, we will first study $D^{I}P_{t}V(\mu)$ and $\nabla_{\phi}X_{t}^{\mu}$.
To this end, let $u_{t}=u_{t}^{\lambda_{0},\mu},\Theta_{t}=\Theta_{t}^{\lambda_{0},\mu}$ be defined in (2.7) and (2.10) for large enough $\lambda_{0}>0$ such that (2.9) holds for $\lambda=\lambda_{0}$, and let
$\mathscr{C}_{k}:=\left\{\tilde{v}\in L^{k}\big{(}\Omega\rightarrow C([0,T];\mathbb{R}^{d}),\mathbb{P}\big{)}\ \text{is\ }\mathscr{F}_{t}\text{-adapted,\ }\tilde{v}_{0}=\nabla_{\phi(X_{0}^{\mu})}\Theta_{0}(X_{0}^{\mu})\right\},$
which is complete under the metric
$\rho_{\delta}(\tilde{v}^{1},\tilde{v}^{2}):=\left(\mathbb{E}\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta kt}|\tilde{v}_{t}^{1}-\tilde{v}^{2}_{t}|^{k}\right)^{\frac{1}{k}},\ \ \tilde{v}^{1},\tilde{v}^{2}\in\mathscr{C}_{k},\delta>0.$
For any $s\in[0,T]$, $\eta\in L^{k}(\Omega\rightarrow\mathbb{R}^{d},\mathbb{P})$, define
$\psi_{s}(v):=\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{*}\mu)(X_{s}^{\mu}),(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\eta\big{\rangle}\big{]}|_{y=X_{s}^{\mu}}.$
Recall that $\tilde{\alpha}$ is defined in (2.35). For any $(v,\tilde{v})\in C([0,T];\mathbb{B})\times\mathscr{C}_{k}$, we denote
(3.13) $\begin{split}&N_{s}(v):=\nabla_{v_{s}}^{\mathbb{B}}B_{s}(X_{s}^{\mu},P_{s}^{*}\mu,\cdot)(P_{s}V(\mu)),\ \ \tilde{N}_{s}(\tilde{v}):=\psi_{s}(\tilde{v}_{s}),\end{split}$
and write
$\begin{split}&\tilde{I}_{t}^{V}(\mu,v,\tilde{v})=\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})+\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v}),\\ &\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v}):=\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})}{\{\tilde{\alpha}(s^{\frac{1}{2}})s\}^{\frac{1}{2}}}\Big{\langle}\zeta_{s}(X_{s}^{\mu})N_{s}(v),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\\ &\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v}):=\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\tilde{N}_{s}(\tilde{v}),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\ \ t\in[0,T].\end{split}$
Consider the following equation for $(v,\tilde{v})\in C([0,T];\mathbb{B})\times\mathscr{C}_{k}$:
(3.14) $\begin{split}&\text{\rm{d}}\tilde{v}_{t}=\Big{[}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}B_{t}(\cdot,P_{t}^{*}\mu,P_{t}V(\mu))+\lambda_{0}u_{t}\big{\}}(X_{t}^{\mu})\Big{]}\text{\rm{d}}t\\ &+\nabla_{\psi_{t}(\tilde{v}_{t})}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t+{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\nabla_{N_{t}(v)}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t\\ &+\Big{[}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X^{\mu}_{t})\Big{]}\,\text{\rm{d}}W_{t},\ \ \tilde{v}_{0}=\nabla_{\phi(X_{0}^{\mu})}\Theta_{0}(X_{0}^{\mu}),\\ &v_{t}={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{V}(\mu,\phi)+\tilde{I}_{t}^{V}(\mu,v,\tilde{v})\big{\}},\ \ t\in[0,T].\end{split}$
In the following two subsections, we prove the well-posedness of (3.14) and the formulas
(3.15) $\begin{split}&D_{\phi}^{I}P_{t}V(\mu)={\frac{\alpha(t^{\frac{1}{2}})}{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}}\,v_{t},\ \ t\in(0,T],\\ &\nabla_{\phi}X_{t}^{\mu}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{\mu_{\varepsilon}}-X_{t}^{\mu}}{\varepsilon}=(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t},\ \ \ t\in[0,T].\end{split}$
3.2 Well-posedness of (3.14)
Lemma 3.4.
[…]
(3.16) $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|v_{t}(\mu,\phi)\|_{\mathbb{B}}\leq{c\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ \mu\in\mathscr{P}_{k},t\in[0,T],$
and for any $j\geq 1$, there exists a constant $\tilde{c}>0$ such that
(3.17) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\tilde{v}_{t}(\mu,\phi)|^{j}|\mathscr{F}_{0}\Big{]}\leq\tilde{c}|\phi(X_{0}^{\mu})|^{j}+{\tilde{c}\|\phi\|_{L^{k}(\mu)}^{j}},\ \ \mu\in\mathscr{P}_{k},t\in[0,T].$
Proof.
By Jensen’s inequality, for estimate (3.17) we only need to consider $j\geq k.$
As in the proof of [29, Lemma 4.1], there exists a constant $\theta>1$ and $0<f\in\mathscr{B}([0,T]\times\mathbb{R}^{d})$ such that
(3.18) [$\big{\|}\nabla\big{[}\{\nabla\Theta\}\sigma\big{]}\big{\|}\leq f,\ \ \ \mathbb{E}\int_{0}^{T}f_{s}(X_{s}^{\mu})^{2\theta}\text{\rm{d}}s<\infty.$
For any $n\geq 0$, define
$\tau_{n}:=T\wedge\inf\left\{t\geq 0:\int_{0}^{t}f_{s}(X_{s}^{\mu})^{2\theta}\text{\rm{d}}s\geq n\right\},$
$\mathscr{C}_{k,n}:=\big{\{}(z_{\cdot\wedge\tau_{n}}),\ \ z\in\mathscr{C}_{k}\big{\}}.$
We have $\lim_{n\rightarrow\infty}\tau_{n}=T$. Let
$\mathscr{V}_{0}:=\big{\{}v\in C([0,T];\mathbb{B}):\ v_{0}=0\big{\}}.$
Recall that $\tilde{\alpha}$ is defined in (2.35). By (3.18), $\|\nabla B\|_{\infty}<\infty$ due to (B), and
(3.19) $\int_{0}^{t}\frac{\alpha(rs^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(rs^{\frac{1}{2}})}\text{\rm{d}}s=2\int_{0}^{rt^{\frac{1}{2}}}\frac{\alpha(s)^{2}}{s\tilde{\alpha}(s)}\text{\rm{d}}s=4\int_{0}^{rt^{\frac{1}{2}}}\tilde{\alpha}^{\prime}(s)\text{\rm{d}}s=4\tilde{\alpha}(rt^{\frac{1}{2}}),\ \ r\geq 0,$
we see that for any $(v,\tilde{v})\in\mathscr{V}_{0}\times\mathscr{C}_{k},$
$H_{s}(v,\tilde{v})$ $:=\tilde{v}_{0}+\int_{0}^{s}\nabla_{\psi_{t}(\tilde{v}_{t})}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t$
$+\int_{0}^{s}\Big{\{}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\{B_{t}(\cdot,P_{t}^{*}\mu,P_{t}V(\mu))+\lambda_{0}u_{t}\}\Big{\}}(X_{t}^{\mu})\text{\rm{d}}t$
$+\int_{0}^{s}{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\nabla_{N_{t}(v)}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t$
$+\int_{0}^{s}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X^{\mu}_{t})\,\text{\rm{d}}W_{t},\ \ s\in[0,T]$
is a continuous adapted process on $\mathbb{R}^{d}$ such that
$H_{\cdot\land\tau_{n}}:\mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{C}_{k,n},\ \ n\geq 1.$
On the other hand, let
(3.20) $\begin{split}&\bar{H}_{t}(v,\tilde{v}):={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{V}(\mu,\phi)+\tilde{I}_{t}^{V}(\mu,v,\tilde{v})\big{\}}.\end{split}$
We claim that it suffices to prove
  1. (i)
    for each $n\geq 1$, the map
    $\Phi^{(n)}:=(\bar{H},H_{\cdot\land\tau_{n}}):\ \mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{V}_{0}\times\mathscr{C}_{k,n}$
    is well-defined and has a unique fixed point $(v^{(n)},{\tilde{v}}^{(n)})$.
  2. (ii)
    there exists a constant $c>0$ such that
    $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|v_{t}^{(n)}\|_{\mathbb{B}}\leq c{\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ (t,\mu)\in[0,T]\times\mathscr{P}_{k},n\geq 1,$
    and for any $j\geq k$, there exists a constant $\tilde{c}>0$ such that
    $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|{\tilde{v}}_{t}^{(n)}|^{j}|\mathscr{F}_{0}\Big{]}\leq\tilde{c}|\phi(X_{0}^{\mu})|^{j}+\tilde{c}{\|\phi\|_{L^{k}(\mu)}^{j}},\ \ (t,\mu)\in[0,T]\times\mathscr{P}_{k},n\geq 1.$
Indeed, by (i), the fixed points are consistent:
$\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{t\land\tau_{n}}=\big{(}v^{(n+i)},{\tilde{v}}^{(n+i)}\big{)}_{t\land\tau_{n}},\ \ n,i\geq 1,t\in[0,T],$
so that $\mathbb{P}$-a.s.
$\big{(}v(\mu,\phi),\tilde{v}(\mu,\phi)\big{)}:=\lim_{n\rightarrow\infty}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{\cdot\land\tau_{n}}=\sum_{n=1}^{\infty}1_{[\tau_{n-1},\tau_{n})}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{\cdot\land\tau_{n}}$
exists in $C([0,T];\mathbb{B}\times\mathbb{R}^{d})$, and by (ii) the estimates (3.16) and (3.17) hold. Combining this with
(3.21) $\Phi^{(n)}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}:=\big{(}\bar{H},H_{\cdot\land\tau_{n}}\big{)}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}=\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)},\ \ n\geq 1,$
we see that $(v(\mu,\phi),\tilde{v}(\mu,\phi))$ solves (3.14), and the uniqueness follows from that of the fixed point for $\Phi^{(n)},n\geq 1.$ Next, we will prove (i) and (ii) one by one.
(1) Proof of (i).
(a) We first verify
(3.22) $\bar{H}:\mathscr{V}_{0}\times\mathscr{C}_{k}\rightarrow\mathscr{V}_{0},$
so that $\Phi^{(n)}:\mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{V}_{0}\times\mathscr{C}_{k,n}$ for every $n\geq 1$.
By (3.3), we have
(3.23) $I_{t}^{V}(\mu,\phi)=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}\big{\{}V(X_{t}^{x,\mu})\big{\}}\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).$
Similar to (2.11)-(2.13), for any $\lambda\geq\lambda_{0}$, $\gamma\in\mathscr{P}_{k}$ and $x\in\mathbb{R}^{d}$, consider
$\text{\rm{d}}\theta^{\lambda,\gamma,B}_{t}(x)=B_{t}((\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma,B}_{t}(x)),P_{t}^{\ast}\gamma,P_{t}V(\gamma))\text{\rm{d}}t,\ \ t\in[0,T],\theta^{\lambda,\gamma,B}_{0}(x)=\Theta^{\lambda,\gamma}_{0}(x),$
and let
$\tilde{\theta}^{\lambda,\gamma,B}_{t}(x)=(\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma,B}_{t}(x)),\ \ t\in[0,T],x\in\mathbb{R}^{d}.$
Then we have
$\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(\tilde{\theta}^{\lambda,\gamma,B}_{t}(x))=B_{t}(\tilde{\theta}^{\lambda,\gamma,B}_{t}(x),P_{t}^{\ast}\gamma,P_{t}V(\gamma))\text{\rm{d}}t,\ \ t\in[0,T],\tilde{\theta}^{\lambda,\gamma,B}_{0}(x)=x\in\mathbb{R}^{d}.$
Repeating the proof of Lemma 2.3, we conclude that for any $p\geq 1$, there exists a constant $c_{p}>0$ such that
(3.24) $\mathbb{E}[|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma,B}_{t}(x)|^{p}]\leq c_{p}t^{\frac{p}{2}},\ \ t\in[0,T],x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},$
and
(3.25) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{v}X_{t}^{x,\gamma}|^{p}\Big{]}\leq c_{p}|v|^{p},\ \ v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}.$
By $[V]_{\alpha}\leq 1$, (2.22) in Lemma 2.5 for $p=2$ and $y=\tilde{\theta}^{\lambda_{0},\mu,B}_{t}(x)$, (3.25), (2.1), (3.24) and (3.23), we find a constant $c_{1}>0$ such that
(3.26) $\begin{split}\|I_{t}^{V}(\mu,\phi)\|_{\mathbb{B}}&\leq\frac{c_{1}}{\sqrt{t}}\int_{\mathbb{R}^{d}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}|\phi(x)|\mu(\text{\rm{d}}x)}\\ &{\leq\frac{c_{1}}{\sqrt{t}}\alpha\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}}.\end{split}$
To estimate $\tilde{I}_{t}^{V}(\mu,v,\tilde{v})$, noting that $X_{t}^{\mu}$ solves (3.9) for $\gamma=\mu$ and initial value $X_{0}^{\mu}$ replacing $x$, by the standard Markov property for solutions of (3.9), we have
$\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}V(X_{t}^{x,\mu})\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})}{\{\tilde{\alpha}(s^{\frac{1}{2}})s\}^{\frac{1}{2}}}}$
×$\qquad\qquad\times\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{v_{s}}^{\mathbb{B}}B_{s}(X_{s}^{x,\mu},P_{s}^{*}\mu,\cdot)(P_{s}V(\mu)),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),$
$\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v})=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}V(X_{t}^{x,\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})$
×$\qquad\qquad\times\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{\ast}\mu)(X_{s}^{\mu}),(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}\big{\rangle}\big{]}|_{y=X_{s}^{x,\mu}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).$
So, by $(B_{2})$, $(B_{3})$, (2.22) in Lemma 2.5 for $p=2$ and $y=\tilde{\theta}^{\lambda_{0},\mu,B}_{t}(x)$, (2.1), (3.24) and Jensen’s inequality, we find a constant $c_{2}>0$ such that
(3.27) $\begin{split}&\|\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})\|_{\mathbb{B}}\leq c_{2}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\bigg{(}\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}}\|v_{s}\|^{2}_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\\ &\|\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v})\|_{\mathbb{B}}\leq c_{2}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\left(\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}|^{k})^{\frac{2}{k}}\text{\rm{d}}s\right)^{\frac{1}{2}}.\end{split}$
Combining this with (3.19), (3.20) and (3.26), we find a constant $c_{3}>0$ such that
(3.28) $\begin{split}\|\bar{H}_{t}(v,\tilde{v})\|_{\mathbb{B}}\leq&\,c_{3}\|\phi\|_{L^{k}(\mu)}\sqrt{\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\\ &+c_{3}\sqrt{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}\big{(}s^{\frac{1}{2}}\big{)}}\|v_{s}\|^{2}_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\\ &+c_{3}\sqrt{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\left(\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}|^{k})^{\frac{2}{k}}\text{\rm{d}}s\right)^{\frac{1}{2}}.\end{split}$
Then (3.22) holds.
(b) We intend to prove that $\Phi^{(n)}$ in (i) has a unique fixed point in $\mathscr{V}_{0}\times\mathscr{C}_{k,n}.$ Obviously, for any $\delta>0$, $\mathscr{V}_{0}\times\mathscr{C}_{k}$ is complete under the metric
$\rho_{\delta,n}\big{(}(v,\tilde{v}),(w,\tilde{w})\big{)}:=\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta t}\|v_{t}-w_{t}\|_{\mathbb{B}}+\rho_{\delta,n}^{(k)}\big{(}\tilde{v},\tilde{w}\big{)},$
$\rho_{\delta,n}^{(k)}\big{(}\tilde{v},\tilde{w}\big{)}:=\bigg{(}\mathbb{E}\Big{[}\sup_{t\in[0,\tau_{n}\land T]}\text{\rm{e}}^{-\delta kt}|\tilde{v}_{t}-\tilde{w}_{t}|^{k}\Big{]}\bigg{)}^{\frac{1}{k}},\ \ v,w\in\mathscr{V}_{0},\tilde{v},\tilde{w}\in\mathscr{C}_{k}.$
So, it suffices to prove the contraction of $\Phi^{(n)}$ in $\rho_{\delta,n}$ for large enough $\delta>0$.
By the definition of $H_{t}$, $\|\nabla(B+\lambda_{0}u)\|_{\infty}<\infty$, and repeating step (2) in the proof of [29, Lemma 4.1], we find a function $c_{n}:(0,\infty)\rightarrow(0,\infty)$ with $c_{n}(\delta)\rightarrow 0$ as $\delta\rightarrow\infty$, such that
(3.29) $\begin{split}&\rho_{\delta,n}^{(k)}(H(v,\tilde{v}),H(w,\tilde{w})):=\bigg{(}\mathbb{E}\Big{[}\sup_{t\in[0,\tau_{n}]}\text{\rm{e}}^{-\delta kt}|H_{t}(v,\tilde{v})-H_{t}(w,\tilde{w})|^{k}\Big{]}\bigg{)}^{\frac{1}{k}}\\ &\leq\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta t}\|v_{t}-w_{t}\|_{\mathbb{B}}\sup_{s\in[0,T]}\int_{0}^{s}\text{\rm{e}}^{-\delta(s-t)}{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\text{\rm{d}}t\\ &+c(\delta,n)\rho_{\delta,n}^{k}(\tilde{v},\tilde{w}),\ \ n\geq 1,(v,\tilde{v}),(w,\tilde{w})\in\mathscr{V}_{0}\times\mathscr{C}_{k,n},\ \delta>0.\end{split}$
Moreover, by (3.20), (3.26) and (3.27), we find a constant $c_{4}>0$ such that
$\|\bar{H}_{t}(v,\tilde{v})-\bar{H}_{t}(w,\tilde{w})\|_{\mathbb{B}}=\frac{\{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\}^{\frac{1}{2}}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\|\tilde{I}_{t}^{V}(\mu,v-w,\tilde{v}-\tilde{w})\|_{\mathbb{B}}$
$\leq c_{4}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}\|v_{s}-w_{s}\|^{2}_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}+c_{4}\left(\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}-\tilde{w}_{s}|^{k})^{\frac{2}{k}}\text{\rm{d}}s\right)^{\frac{1}{2}},$
$\qquad\ v,w\in\mathscr{V}_{0},\tilde{v},\tilde{w}\in\mathscr{C}_{k,n},t\in[0,T].$
Combining this with (3.19) and (3.29), we conclude that $\Phi^{(n)}$ is contractive in the complete metric space $(\mathscr{V}_{0}\times\mathscr{C}_{k,n},\rho_{\delta,n})$ for large enough $\delta>0$, and hence has a unique fixed point.
(2) Proof of (ii). Combining the first equation in (3.14), (B), Itô’s formula, the maximal inequality in [30, Lemma 2.1], Krylov’s estimate (see [32, Theorem 3.1]) and hence Khasminskii’s estimate (see [30, Lemma 4.1]), for any $j\geq k$ we find a constant $c_{1}>0$ such that
d $|{\tilde{v}}^{n}_{t}|^{j+1}$
$\leq c_{1}\bigg{\{}|{\tilde{v}}^{n}_{t}|^{j+1}\text{\rm{d}}\mathscr{L}_{t}+\big{(}\mathbb{E}[|\tilde{v}_{t}^{n}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}t$
$\qquad\quad+|{\tilde{v}}_{t}^{n}|^{j}{\frac{\alpha(t^{\frac{1}{2}})}{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}}\|v_{t}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}t+\text{\rm{d}}M_{t}\bigg{\}},\ \ t\in[0,\tau_{n}]$
for some local martingale $M_{t}$ and an adapted increasing process $\mathscr{L}_{t}$ with $\mathbb{E}[\text{\rm{e}}^{p\mathscr{L}_{T}}]<\infty$ for any $p>0$. Then by the stochastic Gronwall inequality (see [31, Lemma 3.7]), for any $j\geq k$, we find constants $c_{2},c_{3}>0$ such that
$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{2}|\phi(X_{0}^{\mu})|^{j}+c_{2}\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}}$
$\qquad\qquad\quad+c_{2}\left(\mathbb{E}\bigg{(}\int_{0}^{t}|{\tilde{v}}_{t}^{n}|^{j}{\frac{\alpha(s^{\frac{1}{2}})}{(\tilde{\alpha}(s^{\frac{1}{2}})s)^{\frac{1}{2}}}}\|v_{s}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\bigg{)}\right)^{\frac{j}{j+1}}$
$\qquad\qquad\quad\leq c_{3}|\phi(X_{0}^{\mu})|^{j}+c_{3}\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}}$
$\qquad\qquad\quad+c_{3}\left(\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})}{(\tilde{\alpha}(s^{\frac{1}{2}})s)^{\frac{1}{2}}}}\|v_{s}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}s\right)^{j}+\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}.$
Noting that (3.19) and the Schwartz inequality imply
$\bigg{(}\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})}{(\tilde{\alpha}(s^{\frac{1}{2}})s)^{\frac{1}{2}}}}\|v_{s}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{2}\leq 4{\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s,$
we find a constant $c>0$ such that
(3.30) $\begin{split}&\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c|\phi(X_{0}^{\mu})|^{j}+c\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}}\\ &\qquad+c\left({\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\right)^{\frac{j}{2}},\ \ t\in[0,T].\end{split}$
Choosing $j=k$, taking expectations to both sides, and noting that
(3.31) $\begin{split}&c\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{k+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{k}{k+1}}\\ &\leq c_{4}\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\leq\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}+\frac{c_{4}^{2}}{2}\int_{0}^{t}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\text{\rm{d}}s\end{split}$
holds for some constant $c_{4}>c$, we derive
$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}\leq$  $\,2c_{4}\|\phi\|_{L^{k}(\mu)}^{k}+c_{4}^{2}\int_{0}^{t}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\text{\rm{d}}s$
$+2c_{4}\left({\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\right)^{\frac{k}{2}}.$
So, Gronwall’s inequality yields
(3.32) $\begin{split}&\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}\leq 2c_{4}\text{\rm{e}}^{c_{4}^{2}T}\bigg{\{}\|\phi\|_{L^{k}(\mu)}^{k}+\left({\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\right)^{\frac{k}{2}}\bigg{\}}.\end{split}$
On the other hand, by taking $(v,\tilde{v})=(v^{(n)},\tilde{v}^{(n)})$ in (3.28) and noting that $\bar{H}_{t}(v,\tilde{v})=v_{t}^{(n)}$ due to (3.21), we find a constant $c_{5}>0$ such that
$\|v_{t}^{(n)}\|_{\mathbb{B}}^{2}\leq c_{5}\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)}^{2}+c_{5}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}\|v_{s}^{(n)}\|^{2}_{\mathbb{B}}\text{\rm{d}}s+c_{5}\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}^{(n)}|^{k})^{\frac{2}{k}}\text{\rm{d}}s,\ \ t\in[0,T].$
Combining this with (3.19), (3.32) and Gronwall’s inequality, we find a constant $c_{6}>0$ such that for any $n\geq 1$ and $t\in[0,T]$,
$\|v_{t}^{(n)}\|_{\mathbb{B}}\leq c_{6}\|\phi\|_{L^{k}(\mu)}\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},$
$\mathbb{E}\Big{[}\sup_{s\in[0,t\land\tau_{n}]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}\leq c_{6}\|\phi\|_{L^{k}(\mu)}^{k},\ \ \mu\in\mathscr{P}_{k},\ \phi\in T_{\mu,k}.$
This together with (3.30) proves (ii). ∎
3.3 Proof of Theorem 3.2
By Lemma 3.4, the proof of Theorem 3.2(1)-(2) is completed by the following lemma.
Lemma 3.5.
[…]
$\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\bigg{\{}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\|D_{\phi}P_{t}V(\mu)\|_{\mathbb{B}}+\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{\phi}X_{t}^{\mu}|^{k}\Big{]}\bigg{\}}\leq c.$
Proof.
By Lemma 3.4, it suffices to prove (3.15). For any $t\in[0,T]$ and $\varepsilon\in(0,1]$, (3.10) for $V$ replacing $f$ implies
(3.33) $\begin{split}&v_{t}^{\varepsilon}:=\frac{P_{t}V(\mu_{\varepsilon})-P_{t}V(\mu)}{\varepsilon}=v_{t}^{\varepsilon,1}+v_{t}^{\varepsilon,2},\\ &v_{t}^{\varepsilon,1}:=\frac{P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)}{\varepsilon},\ \ v_{t}^{\varepsilon,2}:=\frac{P_{t}^{\mu_{\varepsilon}}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu_{\varepsilon})}{\varepsilon}.\end{split}$
Next, let
(3.34) $\begin{split}&\hat{v}_{t}:={\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}v_{t}(\mu,\phi),\\ &\tilde{v}_{t}:=\tilde{v}_{t}(\mu,\phi),\ \ \tilde{v}_{t}^{\varepsilon}=\frac{\Theta_{t}(X_{t}^{\mu_{\varepsilon}})-\Theta_{t}(X_{t}^{\mu})}{\varepsilon}.\end{split}$
Then (3.15) follows from
(3.35) $\lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\big{\|}v_{t}^{\varepsilon}-\hat{v}_{t}\big{\|}_{\mathbb{B}}=0,\ \ \lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{t}|^{k}\Big{]}=0.$
In the following, we first estimate $\|v_{s}^{\varepsilon,1}\|_{\mathbb{B}}$, $\|v_{s}^{\varepsilon,2}\|_{\mathbb{B}}$, $|\tilde{v}_{s}^{\varepsilon}|$, $\|v_{t}^{\varepsilon}-\hat{v}_{t}\big{\|}_{\mathbb{B}}$ and $|\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{t}|$ in (a)-(e), then we prove (3.35) in (f).
(a) For $v_{t}^{\varepsilon,1}$. Let
$V_{N}:=\frac{(N\land\|V\|_{\mathbb{B}})V}{\|V\|_{\mathbb{B}}},\ \ N\geq 1.$
By (3.11), we have
$\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu}V_{N}(\mu_{\varepsilon})-P_{t}^{\mu}V_{N}(\mu)}{\varepsilon}=I_{t}^{V_{N}}(\mu,\phi),$
$P_{t}^{\mu}V_{N}(\mu_{\varepsilon})-P_{t}^{\mu}V_{N}(\mu)=\int_{0}^{\varepsilon}\frac{\text{\rm{d}}}{\text{\rm{d}}r}P_{t}^{\mu}V_{N}(\mu_{r})\text{\rm{d}}r=\int_{0}^{\varepsilon}I_{t}^{V_{N}}(r,\mu,\phi)\text{\rm{d}}r,\ \ N\geq 1.$
Since $(B_{2})$ implies
(3.36) ${\|V_{N}\|_{\mathbb{B}}\leq\|V\|_{\mathbb{B}}\leq c(1+|\cdot|^{k-1})}$
for some constant $c>0$, by (2.14), (3.24), the same argument to (3.26), and the dominated convergence theorem for $N\rightarrow\infty$, we find a constant $c(\mu,\phi)>0$ depending on $\phi,\mu$ such that
(3.37) $\|v_{t}^{\varepsilon,1}\|_{\mathbb{B}}=\frac{\|P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)\|_{\mathbb{B}}}{\varepsilon}\leq\frac{c(\mu,\phi)\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \ \varepsilon\in(0,1],t\in(0,T],$
and
$\lim_{\varepsilon\downarrow 0}\Big{\|}\frac{P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)}{\varepsilon}-I_{t}^{V}(\mu,\phi)\Big{\|}_{\mathbb{B}}$
$\leq\lim_{N\rightarrow\infty}\lim_{\varepsilon\downarrow 0}\bigg{\{}\|I_{t}^{V-V_{N}}(\mu,\phi)\|_{\mathbb{B}}+\frac{\|P_{t}^{\mu}(V-V_{N})(\mu_{\varepsilon})-P_{t}^{\mu}(V-V_{N})(\mu)\|_{\mathbb{B}}}{\varepsilon}\bigg{\}}$
$\leq\lim_{N\rightarrow\infty}\frac{c\|\phi\|_{L^{k}(\mu)}}{\sqrt{t}}\bigg{(}\int_{\mathbb{R}^{d}}\mathbb{E}[\|(V-V_{N})(X_{t}^{x,\mu})\|_{\mathbb{B}}^{k^{*}}]\mu(\text{\rm{d}}x)\bigg{)}^{\frac{1}{k^{*}}}=0,\ \ t\in(0,T].$
Hence,
$h_{t,1}(\varepsilon):=\big{\{}v_{t}^{\varepsilon,1}-I_{t}^{V}(\mu,\phi)\big{\}}\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}$
satisfies
(3.38) $\lim_{\varepsilon\rightarrow 0}\|h_{t,1}(\varepsilon)\|_{\mathbb{B}}=0,\ \ \sup_{\varepsilon\in(0,1]}\|h_{t,1}(\varepsilon)\|_{\mathbb{B}}\leq 2c(\mu,\phi)\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},\ \ t\in[0,T].$
(b) For $v_{t}^{\varepsilon,2}$ and $v_{t}^{\varepsilon}$. Recall that $X_{t}^{\mu_{\varepsilon},\mu}$ solves (3.2) with initial value $X_{0}^{\varepsilon}$. Let
(3.39) $\begin{split}&R_{t}^{\varepsilon}:=\text{\rm{e}}^{\int_{0}^{t}\langle\eta_{s}^{\varepsilon},\text{\rm{d}}W_{s}\rangle-\frac{1}{2}\int_{0}^{t}|\eta_{s}^{\varepsilon}|^{2}\text{\rm{d}}s},\\ &\eta_{s}^{\varepsilon}:=\zeta_{s}(X_{s}^{\mu_{\varepsilon},\mu})\\ &\qquad\quad\times\big{\{}B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{\ast}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{\ast}\mu,P_{s}V(\mu))\big{\}},\ \ s,t\in[0,T].\end{split}$
Since (B) implies (A), (2.24) holds so that there exists a constant $c>0$ such that
$\sup_{x\in\mathbb{R}^{d},z\in\mathbb{B},t\in[0,T]}|B_{t}(x,P_{s}^{*}\mu_{\varepsilon},z)-B_{t}(x,P_{s}^{*}\mu,z)|\leq c\mathbb{W}_{k}(\mu_{\varepsilon},\mu)\leq c\varepsilon\|\phi\|_{L^{k}(\mu)},\ \ \varepsilon\in[0,1].$
Hence, by (B) we find a constant $c(\mu,\phi)>0$ such that
(3.40) $\begin{split}|\eta_{s}^{\varepsilon}|&\leq K\|P_{s}V(\mu_{\varepsilon})-P_{s}V(\mu)\|_{\mathbb{B}}+c\varepsilon\|\phi\|_{L^{k}(\mu)}\\ &\leq c(\mu,\phi)\varepsilon\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}\big{)},\ \ s\in[0,T],\ \varepsilon\in(0,1].\end{split}$
Then by Girsanov’s theorem,
$W_{t}^{\varepsilon}:=W_{t}-\int_{0}^{t}\eta_{s}^{\varepsilon}\text{\rm{d}}s,\ \ s\in[0,T]$
is a Brownian motion under $\mathbb{Q}:=R_{T}^{\varepsilon}\mathbb{P}$. Reformulate (3.9) with $\gamma=\mu$ as
$\text{\rm{d}}X_{t}^{\mu_{\varepsilon},\mu}$ =$=\Big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\mu})+B_{t}(X_{t}^{\mu_{\varepsilon},\mu},P_{t}^{*}\mu,P_{t}V(\mu))\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\mu})\text{\rm{d}}W_{t}$
=$=\Big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\mu})+B_{t}(X_{t}^{\mu_{\varepsilon},\mu},P_{t}^{*}\mu_{\varepsilon},P_{t}V(\mu_{\varepsilon}))\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\mu})\text{\rm{d}}W_{t}^{\varepsilon},$
$\qquad\ X_{0}^{\mu_{\varepsilon},\mu}=X_{0}^{\mu_{\varepsilon},\mu_{\varepsilon}}=X_{0}^{\mu_{\varepsilon}}=X_{0}^{\mu}+\varepsilon\phi(X_{0}^{\mu}).$
By the weak uniqueness of (3.9) with $\gamma=\mu_{\varepsilon},$ we get
(3.41) $v_{t}^{\varepsilon,2}:=\frac{P_{t}^{\mu_{\varepsilon}}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu_{\varepsilon})}{\varepsilon}=\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu_{\varepsilon},\mu})(R_{t}^{\varepsilon}-1)],\ \ t\in[0,T].$
Combinig this with $[V]_{\alpha}\leq 1,$ and using Lemma 2.5 for $p=k^{\ast}$, we obtain
(3.42) $\begin{split}&\Big{\|}v_{t}^{\varepsilon,2}-\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu})(R_{t}^{\varepsilon}-1)]\Big{\|}_{\mathbb{B}}\leq\frac{1}{\varepsilon}\mathbb{E}[\alpha(|X_{t}^{\mu}-X_{t}^{\mu_{\varepsilon},\mu}|)|R_{t}^{\varepsilon}-1|]\\ &\leq\frac{1}{\varepsilon}\Big{(}\mathbb{E}[|R_{t}^{\varepsilon}-1|^{k^{*}}]\Big{)}^{\frac{1}{k^{*}}}\alpha\Big{(}\big{(}\mathbb{E}[|X_{t}^{\mu}-X_{t}^{\mu_{\varepsilon},\mu}|^{k}]\big{)}^{\frac{1}{k}}\Big{)}.\end{split}$
By (3.24) and (3.36), we find a constant $c_{0}(\mu,\phi)>0$ such that
(3.43) $\mathbb{E}[\|V(X_{t}^{\mu_{\varepsilon},\mu})\|_{\mathbb{B}}^{k}]\leq c_{0}(\mu,\phi),\ \ t\in[0,T],\varepsilon\in(0,1].$
On the other hand, by (3.40), for any $p\geq 1$ there exists a constant $c(p,\mu,\phi)>0$ such that
(3.44) $\mathbb{E}[|R_{t}^{\varepsilon}-1|^{p}]\leq c(p,\mu,\phi)\varepsilon^{p}\bigg{(}\int_{0}^{t}\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\big{)}\text{\rm{d}}s\bigg{)}^{\frac{p}{2}},\ \ t\in[0,T],\varepsilon\in[0,1].$
Combining this with (3.41) and (3.43), we find a constant $c_{1}(\mu,\phi)>0$ such that
$\|v_{t}^{\varepsilon,2}\|_{\mathbb{B}}$ $\leq\big{(}\mathbb{E}[\|V(X_{t}^{\mu_{\varepsilon},\mu})\|_{\mathbb{B}}^{k}]\big{)}^{\frac{1}{k}}\frac{\big{(}\mathbb{E}[|R_{t}^{\varepsilon}-1|^{k^{*}}]\big{)}^{\frac{1}{k^{*}}}}{\varepsilon}$
$\leq c_{1}(\mu,\phi)\bigg{(}\int_{0}^{t}\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\big{)}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}.$
This together with (3.33) and (3.37) yields that for some constant $c_{2}(\mu,\phi)>0$,
$\|v_{t}^{\varepsilon}\|_{\mathbb{B}}^{2}\leq 2\|v_{t}^{\varepsilon,1}\|_{\mathbb{B}}^{2}+2\|v_{t}^{\varepsilon,2}\|_{\mathbb{B}}^{2}\leq c_{2}(\mu,\phi)\bigg{(}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}+\int_{0}^{t}\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)},\ \ t\in[0,T].$
By Grownwall’s inequality, we find a constant $c_{3}(\mu,\phi)>0$ such that
(3.45) $\begin{split}&\|v_{t}^{\varepsilon}\|_{\mathbb{B}}^{2}\leq c_{2}(\mu,\phi)\bigg{(}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}+\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\text{\rm{e}}^{c_{2}(\mu,\phi)(t-s)}\text{\rm{d}}s\bigg{)}\\ &\leq c_{3}(\mu,\phi)\frac{\alpha(t^{\frac{1}{2}})^{2}}{t},\ \ t\in[0,T],\ \ \varepsilon\in(0,1].\end{split}$
(c) For $|\tilde{v}_{s}^{\varepsilon}|$. Recall that $\Theta_{t}=u_{t}+id$ for $u_{t}=u_{t}^{\lambda_{0},\mu}$. Similarly to (2.16) and (2.21), by Itô’s formula for $\Theta_{t}(X_{t}^{\mu_{\varepsilon}})-\Theta_{t}(X_{t}^{\mu})$, we obtain
(3.46) $\begin{split}&\text{\rm{d}}\tilde{v}_{t}^{\varepsilon}=\frac{1}{\varepsilon}\Big{\{}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X_{t}^{\mu_{\varepsilon}})-\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X_{t}^{\mu})\Big{\}}\text{\rm{d}}W_{t}\\ &+\frac{1}{\varepsilon}\Big{[}\lambda_{0}u_{t}(X_{t}^{\mu_{\varepsilon}})-\lambda_{0}u_{t}(X_{0}^{\mu})+\nabla_{B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu_{\varepsilon},P_{t}V(\mu_{\varepsilon}))-B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu,P_{t}V(\mu))}\Theta_{t}(X_{t}^{\mu_{\varepsilon}})\Big{]}\text{\rm{d}}t\\ &+\frac{1}{\varepsilon}[B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu,P_{t}V(\mu))-B_{t}(X_{t}^{\mu},P_{t}^{*}\mu,P_{t}V(\mu))]\text{\rm{d}}t,\\ &\tilde{v}_{0}^{\varepsilon}=\frac{\Theta_{0}(X_{0}^{\mu_{\varepsilon}})-\Theta_{0}(X_{0}^{\mu})}{\varepsilon}.\end{split}$
Similarly to (3.30), by combining this with (B), (2.9), (3.45) and the stochastic Gronwall inequality, for any $j\geq 1$ we find constants $c(j),c_{4}(\mu,\phi)>0$ such that
(3.47) $\begin{split}&\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c(j)|\phi(X_{0}^{\mu})|^{j}+c_{4}(\mu,\phi)\\ &\qquad+c(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}},\ \ t\in[0,T],\varepsilon\in(0,1].\end{split}$
Letting $j=k$, taking expectation to both sides and noting that
$c(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{\frac{k+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{k}{k+1}}$
$\leq\tilde{c}(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\leq\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}+\frac{\tilde{c}(j)^{2}}{2}\int_{0}^{t}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}\text{\rm{d}}s,$
for some $\tilde{c}(j)\geq c(j)$ as in (3.31), by Gronwall’s inequality we find a constant $c_{5}(\mu,\phi)>0$ such that
(3.48) $\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}\leq c_{5}(\mu,\phi),\ \ \varepsilon\in(0,1],$
so that (3.47) implies
(3.49) $\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\tilde{v}_{s}^{\varepsilon}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{6}(\mu,\phi)+c(j)|\phi(X_{0}^{\mu})|^{j},\ \ \varepsilon\in(0,1]$
for some constant $c_{6}(\mu,\phi)>0.$
(d) For $|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}$. Now, let
$\psi_{t}^{\varepsilon}:=\,$ $\mathbb{E}\Big{[}\big{\langle}\{D^{L}B_{t}(z,\cdot,P_{t}V(\mu))\}(P_{t}^{*}\mu)(X_{t}^{\mu}),(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}^{\varepsilon}\big{\rangle}\Big{]}|_{z=X_{t}^{\mu}}$
+$\qquad+\big{\{}\nabla^{\mathbb{B}}_{v_{t}^{\varepsilon}}B_{t}(X_{t}^{\mu},P_{t}^{*}\mu,\cdot)\big{\}}(P_{t}V(\mu)).$
Then (3.46) can be reformulated as
$\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{0}^{\varepsilon}$
$=\int_{0}^{t}\Big{\{}\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\big{[}B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}\big{]}(X_{s}^{\mu})+\nabla_{\psi_{s}^{\varepsilon}}\Theta_{s}(X_{s}^{\mu})\Big{\}}\text{\rm{d}}s$
$\quad+\int_{0}^{t}\big{[}\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu})\big{]}\text{\rm{d}}W_{s}+\alpha_{t}^{\varepsilon},\ \ t\in[0,T],$
where for $t\in[0,T],$
$\alpha_{t}^{\varepsilon}:=\int_{0}^{t}\xi_{s}^{\varepsilon}\text{\rm{d}}s+\int_{0}^{t}\tilde{\xi}_{s}^{\varepsilon}\,\text{\rm{d}}W_{s},$
$\xi_{s}^{\varepsilon}:=\frac{1}{\varepsilon}\{[B_{s}(\cdot,P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))+\lambda_{0}u_{s}](X_{s}^{\mu_{\varepsilon}})-[B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}](X_{s}^{\mu})\}$
$\qquad\qquad+\nabla_{\frac{1}{\varepsilon}\{B_{s}(X_{s}^{\mu_{\varepsilon}},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon}},P_{s}^{*}\mu,P_{s}V(\mu))\}}u_{s}(X_{s}^{\mu_{\varepsilon}})$
$\qquad-\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\big{[}B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}\big{]}(X_{s}^{\mu})-\nabla_{\psi_{s}^{\varepsilon}}\Theta_{s}(X_{s}^{\mu}),$
$\tilde{\xi}_{s}^{\varepsilon}:=\frac{\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu_{\varepsilon}})-\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu})}{\varepsilon}-\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu}).$
By the same argument leading to [29, (4.23),(4.26)] and (3.30) above, (B) together with (2.9), (3.45), (3.48), (3.49) implies
$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\alpha_{s}(\varepsilon)|^{j}\Big{|}\mathscr{F}_{0}\Big{]}=0,\ \ j\geq 1,$
and by Jensen’s inequality we find a constant $c>0$ such that
$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\leq c|\tilde{v}_{0}^{\varepsilon}-\tilde{v}_{0}|^{k}+c\bigg{(}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\alpha_{s}(\varepsilon)|^{2k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{)}^{\frac{1}{2}}$
(3.50) $\quad+c\bigg{(}\int_{0}^{t}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}+c\bigg{(}\int_{0}^{t}\Big{\{}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\Big{\}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in[0,T].$
(e) For $\|v_{t}^{\varepsilon}-\hat{v}_{t}\|_{\mathbb{B}}$. Similarly to (d), for $N_{s}$ and $\tilde{N}_{s}$ in (3.13), we have
(3.51) $\begin{split}&\frac{R_{t}^{\varepsilon}-1}{\varepsilon}=\int_{0}^{t}R_{s}^{\varepsilon}\big{\langle}\varepsilon^{-1}\eta_{s}^{\varepsilon},\text{\rm{d}}W_{s}\big{\rangle}\\ &=\int_{0}^{t}R_{s}^{\varepsilon}\Big{\langle}\frac{\zeta_{s}(X_{s}^{\mu_{\varepsilon},\mu})[B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu,P_{s}V(\mu))]}{\varepsilon},\ \text{\rm{d}}W_{s}\Big{\rangle}\\ &=h_{t}(\varepsilon)+\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(v^{\varepsilon})+\tilde{N}_{s}(\tilde{v}^{\varepsilon})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle},\end{split}$
where
$h_{t}(\varepsilon):=\int_{0}^{t}\Big{\langle}$ $\zeta_{s}(X_{s}^{\mu})\Big{\{}R_{s}^{\varepsilon}\frac{B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu,P_{s}V(\mu))}{\varepsilon}$
$\qquad\qquad-N_{s}(v^{\varepsilon})-\tilde{N}_{s}(\tilde{v}^{\varepsilon})\Big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}$
satisfies
(3.52) $\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|h_{t}(\varepsilon)|^{p}\Big{|}\mathscr{F}_{0}\Big{]}=0,\ \ p\geq 1.$
Noting that $X_{t}^{\mu}=X_{t}^{\mu,\mu}$, by (3.25) we find a constant $c(\mu,\phi)>0$ such that
(3.53) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\mu_{\varepsilon},\mu}-X_{t}^{\mu}|^{k}\Big{]}\leq c(\mu,\phi)\varepsilon^{k},\ \ \varepsilon\in[0,1],\mu\in\mathscr{P}_{k}.$
Combining these with (3.42), (3.45) and (3.44), we find a constant $k_{1}(\mu,\phi)>0$ such that
(3.54) $\Big{\|}v_{t}^{\varepsilon,2}-\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu})(R_{t}^{\varepsilon}-1)]\Big{\|}_{\mathbb{B}}\leq k_{1}(\mu,\phi){{\alpha(\varepsilon)}},\ \ t\in[0,T],\ \varepsilon\in(0,1).$
By (3.14) for $v_{t}=v_{t}(\mu,\phi),\tilde{v}_{t}=\tilde{v}_{t}(\mu,\phi)$, and by the definition of $\hat{v}_{t}$ in (3.35), we have
$\hat{v}_{t}$ $=I_{t}^{V}(\mu,\phi)+\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(\hat{v})+\tilde{N}_{s}(\tilde{v})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}.$
Combining this with (3.33), (3.38), (3.52), (3.51), (3.54) and the same argument to (3.27), we find a constant $k_{2}(\mu,\phi)$ and a measurable function $\tilde{h}:(0,T]\times[0,1]\rightarrow(0,\infty)$ with
(3.55) $\sup_{\varepsilon\in(0,1],t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)\leq k_{2}(\mu,\phi),\ \ \lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)=0,$
such that
$\|v_{t}^{\varepsilon}-\hat{v}_{t}\|_{\mathbb{B}}$
(3.56) $\leq\tilde{h}_{t}(\varepsilon)+\bigg{\|}\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(v^{\varepsilon}-\hat{v})+\tilde{N}_{s}(\tilde{v}^{\varepsilon}-\tilde{v})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\bigg{\|}_{\mathbb{B}}$
$\leq\tilde{h}_{t}(\varepsilon)+k_{2}(\mu,\phi)\bigg{(}\int_{0}^{t}\Big{\{}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}+\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\big{)}^{\frac{2}{k}}\Big{\}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in(0,T].$
(f) Proof of (3.35). By (3.49)-(3.3) and Fatou’s lemma, we see that
$\begin{split}&\beta_{t}:=\limsup_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{]},\ \ \tilde{\beta}_{t}:=\limsup_{\varepsilon\rightarrow 0}\sup_{s\in(0,t]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}\end{split}$
satisfy
(3.57) $\begin{split}&\beta_{t}\leq\mathbb{E}\bigg{[}\limsup_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{]}\\ &\leq c\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\limsup_{\varepsilon\rightarrow 0}\frac{s}{\alpha(s^{\frac{1}{2}})^{2}}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}\\ &\qquad\quad+c\bigg{(}\int_{0}^{t}\Big{\{}\limsup_{\varepsilon\rightarrow 0}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\Big{\}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\\ &\leq c\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\,\tilde{\beta}_{s}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}+c\left(\int_{0}^{t}\beta_{s}^{2}\text{\rm{d}}s\right)^{\frac{1}{2}},\ \ t\in[0,T].\end{split}$
Noting that (3.16), (3.34) and (3.45) imply that $\tilde{\beta}_{t}$ satisfies
$\sup_{t\in(0,T]}\tilde{\beta}_{t}\leq\sup_{\varepsilon\in(0,1]}\sup_{s\in(0,T]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}=:\tilde{c}(\mu,\phi)<\infty,$
so that by Fatou’s lemma in (3.3) we derive from (3.55) that
$\tilde{\beta}_{t}^{2}\leq Ck_{2}(\mu,\phi)^{2}\int_{0}^{t}\Big{\{}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\tilde{\beta}_{s}^{2}+\beta_{s}^{\frac{2}{k}}\Big{\}}\text{\rm{d}}s,\ \ t\in(0,T],$
where by (2.2),
$C:=\sup_{t\in(0,T]}\frac{t}{\alpha(t^{\frac{1}{2}})^{2}}<\infty.$
Combining this with (3.57), $\int_{0}^{T}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}\text{\rm{d}}t<\infty$, and applying Gronwall’s inequality, we prove (3.35) and hence finish the proof. ∎
We are now ready to prove Theorem 3.2.
Proof of Theorem 3.2.
By (3.10) and (3.12), it suffices to prove that for any $t\in(0,T]$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$,
(3.58) $\begin{split}\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}=\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}.\end{split}$
Let $R_{t}^{\varepsilon}$ be in (3.39). By (3.41) for $f$ replacing $V$, we obtain
(3.59) $\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}=\frac{1}{\varepsilon}\mathbb{E}\big{[}f(X_{t}^{\mu_{\varepsilon},\mu})(R_{t}^{\varepsilon}-1)\big{]},\ \ t\in(0,T].$
Noting that (3.53) implies
$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\mu_{\varepsilon},\mu}-X_{t}^{\mu}|^{k}\Big{]}=0,$
while $(B_{2})$, Lemma 3.5 and (3.39) lead to
$\lim_{\varepsilon\rightarrow 0}\frac{R_{t}^{\varepsilon}-1}{\varepsilon}=\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}$
in $L^{2}(\mathbb{P})$, by taking $\varepsilon\rightarrow 0$ in (3.59) we deduce (3.58) for $f\in C_{b}(\mathbb{R}^{d})$. By an approximation argument as in [29, Proof of (2.3)] for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, this implies (3.59) for $f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$

### v4

@@
 Abstract
-The log-Harnack inequality and Bismut formula are established for McKean-Vlasov SDEs with singularities in all (time, space, distribution) variables, where the drift satisfies a standard integrability condition in time-space, and may be discontinuous in the distance induced by any Dini function.
-The main results considerably improve the existing ones for the case where the drift is $L$-differentiable and Lipschitz continuous in distribution with respect to the quadratic Wasserstein distance.
+The log-Harnack inequality and Bismut formula are established for McKean-Vlasov SDEs with singularities in all (time, space, distribution) variables, where the drift satisfies an integrability condition in time-space, and the continuity in distribution may be weaker than Dini.
+The main results considerably improve the existing ones for the case where the drift is $L$-differentiable and Lipschitz continuous in distribution with respect to the 2-Wasserstein distance.
 AMS subject Classification:  60H10, 60B05.
@@
 where $T>0$ is a fixed time, $\mathscr{L}_{X_{t}}$ is the distribution of $X_{t}$, and
-$b:[0,T]\times\mathbb{R}^{d}\times\mathscr{P}\rightarrow\mathbb{R}^{d},\ \ \sigma:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{m}$
-are measurable. Because of its wide applications, this type SDE has been intensively investigated, see for instance [4, 5, 8, 9, 16, 17, 22] and the survey [11].
+$b:[0,T]\times\mathbb{R}^{d}\times\tilde{\mathscr{P}}\rightarrow\mathbb{R}^{d},\ \ \sigma:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{m}$
+are measurable for some non-empty subspace $\tilde{\mathscr{P}}\subset\mathscr{P}$ equipped with a complete distance $\tilde{\rho}$.
+Because of its wide applications, this type SDE has been intensively investigated, see for instance [4, 5, 8, 9, 16, 17, 23] and the survey [11].
 In this paper, we study the regularity of (1.1) for distributions in
@@
 Write $b_{t}(x,\mu)=b_{t}^{(0)}(x)+b_{t}^{(1)}(x,\mu)$. According to [29], if $b^{(0)}$ satisfies some integrability condition on $(t,x)$, and there exists a constant $K_{b}\geq 0$ such that
-$|b_{t}^{(1)}(x,\mu)-b_{t}^{(1)}(y,\nu)|\leq K_{b}(|x-y|^{2}+\mathbb{W}_{2}(\mu,\nu)^{2}),\ \ (x,\mu),(y,\nu)\in\mathbb{R}^{d}\times\mathscr{P}_{2},t\in[0,T],$
+$|b_{t}^{(1)}(x,\mu)-b_{t}^{(1)}(y,\nu)|\leq K_{b}(|x-y|+\mathbb{W}_{2}(\mu,\nu)),\ \ (x,\mu),(y,\nu)\in\mathbb{R}^{d}\times\mathscr{P}_{2},t\in[0,T],$
 then there exists a constant $c>0$ such that the log-Harnack inequality
@@
 See also [14, 20, 28] for log-Harnack inequalities with more regular $b^{(0)}$, and see [18] for the dimension-free Harnack inequality with power.
-If furthermore $b_{t}^{(1)}(x,\mu)$ is $L$-differentiable in $\mu\in\mathscr{P}_{k}$ (see Definition 3.1 below), the following Bismut type formula has been established in [29] for the intrinsic derivative $D^{I}_{\phi}$ (see Definition 3.1 below):
+If furthermore $b_{t}^{(1)}(x,\mu)$ is $L$-differentiable in $\mu\in\mathscr{P}_{k}$, the following Bismut type formula has been established in [30] for the intrinsic derivative $D^{I}_{\phi}$ (see Definition 3.1 below):
 $D_{\phi}^{I}P_{t}f(\mu)=\mathbb{E}\big{[}f(X_{t}^{\mu})M_{t}^{\mu,\phi}\big{]},\ \ t\in(0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\mu\in\mathscr{P}_{k},\phi\in L^{k}(\mathbb{R}^{d}\rightarrow\mathbb{R}^{d};\mu),$
@@
 $\mathbb{W}_{\varepsilon}(\mu,\nu):=\sup\big{\{}\big{|}\mu(f)-\nu(f)\big{|}:|f(x)-f(y)|\leq|x-y|^{\varepsilon}\big{\}}$
-rather than $\mathbb{W}_{1}$, and hence also has certain singularity in the distribution variable.
+for some $\varepsilon\in(0,1)$ rather than $\mathbb{W}_{1}$, and hence also has certain singularity in the distribution variable.
 This result encourages us to establish the log-Harnack inequality and Bismut formula for McKean-Vlasov SDEs with coefficients singular in all time-spatial-distribution variables.
@@
 $\mathbb{W}_{\alpha}(\mu,\nu):=\sup\big{\{}\big{|}\mu(f)-\nu(f)\big{|}:|f(x)-f(y)|\leq\alpha(|x-y|)\big{\}},$
-where $\alpha$ is the square root of a Dini function, i.e.
-$\alpha$ is concave with $\alpha(0)=0,\alpha^{\prime}(s)>0$ for $s>0$, and $\int_{0}^{1}\frac{\alpha(s)^{2}}{s}\text{\rm{d}}s<\infty$ (Dini condition for $\alpha^{2}$).
-Thus, the continuity in the distribution variable is even weaker than the Dini continuity, so that the existing study in the literature is considerably improved.
-To make this improvement, we develop a new technique by establishing a-priori derivative estimates and Bismut formula for the decoupled singular SDEs.
-The key point is to find a constant $c>0$ such that
-(1.2) $\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{2}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \gamma,\tilde{\gamma}\in\mathscr{P}_{2},$
-and for the drift of type $b_{t}(x,\mu(V))$ with $[V]_{\alpha}:=\sup_{x\neq y}\frac{|V(x)-V(y)|}{\alpha(|x-y|)}\leq 1$,
-(1.3) $\|D^{I}P_{t}V(\mu)\|_{L^{\frac{k}{k-1}}(\mu)}\leq\frac{c\,\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},$
-where $D^{I}$ is the intrinsic derivative (see Definition 3.1 below).
+where $\alpha$ is the square root of a Dini function, i.e. it belongs to class
+[…]
+$\alpha(0)=0,\ \alpha(r)>0\ \text{for}\ r>0,\ \int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r\in(0,\infty)\bigg{\}}.$
+Noting that $\int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r<\infty$ is the Dini condition for $\alpha^{2}$, the continuity in the distribution variable is even weaker than Dini, so that the existing study in the literature is considerably improved.
+The log-Harnack inequality is established in Section 2, where a key step is to derive the estimate (Lemma 2.6 for $k=2$):
+$\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{2}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \gamma,\tilde{\gamma}\in\mathscr{P}_{2},t\in(0,T]$
+for some constant $c>0$.
+The Bismut formula for the intrinsic derivative of $P_{t}f$ is presented in Section 3, for which we develop new techniques to control the intrinsic derivative $D^{I}$ and the extrinsic derivative $D^{E}$ of the drift term in the distribution variable (Theorem 3.3(1)):
+$\|D^{I}P_{t}[D^{E}b_{t}(y,\nu)(\cdot)](\mu)\|_{L^{\frac{k}{k-1}}(\mu)}\leq\frac{c\,\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.$
 2 Log-Harnack Inequality
-Since $\mathbb{W}_{2}$ is involved in the log-Harnack inequality, in this section we mainly consider (1.1) for distributions in $\mathscr{P}_{2}$, but the drift may be discontinuous in $\mathbb{W}_{k}$ for any $k>0$.
+Since $\mathbb{W}_{2}$ is involved in the log-Harnack inequality, in this section we mainly consider (1.1) for $(\tilde{\mathscr{P}},\tilde{\rho})=(\mathscr{P}_{2},\mathbb{W}_{2})$, but the drift may be not Lipschitz continuous in $\mathbb{W}_{k}$ for any $k>0$.
 We first state the concrete assumption and the main result on the log-Harnack inequality, then present a complete proof in a separate subsection.
@@
 We will allow $b_{t}(x,\cdot)$ to be merely Lipschitz continuous in the sum of $\mathbb{W}_{2}$ and the Wasserstein distance induced by the square root of a Dini function.
-Let $\alpha$ be in the following class
-$\mathscr{A}:=\bigg{\{}\alpha:[0,\infty)\rightarrow[0,\infty)\ \text{is\ increasing\ and\ concave,\ }\alpha(0)=0,\ \int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r\in(0,\infty)\bigg{\}},$
-where $\int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r<\infty$ is the Dini condition for $\alpha^{2}$. For a (real or Banach valued) function $f$, let
-$[f]_{\alpha}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{\alpha(|x-y|)}$
-be its continuity modulus in $\alpha$. Define the Wasserstein distance induced by $\alpha$:
-$\mathbb{W}_{\alpha}(\mu,\nu):=\sup_{[f]_{\alpha}\leq 1}|\mu(f)-\nu(f)|,\ \ \mu,\nu\in\mathscr{P}_{\alpha}:=\big{\{}\mu\in\mathscr{P}:\mu(\alpha(|\cdot|))<\infty\big{\}},$
-where $f$ are real functions and $\mu(f)=\int_{\mathbb{R}^{d}}f\text{\rm{d}}\mu$.
-By the concavity of $\alpha$, $\mathbb{W}_{\alpha}$ is a complete distance on $\mathscr{P}_{\alpha}$, $\mathscr{P}_{k}\subset\mathscr{P}_{\alpha}$ for $k\geq 1$, and
+Let $\alpha\in\mathscr{A}$. Then it holds
 (2.1) $\alpha(s+t)\leq\alpha(s)+\alpha(t),\ \ \alpha(rt)\leq r\alpha(t),\ \ \\ \ s,t>0,r\geq 1.$
@@
 $\alpha^{\prime}(s+t)\leq\alpha^{\prime}(s),\ \ \frac{\text{\rm{d}}}{\text{\rm{d}}t}\alpha(rt)=r\alpha^{\prime}(rt)\leq r\alpha^{\prime}(t),\ \ s,t\geq 0,r\geq 1.$
-Since $\alpha$ is increasing with $\int_{0}^{1}\frac{\alpha(s)^{2}}{s}\text{\rm{d}}s>0,$ we have $\alpha(1)>0$ so that the second estimate in (2.1) with $r=t^{-1}$ yields
+The second estimate in (2.1) with $r=t^{-1}$ yields
 (2.2) $\alpha(t)\geq\alpha(1)t>0,\ \ \ t\in(0,1].$
@@
 We take $(p,q)$ from the space
-$\mathscr{K}:=\Big{\{}(p,q)\in(2,\infty]^{2},\ \frac{d}{p}+\frac{2}{q}<1\Big{\}},$
+$\mathscr{K}:=\Big{\{}(p,q)\in(2,\infty]^{2}:\ \ \frac{d}{p}+\frac{2}{q}<1\Big{\}},$
 and make the following assumption where $\nabla$ is the gradient in $x\in\mathbb{R}^{d}$.
   1. (A)
-    There exist $\alpha\in\mathscr{A}$, $k\in(1,\infty),K\in(0,\infty),l\in\mathbb{N},$ and
+    Let $(\tilde{\mathscr{P}},\tilde{\rho})=(\mathscr{P}_{k},\mathbb{W}_{k})$ for some $k\in(1,\infty)$. There exist $K\in(0,\infty),l\in\mathbb{N},\alpha\in\mathscr{A}$ and
 […]
@@
     $|b_{t}^{(1)}(x,\mu)-b_{t}^{(1)}(y,\nu)|\leq K\big{\{}|x-y|+\mathbb{W}_{\alpha}(\mu,\nu)+\mathbb{W}_{k}(\mu,\nu)\big{\}}.$
-We first observe that (A) implies the well-posedness of (1.1) for distributions in $\mathscr{P}_{k}$. Since $\alpha(0)=0$ and $\alpha$ is concave, there exists a constant $c>0$ such that
-$\sup_{[f]_{\alpha}\leq 1}|f(x)-f(0)|\leq\alpha(|x|)\leq c+c|x|^{k},\ \ x\in\mathbb{R}^{d}.$
+We first observe that (A) implies the well-posedness of (1.1) for distributions in $\mathscr{P}_{k}$. Let $[\cdot]_{\alpha}$ be the $\alpha$-continuity modulus defined by
+$[f]_{\alpha}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{\alpha(|x-y|)}.$
+Since $\alpha(0)=0$ and $\alpha$ is concave, there exists a constant $c>0$ such that
+$\sup_{[f]_{\alpha}\leq 1}|f(x)-f(0)|\leq\alpha(|x|)\leq\alpha(1)(1+|x|)\leq c+c|x|^{k},\ \ x\in\mathbb{R}^{d}.$
 Thus,
 (2.3) $\frac{1}{c}\mathbb{W}_{\alpha}(\mu,\nu)\leq\mathbb{W}_{k,var}(\mu,\nu):=\sup_{|f|\leq 1+|\cdot|^{k}}|\mu(f)-\nu(f)|.$
-So, by [28, Theorem 3.1(1)] for $D=\mathbb{R}^{d}$, under assumption (A), (1.1) is well-posed for distributions in $\mathscr{P}_{k}$, and for any $n\geq 1$ there exists a constant $c_{n}>0$ such that
-$\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}|^{n}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{n}(1+|X_{0}|^{n}).$
+So, by [29, Theorem 3.1(1)] for $D=\mathbb{R}^{d}$, under assumption (A), (1.1) is well-posed for distributions in $\mathscr{P}_{k}$, and for any $n\geq 1$ there exists a constant $c_{n}>0$ such that
+(2.4) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}|^{n}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{n}(1+|X_{0}|^{n}).$
 Consequently,
-(2.4) $\sup_{t\in[0,T]}\|P_{t}^{*}\gamma\|_{n}^{n}\leq\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\gamma}|^{n}\Big{]}\leq c_{n}(1+\|\gamma\|_{n}^{n}),\ \ n\geq 1.$
+(2.5) $\sup_{t\in[0,T]}\|P_{t}^{*}\gamma\|_{k}^{k}=\sup_{t\in[0,T]}(P_{t}^{*}\gamma)(|\cdot|^{k})\leq\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\gamma}|^{k}\Big{]}\leq c_{k}(1+\|\gamma\|_{k}^{k}).$
 Theorem 2.1.
@@
 Example 2.2.
-Let $h:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfy
+Let $h:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ satisfy
 $|h(x_{1},y_{1})-h(x_{2},y_{2})|\leq K_{h}|x_{1}-x_{2}|+\alpha(|y_{1}-y_{2}|),\ \ x_{1},x_{2},y_{1},y_{2}\in\mathbb{R}^{d}$
@@
 (2.7) $\text{\rm{d}}X^{x,\gamma}_{t}=b_{t}(X^{x,\gamma}_{t},P_{t}^{*}\gamma)\text{\rm{d}}t+\sigma_{t}(X^{x,\gamma}_{t})\text{\rm{d}}W_{t},\ \ X^{x,\gamma}_{0}=x.$
+By [29, Theorem 3.1(1)] for $D=\mathbb{R}^{d}$, this SDE is well-posed and (2.4) also holds for $X_{t}^{x,\gamma}$ in place of $X_{t}$, i.e.
+for any $n\geq 1$ there exists a constant $c_{n}(\gamma)>0$ such that
+(2.8) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{x,\gamma}|^{n}]\leq c_{n}(\gamma)(1+|x|^{n}),\ \ x\in\mathbb{R}^{d}.$
+Let $P_{t}^{\gamma}$ be the associated Markov semigroup, i.e.
+$P_{t}^{\gamma}f(x):=\mathbb{E}[f(X_{t}^{x,\gamma})],\ \ t\in[0,T],x\in\mathbb{R}^{d},f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
+We first present the following generalized Hölder inequality with a concave function $\alpha$.
+Lemma 2.3.
+Let $\alpha:[0,\infty)\rightarrow[0,\infty)$ be concave. Then for any non-negative random variables $\xi$ and $\eta$,
+(2.9) $\mathbb{E}[\alpha(\xi)\eta]\leq\|\eta\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\xi\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1.$
+Consequently, for any random variable $\bar{\xi}$ on $\mathbb{R}^{d}$, $f\in C(\mathbb{R}^{d};\mathbb{B})$ for a Banach space $(\mathbb{B},\|\cdot\|_{\mathbb{B}})$ with $[f]_{\alpha}<\infty$, and any real random variable $\bar{\eta}$ with $\mathbb{E}[\bar{\eta}]=0$,
+(2.10) $\big{\|}\mathbb{E}[f(\bar{\xi})\bar{\eta}]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\|\bar{\eta}\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\bar{\xi}-z\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1,z\in\mathbb{R}^{d}.$
+Proof.
+Since the assertion holds trivially for $p=\infty$, we only prove for $p<\infty$.
+It suffices to prove for $\mathbb{E}[\eta^{p}]\in(0,\infty)$.
+Let $\mathbb{Q}:=\frac{\eta}{\mathbb{E}[\eta]}\mathbb{P}.$ By Jensen’s and Hölder’s inequalities, and using the second inequality in (2.1), we obtain
+$\mathbb{E}[\alpha(\xi)\eta]=\mathbb{E}[\eta]\mathbb{E}_{\mathbb{Q}}[\alpha(\xi)]\leq\mathbb{E}[\eta]\alpha(\mathbb{E}_{\mathbb{Q}}[\xi])\leq\mathbb{E}[\eta]\alpha\bigg{(}\frac{(\mathbb{E}[\eta^{p}])^{\frac{1}{p}}}{\mathbb{E}[\eta]}\big{(}\mathbb{E}[\xi^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\bigg{)}$
+$\leq\mathbb{E}[\eta]\bigg{\{}\frac{(\mathbb{E}[\eta^{p}])^{\frac{1}{p}}}{\mathbb{E}[\eta]}\ \alpha\Big{(}\big{(}\mathbb{E}[\xi^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}\bigg{\}}=\big{(}\mathbb{E}[\eta^{p}]\big{)}^{\frac{1}{p}}\alpha\Big{(}\big{(}\mathbb{E}[\xi^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}.$
+Then the second inequality follows by noting that $\mathbb{E}[\bar{\eta}]=0$ implies
+$\big{\|}\mathbb{E}[f(\bar{\xi})\bar{\eta}]\big{\|}_{\mathbb{B}}=\big{\|}\mathbb{E}[\{f(\bar{\xi})-f(z)\}\bar{\eta}]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\mathbb{E}[\alpha(|\bar{\xi}-z|)|\bar{\eta}|].$
+Therefore, the proof is completed. ∎
 To characterize properties of (2.7), consider the following PDE for $u:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$:
@@
 where $\lambda>0$ is a constant, and
-(2.8) $\mathscr{L}_{t}^{\gamma}:=\frac{1}{2}\mathrm{tr}\big{\{}(\sigma_{t}\sigma_{t}^{\ast})\nabla^{2}\big{\}}+\nabla_{b_{t}(\cdot,P_{t}^{*}\gamma)}.$
-By [32, Theorem 2.1] and (A), for large enough constants $\lambda,c>0$ independent of $\gamma$, (2.7) has a unique solution ${u}^{\lambda,\gamma}$ satisfying
+(2.12) $\mathscr{L}_{t}^{\gamma}:=\frac{1}{2}\mathrm{tr}\big{\{}(\sigma_{t}\sigma_{t}^{\ast})\nabla^{2}\big{\}}+b_{t}(\cdot,P_{t}^{*}\gamma)\cdot\nabla.$
+By [33, Theorem 2.1] and (A), for large enough constants $\lambda,c>0$ independent of $\gamma$, (2.11) has a unique solution ${u}^{\lambda,\gamma}$ satisfying
 (2.13) $\|u^{\lambda,\gamma}\|_{\infty}+\|\nabla{u}^{\lambda,\gamma}\|_{\infty}\leq\frac{1}{5},\ \ \|\nabla^{2}{u}^{\lambda,\gamma}\|_{\tilde{L}_{p_{0}}^{q_{0}}(T)}\leq c.$
@@
 is a homeomorphism on $\mathbb{R}^{d}$.
-Moreover, for any $\gamma\in\mathscr{P}_{k}$, consider
-(2.11) $\text{\rm{d}}\theta^{\lambda,\gamma}_{t}(x)=b^{(1)}_{t}((\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma}_{t}(x)),P_{t}^{\ast}\gamma)\text{\rm{d}}t,\ \ t\in[0,T],\theta^{\lambda,\gamma}_{0}(x)=\Theta^{\lambda,\gamma}_{0}(x),x\in\mathbb{R}^{d},$
+Moreover, for any $\gamma\in\mathscr{P}_{k}$, $t\in[0,T]$, consider
+(2.15) $\text{\rm{d}}\theta^{\lambda,\gamma}_{t}(x)=b^{(1)}_{t}((\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma}_{t}(x)),P_{t}^{\ast}\gamma)\text{\rm{d}}t,\ \ \theta^{\lambda,\gamma}_{0}(x)=\Theta^{\lambda,\gamma}_{0}(x),x\in\mathbb{R}^{d},$
 and let
@@
 Lemma 2.4.
-Let $\sigma$ and $b$ satisfy (A). The the following assertions hold.
-(1) The SDE (2.6) is well-posed and for any $v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}$, the limit
-$\nabla_{v}X_{t}^{x,\gamma}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{x+\varepsilon v,\gamma}-X_{t}^{x,\gamma}}{\varepsilon},\ \ \ \ t\in[0,T]$
-exists in $L^{p}(\Omega\rightarrow C([0,T];\mathbb{R}^{d});\mathbb{P})$ for any $p\geq 1$, and there exists a constant $c_{p}>0$ such that
-(2.14) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{v}X_{t}^{x,\gamma}|^{p}\Big{]}\leq c_{p}|v|^{p},\ \ v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}.$
-Moreover, letting $\zeta_{s}:=\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1}$ and $P_{t}^{\gamma}$ be the associated Markov semigroup, i.e.
-$P_{t}^{\gamma}f(x):=\mathbb{E}[f(X_{t}^{x,\gamma})],\ \ t\in[0,T],x\in\mathbb{R}^{d},f\in\mathscr{B}_{b}(\mathbb{R}^{d}),$
-then the Bismut formula for $P^{\gamma}_{t}$ holds:
-$\nabla_{v}P^{\gamma}_{t}f(x)=\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right],\ \ x,v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},t\in(0,T].$
-(2) For any $p\geq 1$, there exists a constant $c_{p}>0$ such that
-(2.15) $\mathbb{E}[|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|^{p}]\leq c_{p}t^{\frac{p}{2}},\ \ t\in[0,T],x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},$
+Let $\sigma$ and $b$ satisfy (A). Then the following assertions hold.
+  1. $(1)$
+    For any $p\geq 1$, there exists a constant $c_{p}>0$ such that
+    (2.18) $\mathbb{E}[|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|^{p}]\leq c_{p}t^{\frac{p}{2}},\ \ t\in[0,T],x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k}.$
+  2. $(2)$
+    For any $\alpha\in\mathscr{A}$, there exists a constant $c>0$ such that the gradient estimate holds:
+    (2.19) $\begin{split}&|\nabla P^{\gamma}_{t}f|(x):=\limsup_{|y-x|\rightarrow 0}\frac{|P^{\gamma}_{t}f(y)-P^{\gamma}_{t}f(x)|}{|y-x|}\\ &\leq\frac{c\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \ \ [f]_{\alpha}\leq 1,\ x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},t\in(0,T].\end{split}$
 Proof.
-By [29, Theorem 2.1] for $\beta_{s}=\frac{s}{t}$, (A) implies (1). It remains to prove (2.15). To this end, we use Zvonkin’s transform.
-By Itô’s formula (see [32, Lemma 3.3]), we derive
+(1) We will use Zvonkin’s transform defined in (2.14). By Itô’s formula (see [33, Lemma 3.3]), we derive
 (2.20) $\begin{split}\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})&=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})+b_{t}^{(1)}(X^{x,\gamma}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{x,\gamma}_{t})\,\text{\rm{d}}W_{t}.\end{split}$
@@
 $C^{-1}|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|\leq\big{|}\Theta^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(\tilde{\theta}^{\lambda,\gamma}_{t}(x))\big{|}\leq C|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|,$
-$\big{|}b_{t}^{(1)}(X^{x,\gamma}_{t})-b_{t}^{(1)}(\tilde{\theta}^{\lambda,\gamma}_{t}(x))\big{|}\leq C|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|,$
+$\big{|}b_{t}^{(1)}(X^{x,\gamma}_{t},P_{t}^{*}\gamma)-b_{t}^{(1)}(\tilde{\theta}^{\lambda,\gamma}_{t}(x),P_{t}^{*}\gamma)\big{|}\leq C|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|,$
 $\big{|}\lambda u^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})\big{|}+\big{\|}\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{x,\gamma}_{t})\big{\|}\leq C,\ \ (t,x,\gamma)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}.$
-This together with (2.13), (2.16) and Gronwall’s inequality implies (2.15). ∎
+This together with (2.17), (2.20) and Gronwall’s inequality implies (2.18).
+(2) For any measurable $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ with $[f]_{\alpha}\leq 1$, take
+$f_{n}:=[(-n)\lor f]\land n,\ \ n\geq 1.$
+By an approximation technique, it is sufficient to prove (2.19) for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$ with $[f]_{\alpha}\leq 1$.
+According to [33, Theorem 4.1], there exists a constant $c_{0}>0$ such that for any $\gamma\in\mathscr{P}_{k}$, the log-Harnack inequality
+$P_{t}^{\gamma}\log f(x)\leq\log P_{t}^{\gamma}f(y)+\frac{c_{0}}{t}|x-y|^{2},\ \ x,y\in\mathbb{R}^{d},t\in(0,T],f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d})$
+holds, so that [27, Proposition 1.3.8] implies
+$|\nabla P^{\gamma}_{t}f|\leq\frac{\sqrt{2c_{0}}}{\sqrt{t}}\{P_{t}^{\gamma}|f|^{2}\}^{\frac{1}{2}},\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),t\in(0,T],\gamma\in\mathscr{P}_{k}.$
+Observe that for any $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$ with $[f]_{\alpha}\leq 1$,
+$|\nabla P^{\gamma}_{t}f|(x)$ $\leq\inf_{z\in\mathbb{R}}\frac{\sqrt{2c_{0}}}{\sqrt{t}}\{P_{t}^{\gamma}(|f-z|^{2})(x)\}^{\frac{1}{2}}$
+(2.21) $\leq\frac{\sqrt{2c_{0}}}{\sqrt{t}}\{\mathbb{E}(\alpha(|X_{t}^{x,\gamma}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|)^{2})\}^{\frac{1}{2}}$
+$\leq\frac{\sqrt{2c_{0}}}{\sqrt{t}}\alpha(\{\mathbb{E}(|X_{t}^{x,\gamma}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|^{2})\}^{\frac{1}{2}}),\ \ x\in\mathbb{R}^{d},t\in(0,T],$
+where in the last step, we used (2.9) for $\eta=\alpha(\xi)$ with $\xi=|X_{t}^{x,\gamma}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|$ and $p=2$. Therefore, (2.19) follows from (2.2), (2.18) and (2.1). ∎
 To verify (2.6), in the following Lemma 2.5 and Lemma 2.6 we will prove
@@
 and
-(2.21) $\begin{split}&\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})+b_{t}^{(1)}(X^{\tilde{\gamma}}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t\\ &\qquad+\nabla_{b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\tilde{\gamma})-b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\gamma)}\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{\tilde{\gamma}}_{t})\,\text{\rm{d}}W_{t}.\end{split}$
+(2.26) $\begin{split}&\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})+b_{t}^{(1)}(X^{\tilde{\gamma}}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t\\ &\qquad+\nabla\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})[b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\tilde{\gamma})-b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\gamma)]\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{\tilde{\gamma}}_{t})\,\text{\rm{d}}W_{t}.\end{split}$
 Combining this with (2.25) and (A), we prove the desired estimate by using the maximal functional inequality, Khasminskii’s estimate and stochastic Gronwall’s inequality, see for instance the proof of [15, Lemma 2.1] for details.
@@
 $\leq K\big{\{}|X^{\gamma}_{t}-X^{\tilde{\gamma}}_{t}|+\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{t}^{*}\gamma,P_{t}^{*}\tilde{\gamma})\big{\}}.$
-Combining this with (2.20), (2.21), $(A_{1})$, the maximal functional inequality and Khasminskii’s estimate (see [30, Lemma 2.1 and Lemma 4.1]), we derive
+Combining this with (2.25), (2.26), $(A_{1})$, the maximal functional inequality and Khasminskii’s estimate (see [31, Lemma 2.1 and Lemma 4.1]), we derive
 $\text{\rm{d}}\big{|}\Theta^{\lambda,\gamma}_{t}(X^{\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\big{|}^{k+1}\leq\text{\rm{d}}M_{t}+|X^{\gamma}_{t}-X^{\tilde{\gamma}}_{t}|^{k+1}\text{\rm{d}}\mathscr{L}_{t}$
@@
 $\frac{1}{2}|X^{\tilde{\gamma}}_{t}-X^{\gamma}_{t}|\leq|\Theta^{\lambda,\gamma}_{t}(X^{\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\big{|}\leq 2|X^{\tilde{\gamma}}_{t}-X^{\gamma}_{t}|,$
-by the stochastic Gronwall inequality (see [31, Lemma 3.7]), we find a constant $c_{2}>1$ such that
+by the stochastic Gronwall inequality (see [32, Lemma 3.7]), we find a constant $c_{2}>1$ such that
 $\bigg{\{}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{\}}^{1+k^{-1}}-c_{2}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k+1}$
@@
 By Gronwall’s inequality, this implies the desired estimate for some constant $c>0$. ∎
-To build the estimate (1.2), we present the following Hölder inequality for concave functions.
-Lemma 2.5.
-Let $\alpha:[0,\infty)\rightarrow[0,\infty)$ be concave. For any non-negative random variables $\xi$ and $\eta$,
-$\mathbb{E}[\alpha(\xi)\eta]\leq\|\eta\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\xi\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1.$
-Consequently, for any random variable $\xi$ on $\mathbb{R}^{d}$, $f\in C(\mathbb{R}^{d};\mathbb{B})$ for a Banach space $(\mathbb{B},\|\cdot\|_{\mathbb{B}})$ with $[f]_{\alpha}<\infty$, and any real random variable $\eta$ with $\mathbb{E}[\eta]=0$,
-(2.22) $\big{\|}\mathbb{E}[f(\xi)\eta]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\|\eta\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\xi-y\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1,y\in\mathbb{R}^{d}.$
-Proof.
-Since the assertion holds trivially for $p=\infty$, we only prove for $p<\infty$.
-It suffices to prove for $\mathbb{E}[|\eta|^{p}]\in(0,\infty)$.
-Let $\mathbb{Q}:=\frac{\eta}{\mathbb{E}[\eta]}\mathbb{P}.$ By Jensen’s and Hölder’s inequalities, and using (2.1), we obtain
-$\mathbb{E}[\alpha(\xi)\eta]=\mathbb{E}[|\eta|]\mathbb{E}_{\mathbb{Q}}[\alpha(\xi)]\leq\mathbb{E}[|\eta|]\alpha(\mathbb{E}_{\mathbb{Q}}[\xi])\leq\mathbb{E}[|\eta|]\alpha\bigg{(}\frac{(\mathbb{E}[|\eta|^{p}])^{\frac{1}{p}}}{\mathbb{E}[|\eta|]}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\bigg{)}$
-$\leq\mathbb{E}[|\eta|]\bigg{\{}\frac{(\mathbb{E}[|\eta|^{p}])^{\frac{1}{p}}}{\mathbb{E}[|\eta|]}\ \alpha\Big{(}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}\bigg{\}}=\big{(}\mathbb{E}[|\eta|^{p}]\big{)}^{\frac{1}{p}}\alpha\Big{(}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}.$
-Then the second inequality follows by noting that $\mathbb{E}[\eta]=0$ implies
-$\big{\|}\mathbb{E}[f(\xi)\eta]\big{\|}_{\mathbb{B}}=\big{\|}\mathbb{E}[\{f(\xi)-f(y)\}\eta]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\mathbb{E}[\alpha(|\xi-y|)|\eta|].$
 Noting that $X_{t}^{x,\gamma}$ solves (1.1) if the initial value $x$ is random with distribution $\gamma$, by the standard Markov property of $X_{t}^{x,\gamma}$, we have
 (2.27) $P_{t}f(\gamma):=\int_{\mathbb{R}^{d}}f(x)(P_{t}^{*}\gamma)(\text{\rm{d}}x)=\int_{\mathbb{R}^{d}}P^{\gamma}_{t}f(x)\gamma(\text{\rm{d}}x),\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
-The following lemma provides an extension to (1.2), which combined with Lemma 2.4 implies the desired (2.17).
+The following lemma provides a regularity estimate on $P_{t}^{*}$, which together with Lemma 2.5 implies the desired (2.22).
 Lemma 2.6.
 Assume (A). Then there exists a constant $c>0$ such that
-$\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
+(2.28) $\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
 Consequently, there exists a constant $c>0$ such that for any $\gamma,\tilde{\gamma}\in\mathscr{P}_{k},$
@@
 Proof.
-It suffices to prove the first estimate, since it implies (2.24) according to Lemma 2.4.
+By Lemma 2.5, (2.29) follows from (2.28) and the fact $\int_{0}^{T}\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}}\text{\rm{d}}t<\infty$. So, we only need to prove (2.28).
 Let $X_{0}^{\gamma}$ and $X_{0}^{\tilde{\gamma}}$ be in (2.24). For any $\varepsilon\in[0,2]$, let
@@
 so that
-(2.29) $\begin{split}&\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon+r},P_{t}^{*}\gamma^{\varepsilon})^{2}=\sup_{[f]_{\alpha}\leq 1}\big{|}P_{t}f(\gamma^{\varepsilon+r})-P_{t}f(\gamma^{\varepsilon})\big{|}^{2}\leq I_{1}+I_{2},\\ &I_{1}:=2\sup_{[f]_{\alpha}\leq 1}\bigg{|}\int_{\mathbb{R}^{d}}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{|}^{2},\\ &I_{2}:=2\bigg{|}\mathbb{E}\int_{0}^{r}\frac{\text{\rm{d}}}{\text{\rm{d}}\theta}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})\text{\rm{d}}\theta\bigg{|}^{2}=2\bigg{|}\mathbb{E}\int_{0}^{r}\Big{\{}\nabla_{X_{0}^{\tilde{\gamma}}-X_{0}^{\gamma}}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})\Big{\}}\text{\rm{d}}\theta\bigg{|}^{2}.\end{split}$
+(2.34) $\begin{split}&\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon+r},P_{t}^{*}\gamma^{\varepsilon})^{2}=\sup_{[f]_{\alpha}\leq 1}\big{|}P_{t}f(\gamma^{\varepsilon+r})-P_{t}f(\gamma^{\varepsilon})\big{|}^{2}\leq I_{1}+I_{2},\\ &I_{1}:=2\sup_{[f]_{\alpha}\leq 1}\bigg{|}\int_{\mathbb{R}^{d}}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{|}^{2},\\ &I_{2}:=2\sup_{[f]_{\alpha}\leq 1}\bigg{|}\mathbb{E}\big{[}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+r}})-P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon}})\big{]}\bigg{|}^{2}.\end{split}$
 Below we estimate $I_{1}$ and $I_{2}$ respectively.
 By (2.33), we obtain
-(2.30) $\begin{split}&\mathbb{E}|R_{t}^{\varepsilon,r}-1|^{2}=\mathbb{E}\big{[}(R_{t}^{\varepsilon,r})^{2}-1\big{]}\leq\mathrm{esssup}_{\Omega}(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}-1)\\ &\leq\mathrm{esssup}_{\Omega}\left(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s\right)\\ &\leq\psi(\varepsilon,r)\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\big{\}}\text{\rm{d}}s,\end{split}$
+$\mathbb{E}|R_{t}^{\varepsilon,r}-1|^{2}=\mathbb{E}\big{[}(R_{t}^{\varepsilon,r})^{2}-1\big{]}\leq\mathrm{esssup}_{\Omega}(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}-1)$
+(2.35) $\leq\mathrm{esssup}_{\Omega}\left(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s\right)$
+$\leq\psi(\varepsilon,r)\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\big{\}}\text{\rm{d}}s,$
 where for $c_{2}:=2c_{1}^{2}$,
 (2.36) $\psi(\varepsilon,r):=c_{2}\text{\rm{e}}^{c_{2}\int_{0}^{T}\{\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\}\text{\rm{d}}s}.$
-By (2.3) and (2.4) for $n=k$, we have
+By (2.3) and (2.5), we have
 (2.37) $\bar{\psi}:=\sup_{\varepsilon,r\in[0,1]}\psi(\varepsilon,\gamma)<\infty.$
-Combining this with (2.1), (2.15), (2.30) and (2.22) with $y=\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$, where $\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$ is defined in (2.12) with $\gamma^{\varepsilon}$ replacing $\gamma$, we can find constants $k_{1},k_{2}>1$ such that
+Combining this with (2.1), (2.18), (2.2) and (2.10) with $z=\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$, where $\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$ is defined in (2.16) with $\gamma^{\varepsilon}$ replacing $\gamma$, we can find constants $k_{1},k_{2}>1$ such that
 $\bigg{(}\int_{\mathbb{R}^{d}}\sup_{[f]_{\alpha}\leq 1}\Big{|}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\Big{|}\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{)}^{2}$
@@
 (2.38) $I_{1}\leq c_{4}\alpha\big{(}\sqrt{T}\big{)}^{2}\psi(\varepsilon,r)\bigg{(}r^{2}\mathbb{W}_{k}(\gamma,\tilde{\gamma})^{2}+\int_{0}^{t}\Gamma_{s}(\varepsilon,r)\text{\rm{d}}s\bigg{)},\ \ t\in[0,T].$
-Combining this with Lemma 2.3(1), (2.15), (A) and (2.22) with $y=\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$, we find constants $c_{5}>0$ such that
+By (2.19), we find a constant $c_{5}>0$ such that
 $\begin{split}\sup_{[f]_{\alpha}\leq 1}\big{|}\nabla P^{\gamma^{\varepsilon}}_{t}f\big{|}(x)&\leq\frac{c_{5}}{\sqrt{t}}\alpha\big{(}t^{\frac{1}{2}}\big{)}.\end{split}$
 Combining this with (2.1), we find a constant $c_{6}>0$ such that
-(2.34) $\begin{split}I_{2}&\leq 2\bigg{(}\mathbb{E}\bigg{[}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|\int_{0}^{r}|\nabla P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})|\text{\rm{d}}\theta\bigg{]}\bigg{)}^{2}\\ &\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}}{t}r^{2}(\mathbb{E}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|)^{2}\\ &\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}r^{2}}{t}\big{(}\mathbb{E}[|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k}]\big{)}^{\frac{2}{k}}.\end{split}$
+$I_{2}$ $\leq 2\sup_{[f]_{\alpha}\leq 1}\bigg{(}\mathbb{E}\bigg{[}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|\int_{0}^{r}|\nabla P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})|\text{\rm{d}}\theta\bigg{]}\bigg{)}^{2}$
+(2.39) $\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}}{t}r^{2}(\mathbb{E}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|)^{2}$
+$\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}r^{2}}{t}\big{(}\mathbb{E}[|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k}]\big{)}^{\frac{2}{k}}.$
 Let
@@
 $\Big{|}\frac{\text{\rm{d}}}{\text{\rm{d}}\varepsilon}\mathbb{W}_{\alpha}(P_{t}^{*}\gamma,P_{t}^{*}\gamma^{\varepsilon})\Big{|}\leq\limsup_{r\downarrow 0}\frac{\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r})}{r}\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\bigg{\{}\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}}+1\bigg{\}},\ \ \varepsilon\in[0,1].$
-Noting that $\gamma^{1}=\tilde{\gamma}$, this implies the desired estimate. ∎
+This implies (2.28) by noting that $\gamma^{1}=\tilde{\gamma}$ and $\sup_{t\in[0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\leq\frac{\sqrt{T}\vee 1}{\alpha(1)}$ due to (2.2) . ∎
 Proof of Theorem 2.1.
 Let $k=2$.
-According to [28, Theorem 2.5] for $D=\mathbb{R}^{d}$, see also [32, Theorem 4.1], (A) implies the following log-Harnack inequality for some constant $c_{0}>0$ and any $\gamma\in\mathscr{P}_{2}$:
+According to [29, Theorem 2.5] for $D=\mathbb{R}^{d}$, see also [33, Theorem 4.1], (A) implies the following log-Harnack inequality for some constant $c_{0}>0$ and any $\gamma\in\mathscr{P}_{2}$:
 $P_{t}^{\gamma}\log f(x)\leq\log P_{t}^{\gamma}f(y)+\frac{c_{0}}{t}|x-y|^{2},\ \ x,y\in\mathbb{R}^{d},t\in(0,T],f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d}).$
-Then by [28, (4.13)], see also [13, Theorem 2.1], it suffices to find a constant $c>0$ such that
+Then by [29, (4.13)], see also [13, Theorem 2.1], it suffices to find a constant $c>0$ such that
 (2.44) $\begin{split}&\sup_{t\in(0,T]}\log\mathbb{E}[|R_{t}^{\gamma,\tilde{\gamma}}|^{2}]\leq c\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ \gamma,\tilde{\gamma}\in\mathscr{P}_{2},\end{split}$
@@
 Let $k\in(1,\infty)$ and denote $k^{*}:=\frac{k}{k-1}$.
-In this part, we establish the Bismut formula for the intrinsic derivative of $P_{t}f(\gamma)$ for $\gamma\in\mathscr{P}_{k}.$ To this end, we assume
-(3.1) $b_{t}(x,\mu)=b_{t}^{(0)}(x)+B_{t}(x,\mu,\mu(V)),\ \ t\in[0,T],x\in\mathbb{R}^{d},\mu\in\mathscr{P},$
-where for a Banach space $(\mathbb{B},\|\cdot\|_{\mathbb{B}})$,
-$V:\mathbb{R}^{d}\rightarrow\mathbb{B},\ \ \ \ B:[0,T]\times\mathbb{R}^{d}\times\mathscr{P}\times\mathbb{B}\rightarrow\mathbb{R}^{d}$
-are measurable such that $[V]_{\alpha}\leq 1$ for some $\alpha\in\mathscr{A}$, i.e.
-$V$ is only square root Dini continuous and hence $B_{t}(x,\mu,\mu(V))$ may be not intrinsically differentiable in $\mu$, which produces essential difficulty.
-The following definition is taken from [3] using the idea of [1].
+In this part, we consider the SDE (1.1) with $(\tilde{\mathscr{P}},\tilde{\rho})=(\mathscr{P}_{k},\mathbb{W}_{k})$, where as in $(A_{2})$ the drift $b$ is decomposed as
+(3.1) $b_{t}(x,\nu)=b_{t}^{(0)}(x)+b_{t}^{(1)}(x,\nu),\ \ t\in[0,T],x\in\mathbb{R}^{d},\ \nu\in\mathscr{P}_{k}.$
+We aim to establish Bismut type formula for the intrinsic derivative of $\mathscr{P}_{k}\ni\mu\mapsto P_{t}f(\mu)$ for bounded measurable functions $f$ on $\mathbb{R}^{d}$, by only assuming that the extrinsic derivative $D^{E}b_{t}(x,\mu)(z)$ of the drift has a half-Dini continuity in $z\in\mathbb{R}^{d}$.
+To this end, we first recall the notions of intrinsic and extrinsic derivatives which go back to [1], see [3] and [21].
 Definition 3.1.
@@
 $\|D^{I}f(\mu)\|_{L^{k^{*}}(\mu)}:=\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|D^{I}_{\phi}f(\mu)\|_{\mathbb{B}}.$
-If moreover
-$\lim_{\|\phi\|_{T_{\mu,k}}\downarrow 0}\frac{\|f(\mu\circ(id+\phi)^{-1})-f(\mu)-D_{\phi}^{I}f(\mu)\|_{\mathbb{B}}}{\|\phi\|_{T_{\mu,k}}}=0,$
-then $f$ is called $L$-differentiable at $\mu$. The function $f$ is called intrinsically differentiable or $L$-differentiable on $\mathscr{P}_{k}$, if it is so at any $\mu\in\mathscr{P}_{k}$.
-Recall that a real function $f$ on a Banach space $\mathbb{B}$ is called Gateaux differentiable, if for any $z\in\mathbb{B}$,
-$v\mapsto\nabla_{v}^{\mathbb{B}}f(z):=\lim_{\varepsilon\downarrow 0}\frac{f(z+\varepsilon v)-f(z)}{\varepsilon}$
-is a well-defined bounded linear functional. In this case we denote
-$\|\nabla^{\mathbb{B}}f(z)\|_{\mathbb{B}^{*}}:=\sup_{\|v\|_{\mathbb{B}}\leq 1}|\nabla_{v}^{\mathbb{B}}f(z)|.$
-Moreover, $f$ is called Fréchet differentiable if it is Gateaux differentiable and
-$\lim_{\|v\|_{\mathbb{B}}\downarrow 0}\frac{|f(z+v)-f(z)-\nabla_{v}^{\mathbb{B}}f(z)|}{\|v\|_{\mathbb{B}}}=0,\ \ z\in\mathbb{B}.$
-It is well-known that a Gateaux differentiable function $f$ is Fréchet differentiable provided $\nabla_{v}^{\mathbb{B}}f(z)$ is continuous in $(v,z)\in\mathbb{B}\times\mathbb{B}$.
-When $\mathbb{B}=\mathbb{R}^{l}$ for some $l\geq 1$, we simply denote $\nabla^{\mathbb{B}}=\nabla$.
+The function $f$ is called intrinsically differentiable on $\mathscr{P}_{k}$, if it is so at any $\mu\in\mathscr{P}_{k}$.
+Next, we recall the (convexity) extrinsic derivative, see e.g. [21, Definition 1.2].
+Definition 3.2.
+A real function $f$ on $\mathscr{P}_{k}$ is called extrinsically differentiable on $\mathscr{P}_{k}$ with derivative $D^{E}f$ if
+$D^{E}f(\mu)(x)=\lim_{\varepsilon\downarrow 0}\frac{f((1-\varepsilon)\mu+\varepsilon\delta_{x})-f(\mu)}{\varepsilon}\in\mathbb{R}$
+exists for all $(x,\mu)\in\mathbb{R}^{d}\times\mathscr{P}_{k}$.
+When $f=(f^{1},f^{2},\cdots,f^{d})$ is an $\mathbb{R}^{d}$-valued function on $\mathscr{P}_{k}$, we denote $D^{E}f=(D^{E}f^{1},D^{E}f^{2},\cdots,D^{E}f^{d})$.
 3.1 Main result
@@
   3. $(B_{2})$
-    There exists $\alpha\in\mathscr{A}$ and $K_{V}\geq 0$ such that
-    $[V]_{\alpha}:=\sup_{x\neq y}\frac{\|V(x)-V(y)\|_{\mathbb{B}}}{\alpha(|x-y|)}\leq 1,\ \ \|V(x)\|_{\mathbb{B}}\leq K_{V}(1+|x|^{k-1}),\ \ x\in\mathbb{R}^{d}.$
+    For any $t\in[0,T],y\in\mathbb{R}^{d}$, $b_{t}^{(1)}(y,\cdot)$ is extrinsically differentiable in $\mathscr{P}_{k}$ with the extrinsic derivative $D^{E}b_{t}^{(1)}(y,\nu)(z)$ being continuous in $(y,\nu,z)\in\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{R}^{d}$.
+Moreover, there exists $\alpha\in\mathscr{A}$ with $\alpha\leq c_{0}(1+|\cdot|^{k-1})$ for some $c_{0}>0$ such that
+    $|D^{E}b_{t}^{(1)}(y,\nu)(z)-D^{E}b_{t}^{(1)}(y,\nu)(\bar{z})|\leq\alpha(|z-\bar{z}|),$
+    $\qquad\quad z,\bar{z}\in\mathbb{R}^{d},t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.$
   4. $(B_{3})$
-    For any $t\in[0,T]$, we have $B_{t}\in C(\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B})$.
-$B_{t}(x,\mu,z)$ is differentiable in $x$, $L$-differentiable in $\mu\in\mathscr{P}_{k}$, and Fréchet differentiable in $z\in\mathbb{B}$ such that $\nabla_{v}^{\mathbb{B}}B_{t}(x,\mu,z)$ is continuous in $(v,z)\in\mathbb{B}\times\mathbb{B}$.
-Next, there exists a constant $K>0$ such that
-    $|B_{t}(0,\delta_{0},0)|\leq K,$
-    $|\nabla B_{t}(\cdot,\mu,z)(x)|+\|D^{L}B_{t}(x,\cdot,z)(\mu)\|_{L^{k^{*}}(\mu)}+\|\nabla^{\mathbb{B}}B_{t}(x,\mu,\cdot)(z)\|_{\mathbb{B}^{*}}\leq K,$
-    $\ \ (t,x,\mu,z)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B}.$
-    Moreover, for any $(t,x,\mu,z)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B}$, there exists a constant $c(t,x,\mu,z)>0$ such that
-    |$|\{D^{L}B_{t}(x,\cdot,z)(\mu)(y)|\leq c(t,x,\mu,z)(1+|y|^{k-1}),\ \ y\in\mathbb{R}^{d}.$
+    For any $t\in[0,T]$, $\nu\in\mathscr{P}_{k}$, $b_{t}^{(1)}(\cdot,\nu)$ is differentiable and there exists a constant $\tilde{K}>0$ such that
+    $|b_{t}^{(1)}(0,\delta_{0})|\leq\tilde{K},\ \ |\nabla b_{t}^{(1)}(y,\nu)|\leq\tilde{K},\ \ (t,y,\nu)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}.$
+As indicated in Introduction that existing results on Bismut type formulas for the intrinsic derivative of $P_{t}f(\mu)$ are established under upper bound conditions on the $L$-derivative of $b_{t}(y,\nu)$ in $\nu$.
+Noting that under a mild condition, the $L$-derivative equals to the gradient of the extrinsic derivative, so the above condition on the $\alpha$-continuity of $D^{E}b_{t}^{(1)}(y,\nu)(z)$ in $z$ is much weaker.
+To see this, we present below a simple example.
 Example 3.1.
-Let $l$ be a positive integer and let $\alpha(s)=s^{\varepsilon}$ for some $\varepsilon\in(0,1\wedge(k-1))$. Then $V:\mathbb{R}^{d}\rightarrow\mathbb{R}^{l}$ with
-$|V(x)-V(y)|\leq\alpha(|x-y|),\ \ x,y\in\mathbb{R}^{d}$
-satisfies ($B_{2}$).
-Moreover, let $g\in C^{1}(\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{l};\mathbb{R}^{d})$ with bounded gradient.
-Then $B_{t}(x,\mu,z)=\int_{\mathbb{R}^{d}}g(x,y,z)\mu(\text{\rm{d}}y)$ satisfies ($B_{3}$).
+Let $\alpha(s)=s^{\varepsilon}$ for some $\varepsilon\in(0,1\wedge(k-1))$. Let $g:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ satisfy
+$|g(y,z)-g(y,\bar{z})|\leq\alpha(|z-\bar{z}|),\ \ |\nabla g(\cdot,z)|\leq K,\ \ y,z,\bar{z}\in\mathbb{R}^{d}$
+for some $K\geq 0$. Let $b_{t}^{(1)}(y,\nu)=\int_{\mathbb{R}^{d}}g(y,z)\nu(\text{\rm{d}}z)$. By Definition 3.2, it holds
+$D^{E}b_{t}^{(1)}(y,\nu)(z)=g(y,z)-\int_{\mathbb{R}^{d}}g(y,z)\nu(\text{\rm{d}}z),\ \ y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k},z\in\mathbb{R}^{d}.$
+However, by Definition 3.1, $b_{t}^{(1)}(y,\nu)$ is not intrinsically differentiable in $\nu$.
+In fact, since $g(y,z)$ is not differentiable in $z$, for any $y\in\mathbb{R}^{d}$, $\nu\in\mathscr{P}_{k}$, $\phi\in L^{k}(\mathbb{R}^{d}\rightarrow\mathbb{R}^{d};\nu)$, the limit
+$\lim_{r\downarrow 0}\frac{\int_{\mathbb{R}^{d}}g(y,z+r\phi(z))\nu(\text{\rm{d}}z)-\int_{\mathbb{R}^{d}}g(y,z)\nu(\text{\rm{d}}z)}{r}$
+does not exist. Moreover, it holds
+$|D^{E}b_{t}^{(1)}(y,\nu)(z)-D^{E}b_{t}^{(1)}(\bar{y},\bar{\nu})(\bar{z})|$
+$=|g(y,z)-g(\bar{y},\bar{z})|+\left|\int_{\mathbb{R}^{d}}g(y,z)\nu(\text{\rm{d}}z)-\int_{\mathbb{R}^{d}}g(\bar{y},z)\bar{\nu}(\text{\rm{d}}z)\right|$
+$\leq\alpha(|z-\bar{z}|)+2K|y-\bar{y}|+\mathbb{W}_{\alpha}(\nu,\bar{\nu}),\ \ y,\bar{y}\in\mathbb{R}^{d},\nu,\bar{\nu}\in\mathscr{P}_{k},z,\bar{z}\in\mathbb{R}^{d}.$
+Note that Jensen’s inequality implies that
+$\mathbb{W}_{\alpha}(\mu,\nu)\leq\inf_{\pi\in\mathscr{C}(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\alpha(|x-y|)\pi(\text{\rm{d}}x,\text{\rm{d}}y)\leq\alpha(\mathbb{W}_{1}(\mu,\nu))\leq\alpha(\mathbb{W}_{k}(\mu,\nu)),\ \ \mu,\nu\in\mathscr{P}_{k}.$
+So, $D^{E}b_{t}^{(1)}(y,\nu)(z)$ is continuous in $(y,\nu,z)\in\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{R}^{d}$. Finally, by the dominated convergence theorem, we have
+$\nabla b_{t}^{(1)}(\cdot,\nu)=\int_{\mathbb{R}^{d}}\nabla g(\cdot,z)\nu(\text{\rm{d}}z),\ \ \nu\in\mathscr{P}_{k}.$
+Therefore, $b^{(1)}$ satisfies $(B_{2})$-$(B_{3})$.
 Since (B) implies (A), as explained before that under this assumption (1.1) is well-posed for distributions in $\mathscr{P}_{k}.$
 For $\mu\in\mathscr{P}_{k}$, consider the decoupled SDE
-(3.2) $\begin{split}&\text{\rm{d}}X_{t}^{x,\mu}=\big{\{}b_{t}^{(0)}(X_{t}^{x,\mu})+B_{t}(X_{t}^{x,\mu},P_{t}^{*}\mu,P_{t}V(\mu))\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\mu})\text{\rm{d}}W_{t},\\ &\qquad\qquad\qquad X_{0}^{x,\mu}=x,\ \ t\in[0,T].\end{split}$
+(3.2) $\begin{split}&\text{\rm{d}}X_{t}^{x,\mu}=\big{\{}b_{t}^{(0)}(X_{t}^{x,\mu})+b_{t}^{(1)}(X_{t}^{x,\mu},P_{t}^{*}\mu)\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\mu})\text{\rm{d}}W_{t},\\ &\qquad\qquad\qquad X_{0}^{x,\mu}=x,\ \ t\in[0,T].\end{split}$
+Let
+$\mathscr{B}_{k,b}(\mathbb{R}^{d}):=\bigg{\{}f:\ \frac{f}{1+|\cdot|^{k}}\in\mathscr{B}_{b}(\mathbb{R}^{d})\bigg{\}}.$
+We first give a lemma on Bismut formula of $P_{t}^{\mu}f$ for $f\in\mathscr{B}_{k,b}(\mathbb{R}^{d})$.
+Lemma 3.2.
+Let $\sigma$ and $b$ satisfy (B). Then for any $v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}$, the limit
+$\nabla_{v}X_{t}^{x,\gamma}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{x+\varepsilon v,\gamma}-X_{t}^{x,\gamma}}{\varepsilon},\ \ \ \ t\in[0,T]$
+exists in $L^{p}(\Omega\rightarrow C([0,T];\mathbb{R}^{d});\mathbb{P})$ for any $p\geq 1$, and there exists a constant $c_{p}>0$ such that
+(3.3) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{v}X_{t}^{x,\gamma}|^{p}\Big{]}\leq c_{p}|v|^{p},\ \ v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}.$
+Moreover, the Bismut formula for $P^{\gamma}_{t}$ holds:
+(3.4) $\begin{split}&\nabla_{v}P^{\gamma}_{t}f(x)=\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right],\\ &\qquad\zeta_{s}:=\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1},\ \ \ f\in\mathscr{B}_{k,b}(\mathbb{R}^{d}),\ x,v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},t\in(0,T].\end{split}$
+Proof.
+By [30, Theorem 2.1] for $\beta_{s}=\frac{s}{t}$, (B) implies (3.3) and (3.4) for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$. To deduce (3.4) for any $f\in\mathscr{B}_{b,k}(\mathbb{R}^{d}),$ let
+$f_{n}:=[(-n)\lor f]\land n,\ \ n\geq 1.$
+By (2.8), (3.3) and the boundedness of $\zeta_{s}$, we find constants $c_{0},c_{1}(\gamma)>0$ such that
+(3.5) $\begin{split}&\mathbb{E}\bigg{[}\big{(}1+|X_{t}^{x+rv,\gamma}|^{k}\big{)}\bigg{|}\int_{0}^{t}\big{\langle}\zeta_{s}(X^{x+rv,\gamma}_{s})\nabla_{v}X^{x+rv,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\bigg{|}\bigg{]}\\ &\leq c_{0}\sqrt{t}\Big{(}\mathbb{E}\big{[}2+|X_{t}^{x+rv,\gamma}|^{2k}\big{]}\Big{)}^{\frac{1}{2}}\Big{(}\mathbb{E}\big{[}\sup_{s\in[0,T]}\big{|}\nabla_{v}X^{x+rv,\gamma}_{s}\big{|}^{2}\big{]}\Big{)}^{\frac{1}{2}}\\ &\leq c_{1}(\gamma)\sqrt{t}|v|\big{(}1+|x|^{k}+|v|^{k}\big{)},\ \ t\in(0,T],x,v\in\mathbb{R}^{d},r\in[0,1].\end{split}$
+By (3.4) for $f_{n}$ in place of $f$, we obtain
+$\frac{P^{\gamma}_{t}f_{n}(x+\varepsilon v)-P^{\gamma}_{t}f_{n}(x)}{\varepsilon}$
+$=\frac{1}{\varepsilon}\int_{0}^{\varepsilon}\mathbb{E}\left[f_{n}(X^{x+rv,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x+rv,\gamma}_{s})\nabla_{v}X^{x+rv,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\text{\rm{d}}r.$
+Since $f\in\mathscr{B}_{k,b}(\mathbb{R}^{d})$, by (2.8), (3.3) and (3.5), we may apply the dominated convergence theorem such that the above formula with $n\rightarrow\infty$ implies
+(3.6) $\begin{split}&\frac{P^{\gamma}_{t}f(x+\varepsilon v)-P^{\gamma}_{t}f(x)}{\varepsilon}\\ &=\frac{1}{\varepsilon}\int_{0}^{\varepsilon}\mathbb{E}\left[f(X^{x+rv,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x+rv,\gamma}_{s})\nabla_{v}X^{x+rv,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\text{\rm{d}}r,\\ &\qquad f\in\mathscr{B}_{k,b}(\mathbb{R}^{d}),\varepsilon>0,\ x,v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},t\in(0,T].\end{split}$
+Note that (3.4) for $f_{n}$ in place of $f$ yields
+$\lim_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\bigg{|}\frac{P^{\gamma}_{t}f_{n}(x+\varepsilon v)-P^{\gamma}_{t}f_{n}(x)}{\varepsilon}-\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\bigg{|}$
+$=\lim_{n\rightarrow\infty}\bigg{|}\mathbb{E}\left[f_{n}(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]$
+$\qquad\qquad\quad-\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\bigg{|}=0,$
+where the last step follows from the dominated convergence theorem due to $f\in\mathscr{B}_{k,b}(\mathbb{R}^{d})$ and (3.5). This together with (3.6) for $f-f_{n}$ in place of $f$ implies
+(3.7) $\begin{split}&\limsup_{\varepsilon\rightarrow 0}\bigg{|}\frac{P^{\gamma}_{t}f(x+\varepsilon v)-P^{\gamma}_{t}f(x)}{\varepsilon}-\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\bigg{|}\\ &\leq\lim_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\bigg{|}\frac{P^{\gamma}_{t}(f-f_{n})(x+\varepsilon v)-P^{\gamma}_{t}(f-f_{n})(x)}{\varepsilon}\bigg{|}\\ &+\lim_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\bigg{|}\frac{P^{\gamma}_{t}f_{n}(x+\varepsilon v)-P^{\gamma}_{t}f_{n}(x)}{\varepsilon}\\ &\qquad\qquad\qquad\qquad\quad-\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\bigg{|}\\ &\leq\lim_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\frac{1}{t\varepsilon}\int_{0}^{\varepsilon}\mathbb{E}\bigg{|}(f_{n}-f)(X_{t}^{x+rv,\gamma})\\ &\qquad\qquad\qquad\qquad\qquad\quad\times\int_{0}^{t}\big{\langle}\zeta_{s}(X^{x+rv,\gamma}_{s})\nabla_{v}X^{x+rv,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\bigg{|}\text{\rm{d}}r.\end{split}$
+Since $f\in\mathscr{B}_{k,b}(\mathbb{R}^{d})$ implies
+$|(f_{n}-f)(x)|\leq c(1+|x|^{k})1_{\{c(1+|x|^{k})\geq n\}},\ \ n\geq 1$
+for some constant $c>0$, by the same reason leading to (3.5), we find constants $\tilde{c},c_{2}(\gamma)>0$ such that
+$\sup_{r\in[0,1]}\mathbb{E}\bigg{|}(f_{n}-f)(X_{t}^{x+rv,\gamma})\int_{0}^{t}\big{\langle}\zeta_{s}(X^{x+rv,\gamma}_{s})\nabla_{v}X^{x+rv,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\bigg{|}$
+$\leq\tilde{c}\sqrt{t}|v|\sup_{r\in[0,1]}\Big{(}\mathbb{E}\big{[}1+|X_{t}^{x+rv,\gamma}|^{4k}\big{]}\Big{)}^{\frac{1}{2}}n^{-1}$
+$\leq c_{2}(\gamma)\sqrt{t}|v|\big{(}1+|x|^{2k}+|v|^{2k}\big{)}n^{-1}.$
+Therefore, (3.4) follows form (3.7). ∎
 To state the Bismut formula for $P_{t}f$, we introduce the quantity $I_{t}^{f}$: for fixed $t\in(0,T]$, let
-(3.3) $\begin{split}I_{t}^{f}(\mu,\phi)&:=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),\\ \zeta_{s}&:=\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1},\ \ \ s\in[0,t],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}.\end{split}$
-By $(B_{1})$ and (2.14), we find a constant $c>0$ such that
-(3.4) $|I_{t}^{f}(\mu,\phi)|\leq\frac{c}{\sqrt{t}}\big{(}P_{t}|f|^{k^{*}}(\mu)\big{)}^{\frac{1}{k^{*}}}\|\phi\|_{L^{k}(\mu)},\ \ \mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}.$
+(3.8) $\begin{split}I_{t}^{f}(\mu,\phi)&:=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),\\ &s\in[0,t],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k},f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}).\end{split}$
+By (B) and (3.3), we find a constant $c>0$ such that
+(3.9) $|I_{t}^{f}(\mu,\phi)|\leq\frac{c}{\sqrt{t}}\big{(}P_{t}|f|^{k^{*}}(\mu)\big{)}^{\frac{1}{k^{*}}}\|\phi\|_{L^{k}(\mu)},\ \ \mu\in\mathscr{P}_{k},\phi\in T_{\mu,k},f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}).$
 Next, let $X_{0}^{\mu}$ be $\mathscr{F}_{0}$-measurable such that $\mathscr{L}_{X_{0}^{\mu}}=\mu$, and let $X_{t}^{\mu}$ solve (1.1) with initial value $X_{0}^{\mu}$.
@@
 $X_{t}^{\mu}=X_{t}^{\mu_{0}},\ \ P_{t}^{*}\mu_{\varepsilon}=\mathscr{L}_{X_{t}^{\mu_{\varepsilon}}},\ \ t\in[0,T],\varepsilon\geq 0.$
+Now, we present the main result of this part.
 Theorem 3.3.
@@
   1. $(1)$
+    For any $t\in(0,T]$, $y\in\mathbb{R}^{d}$, $\nu\in\mathscr{P}_{k}$, $P_{t}[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)](\mu)$ is intrinsically differentiable on $\mu\in\mathscr{P}_{k}$, and there exists a constant $c>0$ such that
+    $\sup_{(y,\nu)\in\mathbb{R}^{d}\times\mathscr{P}_{k}}\|D^{I}P_{t}[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)](\mu)\|_{L^{\frac{k}{k-1}}(\mu)}\leq\frac{c\,\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\mu\in\mathscr{P}_{k}.$
+  2. $(2)$
+    For any $t\in(0,T]$ and $f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d})$, $P_{t}f$ is intrinsically differentiable on $\mathscr{P}_{k}$. Moreover, for any $\mu\in\mathscr{P}_{k}$ and $\phi\in T_{\mu,k}$,
+    (3.10) $\begin{split}D_{\phi}^{I}P_{t}f(\mu)&=I_{t}^{f}(\mu,\phi)\\ &+\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})N_{s}(\mu,\phi),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),\\ &N_{s}(\mu,\phi):=\big{\{}D_{\phi}^{I}P_{s}[D^{E}b_{s}^{(1)}(y,\nu)(\cdot)](\mu)\big{\}}|_{y=X_{s}^{x,\mu},\nu=P_{s}^{\ast}\mu},\end{split}$
 […]
-  2. $(2)$
-    The limit
-    (3.5) $\nabla_{\phi}X_{t}^{\mu}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{\mu_{\varepsilon}}-X_{t}^{\mu}}{\varepsilon},\ \ \ \ t\in[0,T]$
-    exists in $L^{k}(\Omega\rightarrow C([0,T],\mathbb{R}^{d}),\mathbb{P})$, and there exists a constant $c>0$ such that
-    $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{\phi}X_{t}^{\mu}|^{k}\Big{]}\leq c\|\phi\|_{L^{k}(\mu)}^{k},\ \ \mu\in\mathscr{P}_{k}.$
-  3. $(3)$
-    For any $t\in(0,T]$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, $P_{t}f$ is intrinsically differentiable on $\mathscr{P}_{k}$. Moreover, for any $\mu\in\mathscr{P}_{k}$ and $\phi\in T_{\mu,k}$,
-    (3.6) $\begin{split}&D_{\phi}^{I}P_{t}f(\mu)=I_{t}^{f}(\mu,\phi)+\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\\ &N_{s}:=\big{\{}\nabla^{\mathbb{B}}_{D_{\phi}^{I}P_{s}V(\mu)}B_{s}(X_{s}^{\mu},P_{s}^{\ast}\mu,\cdot)\big{\}}(P_{s}V(\mu)),\\ &\tilde{N}_{s}:=\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{\ast}\mu)(X_{s}^{\mu}),\nabla_{\phi}X_{s}^{\mu}\big{\rangle}\big{]}|_{y=X_{s}^{\mu}},\end{split}$
-[…]
-By (3.4)-(3.6) and ($B_{3}$), we find a constant $c>0$ such that
-$\|D^{I}P_{t}f(\mu)\|_{L^{k^{*}}(\mu)}\leq c\frac{\{P_{t}|f|^{k^{*}}(\mu)\}^{\frac{1}{k^{*}}}}{\sqrt{t}},\ \ t\in(0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\mu\in\mathscr{P}_{k}.$
-Remark 3.3.
-Note that if $\nabla V$ is continuous and bounded, then the chain rule (see for instance [19, Proposition 3.1]) implies
-(3.7) $D^{I}_{\phi}P_{t}V(\mu)=\mathbb{E}\nabla_{\nabla_{\phi}X_{t}^{\mu}}V(X_{t}^{\mu}).$
-However, in the present case, $V$ only satisfies $(B_{2})$ so that (3.7) does not hold and there is essential difficulty in the study of $D^{I}_{\phi}P_{t}V$.
-To explain the main idea in the proof of Theorem 3.2, we first figure out a sketch. By the definition of the intrinsic derivative, we intend to calculate
+By (3.9)-(3.10), we find a constant $c>0$ such that
+$\|D^{I}P_{t}f(\mu)\|_{L^{k^{*}}(\mu)}\leq c\frac{\{P_{t}|f|^{k^{*}}(\mu)\}^{\frac{1}{k^{*}}}}{\sqrt{t}},\ \ t\in(0,T],f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}),\mu\in\mathscr{P}_{k}.$
+To explain the main idea in the proof of Theorem 3.3, we first figure out a sketch.
+By the definition of the intrinsic derivative, we intend to calculate for any $f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d})$,
 (3.11) $D_{\phi}^{I}P_{t}f(\mu):=\lim_{\varepsilon\downarrow 0}\frac{P_{t}f(\mu_{\varepsilon})-P_{t}f(\mu)}{\varepsilon}=\lim_{\varepsilon\downarrow 0}\frac{\mathbb{E}[f(X_{t}^{\mu_{\varepsilon}})-f(X_{t}^{\mu})]}{\varepsilon}.$
-To this end, for any $\gamma\in\mathscr{P}_{k},$ consider the decoupled SDE (3.2) for $\gamma$ replacing $\mu$, i.e.
-(3.9) $\begin{split}\text{\rm{d}}X_{t}^{x,\gamma}=\big{\{}b_{t}^{(0)}(X_{t}^{x,\gamma})+&B_{t}\big{(}X_{t}^{x,\gamma},P_{t}^{\ast}\gamma,P_{t}V(\gamma)\big{)}\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\gamma})\text{\rm{d}}W_{t},\\ &t\in[0,T],X_{0}^{x,\gamma}=x\in\mathbb{R}^{d}.\end{split}$
-Let
-$P_{t}^{\gamma}f(x):=\mathbb{E}[f(X_{t}^{x,\gamma})],\ \ \ x\in\mathbb{R}^{d},$
-$P_{t}^{\gamma}f(\tilde{\gamma}):=\int_{\mathbb{R}^{d}}P_{t}^{\gamma}f\text{\rm{d}}\tilde{\gamma},\ \ t\geq 0,f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
-For $\varepsilon\geq 0$, let $X_{t}^{\mu_{\varepsilon},\gamma}$ be the solution of (3.9) with initial value $X_{0}^{\varepsilon}$, i.e,
-$\begin{split}\text{\rm{d}}X_{t}^{\mu_{\varepsilon},\gamma}=\big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\gamma})+&B_{t}\big{(}X_{t}^{\mu_{\varepsilon},\gamma},P_{t}^{\ast}\gamma,P_{t}V(\gamma)\big{)}\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\gamma})\text{\rm{d}}W_{t},\\ &t\in[0,T],\ X_{0}^{\mu_{\varepsilon},\gamma}=X_{0}^{\varepsilon}.\end{split}$
+To this end, for any $\mu\in\mathscr{P}_{k},x\in\mathbb{R}^{d}$, recall that $X_{t}^{x,\mu}$ solves the decoupled SDE (3.2), and
+$P_{t}^{\mu}f(x)=\mathbb{E}[f(X_{t}^{x,\mu})],\ \ \ x\in\mathbb{R}^{d}.$
+Define
+$P_{t}^{\mu}f(\tilde{\mu}):=\int_{\mathbb{R}^{d}}P_{t}^{\mu}f\text{\rm{d}}\tilde{\mu},\ \ t\geq 0,f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}),\mu,\tilde{\mu}\in\mathscr{P}_{k}.$
+For $\varepsilon\geq 0$, let $X_{t}^{\mu_{\varepsilon},\mu}$ be the solution of (3.2) with initial value $X_{0}^{\varepsilon}$, i.e,
+$\begin{split}\text{\rm{d}}X_{t}^{\mu_{\varepsilon},\mu}=\big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\mu})+&b_{t}^{(1)}\big{(}X_{t}^{\mu_{\varepsilon},\mu},P_{t}^{\ast}\mu\big{)}\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\mu})\text{\rm{d}}W_{t},\\ &t\in[0,T],\ X_{0}^{\mu_{\varepsilon},\mu}=X_{0}^{\varepsilon}.\end{split}$
 Then $X_{t}^{\mu_{\varepsilon},\mu_{\varepsilon}}$ solves (1.1) with initial value $X_{0}^{\varepsilon}$, so that
-$P_{t}f(\mu_{\varepsilon})=P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})=\mathbb{E}[f(X_{t}^{\mu_{\varepsilon},\mu_{\varepsilon}})],\ \ \varepsilon\geq 0,t\in[0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
+$P_{t}f(\mu_{\varepsilon})=P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})=\mathbb{E}[f(X_{t}^{\mu_{\varepsilon},\mu_{\varepsilon}})],\ \ \varepsilon\geq 0,t\in[0,T],f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}).$
 Noting that $\mu_{0}=\mu$, (3.11) reduces to
@@
 $J_{1}f(t,\varepsilon):=\frac{P_{t}^{\mu}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu)}{\varepsilon},\ \ J_{2}f(t,\varepsilon):=\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}.$
-By Lemma 2.3(1), for any $t\in(0,T],\varepsilon\geq 0$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, we have
-(3.11) $\begin{split}&\frac{\text{\rm{d}}}{\text{\rm{d}}\varepsilon}P_{t}^{\mu}f(\mu_{\varepsilon}):=\lim_{r\downarrow 0}\frac{P_{t}^{\mu}f(\mu_{\varepsilon+r})-P_{t}^{\mu}f(\mu_{\varepsilon})}{r}\\ &=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x+\varepsilon\phi(x),\mu})\frac{1}{t}\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+\varepsilon\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x)\\ &=:I_{t}^{f}(\varepsilon,\mu,\phi),\ \ t\in(0,T].\end{split}$
+By Lemma 3.2, for any $t\in(0,T],\varepsilon\geq 0$ and $f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d})$, we have
+$\begin{split}&\frac{\text{\rm{d}}}{\text{\rm{d}}\varepsilon}P_{t}^{\mu}f(\mu_{\varepsilon}):=\lim_{r\downarrow 0}\frac{P_{t}^{\mu}f(\mu_{\varepsilon+r})-P_{t}^{\mu}f(\mu_{\varepsilon})}{r}\\ &=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x+\varepsilon\phi(x),\mu})\frac{1}{t}\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+\varepsilon\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).\end{split}$
 In particular,
@@
 Consequently, it remains to prove
-$\lim_{\varepsilon\rightarrow 0}J_{2}f(t,\varepsilon)=\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}$
-for $N_{s},\tilde{N}_{s}$ defined in (3.6), which involves in $D^{I}P_{t}V(\mu)$ and $\nabla_{\phi}X_{t}^{\mu}$ defined in (3.5).
-Therefore, we will first study $D^{I}P_{t}V(\mu)$ and $\nabla_{\phi}X_{t}^{\mu}$.
-To this end, let $u_{t}=u_{t}^{\lambda_{0},\mu},\Theta_{t}=\Theta_{t}^{\lambda_{0},\mu}$ be defined in (2.7) and (2.10) for large enough $\lambda_{0}>0$ such that (2.9) holds for $\lambda=\lambda_{0}$, and let
-$\mathscr{C}_{k}:=\left\{\tilde{v}\in L^{k}\big{(}\Omega\rightarrow C([0,T];\mathbb{R}^{d}),\mathbb{P}\big{)}\ \text{is\ }\mathscr{F}_{t}\text{-adapted,\ }\tilde{v}_{0}=\nabla_{\phi(X_{0}^{\mu})}\Theta_{0}(X_{0}^{\mu})\right\},$
-which is complete under the metric
-$\rho_{\delta}(\tilde{v}^{1},\tilde{v}^{2}):=\left(\mathbb{E}\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta kt}|\tilde{v}_{t}^{1}-\tilde{v}^{2}_{t}|^{k}\right)^{\frac{1}{k}},\ \ \tilde{v}^{1},\tilde{v}^{2}\in\mathscr{C}_{k},\delta>0.$
-For any $s\in[0,T]$, $\eta\in L^{k}(\Omega\rightarrow\mathbb{R}^{d},\mathbb{P})$, define
-$\psi_{s}(v):=\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{*}\mu)(X_{s}^{\mu}),(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\eta\big{\rangle}\big{]}|_{y=X_{s}^{\mu}}.$
-Recall that $\tilde{\alpha}$ is defined in (2.35). For any $(v,\tilde{v})\in C([0,T];\mathbb{B})\times\mathscr{C}_{k}$, we denote
-(3.13) $\begin{split}&N_{s}(v):=\nabla_{v_{s}}^{\mathbb{B}}B_{s}(X_{s}^{\mu},P_{s}^{*}\mu,\cdot)(P_{s}V(\mu)),\ \ \tilde{N}_{s}(\tilde{v}):=\psi_{s}(\tilde{v}_{s}),\end{split}$
-and write
-$\begin{split}&\tilde{I}_{t}^{V}(\mu,v,\tilde{v})=\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})+\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v}),\\ &\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v}):=\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})}{\{\tilde{\alpha}(s^{\frac{1}{2}})s\}^{\frac{1}{2}}}\Big{\langle}\zeta_{s}(X_{s}^{\mu})N_{s}(v),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\\ &\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v}):=\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\tilde{N}_{s}(\tilde{v}),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\ \ t\in[0,T].\end{split}$
-Consider the following equation for $(v,\tilde{v})\in C([0,T];\mathbb{B})\times\mathscr{C}_{k}$:
-(3.14) $\begin{split}&\text{\rm{d}}\tilde{v}_{t}=\Big{[}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}B_{t}(\cdot,P_{t}^{*}\mu,P_{t}V(\mu))+\lambda_{0}u_{t}\big{\}}(X_{t}^{\mu})\Big{]}\text{\rm{d}}t\\ &+\nabla_{\psi_{t}(\tilde{v}_{t})}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t+{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\nabla_{N_{t}(v)}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t\\ &+\Big{[}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X^{\mu}_{t})\Big{]}\,\text{\rm{d}}W_{t},\ \ \tilde{v}_{0}=\nabla_{\phi(X_{0}^{\mu})}\Theta_{0}(X_{0}^{\mu}),\\ &v_{t}={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{V}(\mu,\phi)+\tilde{I}_{t}^{V}(\mu,v,\tilde{v})\big{\}},\ \ t\in[0,T].\end{split}$
-In the following two subsections, we prove the well-posedness of (3.14) and the formulas
-(3.15) $\begin{split}&D_{\phi}^{I}P_{t}V(\mu)={\frac{\alpha(t^{\frac{1}{2}})}{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}}\,v_{t},\ \ t\in(0,T],\\ &\nabla_{\phi}X_{t}^{\mu}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{\mu_{\varepsilon}}-X_{t}^{\mu}}{\varepsilon}=(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t},\ \ \ t\in[0,T].\end{split}$
+$\lim_{\varepsilon\rightarrow 0}J_{2}f(t,\varepsilon)=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})N_{s}(\mu,\phi),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x)$
+for $N_{s}(\mu,\phi)$ defined in (3.10), which involves in $D_{\phi}^{I}\{P_{s}[D^{E}b_{s}^{(1)}(y,\nu)(\cdot)](\mu)\}$.
+Therefore, we will first study $D_{\phi}^{I}\{P_{s}[D^{E}b_{s}^{(1)}(y,\nu)(\cdot)](\mu)\}$.
+Recall that $\tilde{\alpha}$ is defined in (2.40).
+For any $V\in\mathscr{B}_{b}([0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k};\mathbb{R}^{d})$, the set of bounded and measurable $\mathbb{R}^{d}$-valued functions on $[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}$ and $t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}$, we write
+(3.14) $\begin{split}\tilde{I}_{t}^{V}(y,\nu)&:=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}D^{E}b_{t}^{(1)}(y,\nu)(X_{t}^{x,\mu})\\ &\qquad\qquad\quad\times\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})}{\{\tilde{\alpha}(s^{\frac{1}{2}})s\}^{\frac{1}{2}}}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})V_{s}(X_{s}^{x,\mu},P_{s}^{\ast}\mu),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).\end{split}$
+By $(B_{2})$, for any $t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}$, we have
+$|D^{E}b_{t}^{(1)}(y,\nu)(\cdot)|$ $\leq\alpha(|\cdot|)+|D^{E}b_{t}^{(1)}(y,\nu)(0)|$
+$\leq c_{0}(1+|\cdot|^{k-1})+|D^{E}b_{t}^{(1)}(y,\nu)(0)|.$
+So, $I_{t}^{f}(\mu,\phi)$ for $f=D^{E}b_{t}^{(1)}(y,\nu)(\cdot)$ in (3.8) is well-defined, and we denote
+(3.15) $\begin{split}I_{t}^{\mu,\phi}(y,\nu)&:=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}D^{E}b_{t}^{(1)}(y,\nu)(X_{t}^{x,\mu})\\ &\qquad\quad\times\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).\end{split}$
+Consider the following equation for $V\in\mathscr{B}_{b}([0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k};\mathbb{R}^{d})$:
+(3.16) $\begin{split}&V_{t}(y,\nu)={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{\mu,\phi}(y,\nu)+\tilde{I}_{t}^{V}(y,\nu)\big{\}},\ \ t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.\end{split}$
+If this equation has a unique solution, we denote it by $V_{t}(y,\nu)=v_{t}^{\mu,\phi}(y,\nu)$ for $(t,y,\nu)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}$, to emphasize the dependence on $\mu$ and $\phi$.
+In the following two subsections, we prove the well-posedness of (3.16) and establish the formula
+(3.17) $\begin{split}&D_{\phi}^{I}\{P_{t}[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)]\}(\mu)={\frac{\alpha(t^{\frac{1}{2}})}{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}}\,v_{t}^{\mu,\phi}(y,\nu),\ \ t\in(0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.\end{split}$
 3.2 Well-posedness of (3.16)
@@
 […]
-(3.16) $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|v_{t}(\mu,\phi)\|_{\mathbb{B}}\leq{c\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ \mu\in\mathscr{P}_{k},t\in[0,T],$
-and for any $j\geq 1$, there exists a constant $\tilde{c}>0$ such that
-(3.17) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\tilde{v}_{t}(\mu,\phi)|^{j}|\mathscr{F}_{0}\Big{]}\leq\tilde{c}|\phi(X_{0}^{\mu})|^{j}+{\tilde{c}\|\phi\|_{L^{k}(\mu)}^{j}},\ \ \mu\in\mathscr{P}_{k},t\in[0,T].$
+(3.18) $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\sup_{y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}}|v_{t}^{\mu,\phi}(y,\nu)|\leq{c\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ \mu\in\mathscr{P}_{k},t\in[0,T].$
 Proof.
-By Jensen’s inequality, for estimate (3.17) we only need to consider $j\geq k.$
-As in the proof of [29, Lemma 4.1], there exists a constant $\theta>1$ and $0<f\in\mathscr{B}([0,T]\times\mathbb{R}^{d})$ such that
-(3.18) [$\big{\|}\nabla\big{[}\{\nabla\Theta\}\sigma\big{]}\big{\|}\leq f,\ \ \ \mathbb{E}\int_{0}^{T}f_{s}(X_{s}^{\mu})^{2\theta}\text{\rm{d}}s<\infty.$
-For any $n\geq 0$, define
-$\tau_{n}:=T\wedge\inf\left\{t\geq 0:\int_{0}^{t}f_{s}(X_{s}^{\mu})^{2\theta}\text{\rm{d}}s\geq n\right\},$
-$\mathscr{C}_{k,n}:=\big{\{}(z_{\cdot\wedge\tau_{n}}),\ \ z\in\mathscr{C}_{k}\big{\}}.$
-We have $\lim_{n\rightarrow\infty}\tau_{n}=T$. Let
-$\mathscr{V}_{0}:=\big{\{}v\in C([0,T];\mathbb{B}):\ v_{0}=0\big{\}}.$
-Recall that $\tilde{\alpha}$ is defined in (2.35). By (3.18), $\|\nabla B\|_{\infty}<\infty$ due to (B), and
-(3.19) $\int_{0}^{t}\frac{\alpha(rs^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(rs^{\frac{1}{2}})}\text{\rm{d}}s=2\int_{0}^{rt^{\frac{1}{2}}}\frac{\alpha(s)^{2}}{s\tilde{\alpha}(s)}\text{\rm{d}}s=4\int_{0}^{rt^{\frac{1}{2}}}\tilde{\alpha}^{\prime}(s)\text{\rm{d}}s=4\tilde{\alpha}(rt^{\frac{1}{2}}),\ \ r\geq 0,$
-we see that for any $(v,\tilde{v})\in\mathscr{V}_{0}\times\mathscr{C}_{k},$
-$H_{s}(v,\tilde{v})$ $:=\tilde{v}_{0}+\int_{0}^{s}\nabla_{\psi_{t}(\tilde{v}_{t})}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t$
-$+\int_{0}^{s}\Big{\{}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\{B_{t}(\cdot,P_{t}^{*}\mu,P_{t}V(\mu))+\lambda_{0}u_{t}\}\Big{\}}(X_{t}^{\mu})\text{\rm{d}}t$
-$+\int_{0}^{s}{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\nabla_{N_{t}(v)}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t$
-$+\int_{0}^{s}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X^{\mu}_{t})\,\text{\rm{d}}W_{t},\ \ s\in[0,T]$
-is a continuous adapted process on $\mathbb{R}^{d}$ such that
-$H_{\cdot\land\tau_{n}}:\mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{C}_{k,n},\ \ n\geq 1.$
-On the other hand, let
-(3.20) $\begin{split}&\bar{H}_{t}(v,\tilde{v}):={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{V}(\mu,\phi)+\tilde{I}_{t}^{V}(\mu,v,\tilde{v})\big{\}}.\end{split}$
-We claim that it suffices to prove
+Let
+$\mathscr{V}_{0}:=\big{\{}V\in\mathscr{B}_{b}([0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k};\mathbb{R}^{d}):\ V_{0}=0\big{\}},$
+which is a Banach space under the uniform norm. For $V\in\mathscr{V}_{0},$ let
+$\|V_{t}\|_{\infty}=\sup_{y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}}|V_{t}(y,\nu)|,\ \ t\in[0,T]$
+and for any $t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}$, let
+(3.19) $\begin{split}&\{H(V)\}_{t}(y,\nu):={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{\mu,\phi}(y,\nu)+\tilde{I}_{t}^{V}(y,\nu)\big{\}}.\end{split}$
+Then it suffices to prove
   1. (i)
-    for each $n\geq 1$, the map
-    $\Phi^{(n)}:=(\bar{H},H_{\cdot\land\tau_{n}}):\ \mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{V}_{0}\times\mathscr{C}_{k,n}$
-    is well-defined and has a unique fixed point $(v^{(n)},{\tilde{v}}^{(n)})$.
+[…]
   2. (ii)
     There exists a constant $c>0$ such that
-    $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|v_{t}^{(n)}\|_{\mathbb{B}}\leq c{\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ (t,\mu)\in[0,T]\times\mathscr{P}_{k},n\geq 1,$
-    and for any $j\geq k$, there exists a constant $\tilde{c}>0$ such that
-    $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|{\tilde{v}}_{t}^{(n)}|^{j}|\mathscr{F}_{0}\Big{]}\leq\tilde{c}|\phi(X_{0}^{\mu})|^{j}+\tilde{c}{\|\phi\|_{L^{k}(\mu)}^{j}},\ \ (t,\mu)\in[0,T]\times\mathscr{P}_{k},n\geq 1.$
-Indeed, by (i), the fixed points are consistent:
-$\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{t\land\tau_{n}}=\big{(}v^{(n+i)},{\tilde{v}}^{(n+i)}\big{)}_{t\land\tau_{n}},\ \ n,i\geq 1,t\in[0,T],$
-so that $\mathbb{P}$-a.s.
-$\big{(}v(\mu,\phi),\tilde{v}(\mu,\phi)\big{)}:=\lim_{n\rightarrow\infty}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{\cdot\land\tau_{n}}=\sum_{n=1}^{\infty}1_{[\tau_{n-1},\tau_{n})}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{\cdot\land\tau_{n}}$
-exists in $C([0,T];\mathbb{B}\times\mathbb{R}^{d})$, and by (ii) the estimates (3.16) and (3.17) hold. Combining this with
-(3.21) $\Phi^{(n)}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}:=\big{(}\bar{H},H_{\cdot\land\tau_{n}}\big{)}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}=\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)},\ \ n\geq 1,$
-we see that $(v(\mu,\phi),\tilde{v}(\mu,\phi))$ solves (3.14), and the uniqueness follows from that of the fixed point for $\Phi^{(n)},n\geq 1.$ Next, we will prove (i) and (ii) one by one.
+    $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|v_{t}^{\mu,\phi}\|_{\infty}\leq c{\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ (t,\mu)\in[0,T]\times\mathscr{P}_{k}.$
+Next, we will prove (i) and (ii) one by one.
 (1) Proof of (i).
 (a) We first verify
-(3.22) $\bar{H}:\mathscr{V}_{0}\times\mathscr{C}_{k}\rightarrow\mathscr{V}_{0},$
-so that $\Phi^{(n)}:\mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{V}_{0}\times\mathscr{C}_{k,n}$ for every $n\geq 1$.
-By (3.3), we have
-(3.23) $I_{t}^{V}(\mu,\phi)=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}\big{\{}V(X_{t}^{x,\mu})\big{\}}\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).$
-Similar to (2.11)-(2.13), for any $\lambda\geq\lambda_{0}$, $\gamma\in\mathscr{P}_{k}$ and $x\in\mathbb{R}^{d}$, consider
-$\text{\rm{d}}\theta^{\lambda,\gamma,B}_{t}(x)=B_{t}((\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma,B}_{t}(x)),P_{t}^{\ast}\gamma,P_{t}V(\gamma))\text{\rm{d}}t,\ \ t\in[0,T],\theta^{\lambda,\gamma,B}_{0}(x)=\Theta^{\lambda,\gamma}_{0}(x),$
-and let
-$\tilde{\theta}^{\lambda,\gamma,B}_{t}(x)=(\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma,B}_{t}(x)),\ \ t\in[0,T],x\in\mathbb{R}^{d}.$
-Then we have
-$\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(\tilde{\theta}^{\lambda,\gamma,B}_{t}(x))=B_{t}(\tilde{\theta}^{\lambda,\gamma,B}_{t}(x),P_{t}^{\ast}\gamma,P_{t}V(\gamma))\text{\rm{d}}t,\ \ t\in[0,T],\tilde{\theta}^{\lambda,\gamma,B}_{0}(x)=x\in\mathbb{R}^{d}.$
-Repeating the proof of Lemma 2.3, we conclude that for any $p\geq 1$, there exists a constant $c_{p}>0$ such that
-(3.24) $\mathbb{E}[|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma,B}_{t}(x)|^{p}]\leq c_{p}t^{\frac{p}{2}},\ \ t\in[0,T],x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},$
+(3.20) $H:\mathscr{V}_{0}\rightarrow\mathscr{V}_{0}.$
+Recall that $\theta^{\lambda,\gamma}$ and $\tilde{\theta}^{\lambda,\gamma}$ are defined in (2.15)-(2.17). Since (B) implies (A), we conclude that (2.18) still holds.
+By $[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)]_{\alpha}\leq 1$ due to $(B_{2})$, (2.10) in Lemma 2.3 for $p=2$ and $z=\tilde{\theta}^{\lambda,\mu}_{t}(x)$, (3.3), (2.18), (2.1) and (3.15), we find a constant $c_{1}>0$ such that
+(3.21) $\begin{split}|I_{t}^{\mu,\phi}(y,\nu)|&\leq\frac{c_{1}}{\sqrt{t}}\int_{\mathbb{R}^{d}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}|\phi(x)|\mu(\text{\rm{d}}x)}\\ &{\leq\frac{c_{1}}{\sqrt{t}}\alpha\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}},y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.\end{split}$
+So, by (3.14), $(B_{2})$, (2.10) in Lemma 2.3 for $p=2$ and $z=\tilde{\theta}^{\lambda,\mu}_{t}(x)$, (2.1) and (2.18), we find a constant $c_{2}>0$ such that
+(3.22) $\begin{split}&|\tilde{I}_{t}^{V}(y,\nu)|\leq c_{2}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\bigg{(}\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}}\|V_{s}\|^{2}_{\infty}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.\end{split}$
+Combining this with (3.19) and (3.21), we find a constant $c_{3}>0$ such that
+(3.23) $\begin{split}\|\{H(V)\}_{t}\|_{\infty}\leq&\,c_{3}\|\phi\|_{L^{k}(\mu)}\sqrt{\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}+c_{3}\sqrt{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}\big{(}s^{\frac{1}{2}}\big{)}}\|V_{s}\|^{2}_{\infty}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}.\end{split}$
+Then (3.20) follows from the fact that (2.40) implies
+(3.24) $\int_{0}^{t}\frac{\alpha(rs^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(rs^{\frac{1}{2}})}\text{\rm{d}}s=2\int_{0}^{rt^{\frac{1}{2}}}\frac{\alpha(s)^{2}}{s\tilde{\alpha}(s)}\text{\rm{d}}s=4\int_{0}^{rt^{\frac{1}{2}}}\tilde{\alpha}^{\prime}(s)\text{\rm{d}}s=4\tilde{\alpha}(rt^{\frac{1}{2}}),\ \ r\geq 0.$
+(b) We intend to prove that $H$ in (i) has a unique fixed point in $\mathscr{V}_{0}.$ Obviously, for any $\delta>0$, $\mathscr{V}_{0}$ is complete under the metric
+$\rho_{\delta}\big{(}V,U\big{)}:=\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta t}\|V_{t}-U_{t}\|_{\infty},\ \ V,U\in\mathscr{V}_{0}.$
+So, it suffices to prove the contraction of $H$ in $\rho_{\delta}$ for large enough $\delta>0$.
+By (3.19), (3.21) and (3.22), we find a constant $c_{4}>0$ such that
+|$|\{H(V)\}_{t}(y,\nu)-\{H(U)\}_{t}(y,\nu)|=\frac{\{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\}^{\frac{1}{2}}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}|\tilde{I}_{t}^{V-U}(y,\nu)|$
+$\leq c_{4}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}\|V_{s}-U_{s}\|^{2}_{\infty}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ V,U\in\mathscr{V}_{0},t\in[0,T].$
+Combining this with (3.24), we conclude that $H$ is contractive in the complete metric space $(\mathscr{V}_{0},\rho_{\delta})$ for large enough $\delta>0$, and hence has a unique fixed point denoted by $v^{\mu,\phi}$.
+(2) Proof of (ii). By (3.23) and noting that $H(v^{\mu,\phi})=v^{\mu,\phi}$, we derive
+$\|v_{t}^{\mu,\phi}\|_{\infty}^{2}\leq 2c_{3}^{2}\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)}^{2}+2c_{3}^{2}t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}\|v_{s}^{\mu,\phi}\|^{2}_{\infty}\text{\rm{d}}s,\ \ t\in[0,T].$
+Combining this with (3.24) and Gronwall’s inequality, we find a constant $c_{5}>0$ such that for any $t\in[0,T]$,
+$\|v_{t}^{\mu,\phi}\|_{\infty}\leq c_{5}\|\phi\|_{L^{k}(\mu)}\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},\ \ \mu\in\mathscr{P}_{k},\ \phi\in T_{\mu,k}.$
+This proves (ii). ∎
+3.3 Proof of Theorem 3.3
+By Lemma 3.4, the proof of Theorem 3.3(1) is completed by the following lemma.
+Lemma 3.5.
+Assume (B). Then for any $\mu\in\mathscr{P}_{k}$, $\phi\in T_{\mu,k}$, the function $h:(0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}\rightarrow\mathbb{R}^{d}$ defined by
+$h_{t}(y,\nu):=D_{\phi}^{I}\big{\{}P_{t}[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)](\mu)\big{\}},\ \ t\in(0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}$
+[…]
+$\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\bigg{\{}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\sup_{y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}}|D_{\phi}^{I}\big{\{}P_{t}[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)](\mu)\big{\}}|\bigg{\}}\leq c.$
+Proof.
+(a) By Lemma 3.4, it suffices to prove (3.17). For simplicity, for any $t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}$, let
+(3.25) $U_{t}(y,\nu,z):=D^{E}b_{t}^{(1)}(y,\nu)(z),\ \ z\in\mathbb{R}^{d}.$
+Moreover, simply denote
+$v_{t}^{\varepsilon}(y,\nu):=v_{t}^{\varepsilon,1}(y,\nu)+v_{t}^{\varepsilon,2}(y,\nu),$
+$v_{t}^{\varepsilon,1}(y,\nu):=\frac{P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu)}{\varepsilon},$
+$v_{t}^{\varepsilon,2}(y,\nu):=\frac{P_{t}^{\mu_{\varepsilon}}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})}{\varepsilon}.$
+Next, for $v_{t}^{\mu,\phi}$ in Lemma 3.4, let
+(3.26) $\begin{split}&\hat{v}_{t}(y,\nu):={\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}v_{t}^{\mu,\phi}(y,\nu),\ \ t\in(0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k},\end{split}$
 and
-(3.25) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{v}X_{t}^{x,\gamma}|^{p}\Big{]}\leq c_{p}|v|^{p},\ \ v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}.$
-By $[V]_{\alpha}\leq 1$, (2.22) in Lemma 2.5 for $p=2$ and $y=\tilde{\theta}^{\lambda_{0},\mu,B}_{t}(x)$, (3.25), (2.1), (3.24) and (3.23), we find a constant $c_{1}>0$ such that
-(3.26) $\begin{split}\|I_{t}^{V}(\mu,\phi)\|_{\mathbb{B}}&\leq\frac{c_{1}}{\sqrt{t}}\int_{\mathbb{R}^{d}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}|\phi(x)|\mu(\text{\rm{d}}x)}\\ &{\leq\frac{c_{1}}{\sqrt{t}}\alpha\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}}.\end{split}$
-To estimate $\tilde{I}_{t}^{V}(\mu,v,\tilde{v})$, noting that $X_{t}^{\mu}$ solves (3.9) for $\gamma=\mu$ and initial value $X_{0}^{\mu}$ replacing $x$, by the standard Markov property for solutions of (3.9), we have
-$\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}V(X_{t}^{x,\mu})\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})}{\{\tilde{\alpha}(s^{\frac{1}{2}})s\}^{\frac{1}{2}}}}$
-×$\qquad\qquad\times\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{v_{s}}^{\mathbb{B}}B_{s}(X_{s}^{x,\mu},P_{s}^{*}\mu,\cdot)(P_{s}V(\mu)),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),$
-$\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v})=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}V(X_{t}^{x,\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})$
-×$\qquad\qquad\times\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{\ast}\mu)(X_{s}^{\mu}),(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}\big{\rangle}\big{]}|_{y=X_{s}^{x,\mu}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).$
-So, by $(B_{2})$, $(B_{3})$, (2.22) in Lemma 2.5 for $p=2$ and $y=\tilde{\theta}^{\lambda_{0},\mu,B}_{t}(x)$, (2.1), (3.24) and Jensen’s inequality, we find a constant $c_{2}>0$ such that
-(3.27) $\begin{split}&\|\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})\|_{\mathbb{B}}\leq c_{2}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\bigg{(}\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}}\|v_{s}\|^{2}_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\\ &\|\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v})\|_{\mathbb{B}}\leq c_{2}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\left(\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}|^{k})^{\frac{2}{k}}\text{\rm{d}}s\right)^{\frac{1}{2}}.\end{split}$
-Combining this with (3.19), (3.20) and (3.26), we find a constant $c_{3}>0$ such that
-(3.28) $\begin{split}\|\bar{H}_{t}(v,\tilde{v})\|_{\mathbb{B}}\leq&\,c_{3}\|\phi\|_{L^{k}(\mu)}\sqrt{\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\\ &+c_{3}\sqrt{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}\big{(}s^{\frac{1}{2}}\big{)}}\|v_{s}\|^{2}_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\\ &+c_{3}\sqrt{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\left(\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}|^{k})^{\frac{2}{k}}\text{\rm{d}}s\right)^{\frac{1}{2}}.\end{split}$
-Then (3.22) holds.
-(b) We intend to prove that $\Phi^{(n)}$ in (i) has a unique fixed point in $\mathscr{V}_{0}\times\mathscr{C}_{k,n}.$ Obviously, for any $\delta>0$, $\mathscr{V}_{0}\times\mathscr{C}_{k}$ is complete under the metric
-$\rho_{\delta,n}\big{(}(v,\tilde{v}),(w,\tilde{w})\big{)}:=\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta t}\|v_{t}-w_{t}\|_{\mathbb{B}}+\rho_{\delta,n}^{(k)}\big{(}\tilde{v},\tilde{w}\big{)},$
-$\rho_{\delta,n}^{(k)}\big{(}\tilde{v},\tilde{w}\big{)}:=\bigg{(}\mathbb{E}\Big{[}\sup_{t\in[0,\tau_{n}\land T]}\text{\rm{e}}^{-\delta kt}|\tilde{v}_{t}-\tilde{w}_{t}|^{k}\Big{]}\bigg{)}^{\frac{1}{k}},\ \ v,w\in\mathscr{V}_{0},\tilde{v},\tilde{w}\in\mathscr{C}_{k}.$
-So, it suffices to prove the contraction of $\Phi^{(n)}$ in $\rho_{\delta,n}$ for large enough $\delta>0$.
-By the definition of $H_{t}$, $\|\nabla(B+\lambda_{0}u)\|_{\infty}<\infty$, and repeating step (2) in the proof of [29, Lemma 4.1], we find a function $c_{n}:(0,\infty)\rightarrow(0,\infty)$ with $c_{n}(\delta)\rightarrow 0$ as $\delta\rightarrow\infty$, such that
-(3.29) $\begin{split}&\rho_{\delta,n}^{(k)}(H(v,\tilde{v}),H(w,\tilde{w})):=\bigg{(}\mathbb{E}\Big{[}\sup_{t\in[0,\tau_{n}]}\text{\rm{e}}^{-\delta kt}|H_{t}(v,\tilde{v})-H_{t}(w,\tilde{w})|^{k}\Big{]}\bigg{)}^{\frac{1}{k}}\\ &\leq\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta t}\|v_{t}-w_{t}\|_{\mathbb{B}}\sup_{s\in[0,T]}\int_{0}^{s}\text{\rm{e}}^{-\delta(s-t)}{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\text{\rm{d}}t\\ &+c(\delta,n)\rho_{\delta,n}^{k}(\tilde{v},\tilde{w}),\ \ n\geq 1,(v,\tilde{v}),(w,\tilde{w})\in\mathscr{V}_{0}\times\mathscr{C}_{k,n},\ \delta>0.\end{split}$
-Moreover, by (3.20), (3.26) and (3.27), we find a constant $c_{4}>0$ such that
-$\|\bar{H}_{t}(v,\tilde{v})-\bar{H}_{t}(w,\tilde{w})\|_{\mathbb{B}}=\frac{\{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\}^{\frac{1}{2}}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\|\tilde{I}_{t}^{V}(\mu,v-w,\tilde{v}-\tilde{w})\|_{\mathbb{B}}$
-$\leq c_{4}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}\|v_{s}-w_{s}\|^{2}_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}+c_{4}\left(\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}-\tilde{w}_{s}|^{k})^{\frac{2}{k}}\text{\rm{d}}s\right)^{\frac{1}{2}},$
-$\qquad\ v,w\in\mathscr{V}_{0},\tilde{v},\tilde{w}\in\mathscr{C}_{k,n},t\in[0,T].$
-Combining this with (3.19) and (3.29), we conclude that $\Phi^{(n)}$ is contractive in the complete metric space $(\mathscr{V}_{0}\times\mathscr{C}_{k,n},\rho_{\delta,n})$ for large enough $\delta>0$, and hence has a unique fixed point.
-(2) Proof of (ii).
-Combining the first equation in (3.14), (B), Itô’s formula, the maximal inequality in [30, Lemma 2.1], Krylov’s estimate (see [32, Theorem 3.1]) and hence Khasminskii’s estimate (see [30, Lemma 4.1]), for any $j\geq k$ we find a constant $c_{1}>0$ such that
-d $|{\tilde{v}}^{n}_{t}|^{j+1}$
-$\leq c_{1}\bigg{\{}|{\tilde{v}}^{n}_{t}|^{j+1}\text{\rm{d}}\mathscr{L}_{t}+\big{(}\mathbb{E}[|\tilde{v}_{t}^{n}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}t$
-$\qquad\quad+|{\tilde{v}}_{t}^{n}|^{j}{\frac{\alpha(t^{\frac{1}{2}})}{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}}\|v_{t}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}t+\text{\rm{d}}M_{t}\bigg{\}},\ \ t\in[0,\tau_{n}]$
-for some local martingale $M_{t}$ and an adapted increasing process $\mathscr{L}_{t}$ with $\mathbb{E}[\text{\rm{e}}^{p\mathscr{L}_{T}}]<\infty$ for any $p>0$.
-Then by the stochastic Gronwall inequality (see [31, Lemma 3.7]), for any $j\geq k$, we find constants $c_{2},c_{3}>0$ such that
-$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{2}|\phi(X_{0}^{\mu})|^{j}+c_{2}\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}}$
-$\qquad\qquad\quad+c_{2}\left(\mathbb{E}\bigg{(}\int_{0}^{t}|{\tilde{v}}_{t}^{n}|^{j}{\frac{\alpha(s^{\frac{1}{2}})}{(\tilde{\alpha}(s^{\frac{1}{2}})s)^{\frac{1}{2}}}}\|v_{s}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\bigg{)}\right)^{\frac{j}{j+1}}$
-$\qquad\qquad\quad\leq c_{3}|\phi(X_{0}^{\mu})|^{j}+c_{3}\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}}$
-$\qquad\qquad\quad+c_{3}\left(\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})}{(\tilde{\alpha}(s^{\frac{1}{2}})s)^{\frac{1}{2}}}}\|v_{s}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}s\right)^{j}+\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}.$
-Noting that (3.19) and the Schwartz inequality imply
-$\bigg{(}\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})}{(\tilde{\alpha}(s^{\frac{1}{2}})s)^{\frac{1}{2}}}}\|v_{s}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{2}\leq 4{\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s,$
-we find a constant $c>0$ such that
-(3.30) $\begin{split}&\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c|\phi(X_{0}^{\mu})|^{j}+c\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}}\\ &\qquad+c\left({\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\right)^{\frac{j}{2}},\ \ t\in[0,T].\end{split}$
-Choosing $j=k$, taking expectations to both sides, and noting that
-(3.31) $\begin{split}&c\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{k+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{k}{k+1}}\\ &\leq c_{4}\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\leq\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}+\frac{c_{4}^{2}}{2}\int_{0}^{t}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\text{\rm{d}}s\end{split}$
-holds for some constant $c_{4}>c$, we derive
-$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}\leq$  $\,2c_{4}\|\phi\|_{L^{k}(\mu)}^{k}+c_{4}^{2}\int_{0}^{t}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\text{\rm{d}}s$
-$+2c_{4}\left({\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\right)^{\frac{k}{2}}.$
-So, Gronwall’s inequality yields
-(3.32) $\begin{split}&\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}\leq 2c_{4}\text{\rm{e}}^{c_{4}^{2}T}\bigg{\{}\|\phi\|_{L^{k}(\mu)}^{k}+\left({\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\right)^{\frac{k}{2}}\bigg{\}}.\end{split}$
-On the other hand, by taking $(v,\tilde{v})=(v^{(n)},\tilde{v}^{(n)})$ in (3.28) and noting that $\bar{H}_{t}(v,\tilde{v})=v_{t}^{(n)}$ due to (3.21), we find a constant $c_{5}>0$ such that
-$\|v_{t}^{(n)}\|_{\mathbb{B}}^{2}\leq c_{5}\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)}^{2}+c_{5}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}\|v_{s}^{(n)}\|^{2}_{\mathbb{B}}\text{\rm{d}}s+c_{5}\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}^{(n)}|^{k})^{\frac{2}{k}}\text{\rm{d}}s,\ \ t\in[0,T].$
-Combining this with (3.19), (3.32) and Gronwall’s inequality, we find a constant $c_{6}>0$ such that for any $n\geq 1$ and $t\in[0,T]$,
-$\|v_{t}^{(n)}\|_{\mathbb{B}}\leq c_{6}\|\phi\|_{L^{k}(\mu)}\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},$
-$\mathbb{E}\Big{[}\sup_{s\in[0,t\land\tau_{n}]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}\leq c_{6}\|\phi\|_{L^{k}(\mu)}^{k},\ \ \mu\in\mathscr{P}_{k},\ \phi\in T_{\mu,k}.$
-This together with (3.30) proves (ii). ∎
-3.3 Proof of Theorem 3.2
-By Lemma 3.4, the proof of Theorem 3.2(1)-(2) is completed by the following lemma.
-Lemma 3.5.
-[…]
-$\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\bigg{\{}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\|D_{\phi}P_{t}V(\mu)\|_{\mathbb{B}}+\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{\phi}X_{t}^{\mu}|^{k}\Big{]}\bigg{\}}\leq c.$
-Proof.
-By Lemma 3.4, it suffices to prove (3.15). For any $t\in[0,T]$ and $\varepsilon\in(0,1]$, (3.10) for $V$ replacing $f$ implies
-(3.33) $\begin{split}&v_{t}^{\varepsilon}:=\frac{P_{t}V(\mu_{\varepsilon})-P_{t}V(\mu)}{\varepsilon}=v_{t}^{\varepsilon,1}+v_{t}^{\varepsilon,2},\\ &v_{t}^{\varepsilon,1}:=\frac{P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)}{\varepsilon},\ \ v_{t}^{\varepsilon,2}:=\frac{P_{t}^{\mu_{\varepsilon}}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu_{\varepsilon})}{\varepsilon}.\end{split}$
-Next, let
-(3.34) $\begin{split}&\hat{v}_{t}:={\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}v_{t}(\mu,\phi),\\ &\tilde{v}_{t}:=\tilde{v}_{t}(\mu,\phi),\ \ \tilde{v}_{t}^{\varepsilon}=\frac{\Theta_{t}(X_{t}^{\mu_{\varepsilon}})-\Theta_{t}(X_{t}^{\mu})}{\varepsilon}.\end{split}$
-Then (3.15) follows from
-(3.35) $\lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\big{\|}v_{t}^{\varepsilon}-\hat{v}_{t}\big{\|}_{\mathbb{B}}=0,\ \ \lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{t}|^{k}\Big{]}=0.$
-In the following, we first estimate $\|v_{s}^{\varepsilon,1}\|_{\mathbb{B}}$, $\|v_{s}^{\varepsilon,2}\|_{\mathbb{B}}$, $|\tilde{v}_{s}^{\varepsilon}|$, $\|v_{t}^{\varepsilon}-\hat{v}_{t}\big{\|}_{\mathbb{B}}$ and $|\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{t}|$ in (a)-(e), then we prove (3.35) in (f).
-(a) For $v_{t}^{\varepsilon,1}$. Let
-$V_{N}:=\frac{(N\land\|V\|_{\mathbb{B}})V}{\|V\|_{\mathbb{B}}},\ \ N\geq 1.$
-By (3.11), we have
-$\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu}V_{N}(\mu_{\varepsilon})-P_{t}^{\mu}V_{N}(\mu)}{\varepsilon}=I_{t}^{V_{N}}(\mu,\phi),$
-$P_{t}^{\mu}V_{N}(\mu_{\varepsilon})-P_{t}^{\mu}V_{N}(\mu)=\int_{0}^{\varepsilon}\frac{\text{\rm{d}}}{\text{\rm{d}}r}P_{t}^{\mu}V_{N}(\mu_{r})\text{\rm{d}}r=\int_{0}^{\varepsilon}I_{t}^{V_{N}}(r,\mu,\phi)\text{\rm{d}}r,\ \ N\geq 1.$
-Since $(B_{2})$ implies
-(3.36) ${\|V_{N}\|_{\mathbb{B}}\leq\|V\|_{\mathbb{B}}\leq c(1+|\cdot|^{k-1})}$
-for some constant $c>0$, by (2.14), (3.24), the same argument to (3.26), and the dominated convergence theorem for $N\rightarrow\infty$, we find a constant $c(\mu,\phi)>0$ depending on $\phi,\mu$ such that
-(3.37) $\|v_{t}^{\varepsilon,1}\|_{\mathbb{B}}=\frac{\|P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)\|_{\mathbb{B}}}{\varepsilon}\leq\frac{c(\mu,\phi)\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \ \varepsilon\in(0,1],t\in(0,T],$
+(3.27) $\hat{v}_{t}^{1}(y,\nu):=I_{t}^{U_{t}(y,\nu,\cdot)}(\mu,\phi),\ \ \hat{v}_{t}^{2}(y,\nu):=\hat{v}_{t}(y,\nu)-I_{t}^{U_{t}(y,\nu,\cdot)}(\mu,\phi).$
+Noting that
+$P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu)$
+(3.28) $=\int_{\mathbb{R}^{d}}\Big{[}P_{t}^{\mu}U_{t}(y,\nu,\cdot)(x+\varepsilon\phi(x))-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(x)\Big{]}\mu(\text{\rm{d}}x),$
+by (3.4) for $f=U_{t}(y,\nu,\cdot)$, we obtain
+(3.29) $\lim_{\varepsilon\rightarrow 0}\big{|}v_{t}^{\varepsilon,1}(y,\nu)-\hat{v}_{t}^{1}(y,\nu)\big{|}=0,\ \ t\in(0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.$
+Since (3.16) holds for $V_{t}=v_{t}^{\mu,\phi}$, (3.25)-(3.27) imply that
+(3.30) $\begin{split}&\hat{v}_{t}^{2}(y,\nu)=\hat{v}_{t}(y,\nu)-I_{t}^{U_{t}(y,\nu,\cdot)}(\mu,\phi)\\ &=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}U_{t}(y,\nu,X_{t}^{x,\mu})\\ &\qquad\qquad\quad\times\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})[\hat{v}_{s}^{2}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)+\hat{v}_{s}^{1}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)],\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).\end{split}$
+In view of (3.29), to prove (3.17), it remains to verify
+(3.31) $\lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}|v_{t}^{\varepsilon,2}(y,\nu)-\hat{v}_{t}^{2}(y,\nu)|=0,\ \ y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.$
+In the following, we first estimate $\|v_{t}^{\varepsilon,i}\|_{\infty}$ and $|v_{t}^{\varepsilon,i}-\hat{v}_{t}^{i}|(i=1,2)$ in steps (b)-(c), then verify (3.31) in step (d).
+(b) Estimates on $\|v_{t}^{\varepsilon,i}\|_{\infty},i=1,2$ and $\|v_{t}^{\varepsilon,1}-\hat{v}_{t}^{1}\|_{\infty}$.
+By (3.6) for $f=U_{t}(y,\nu,\cdot)$ and (3.3), we obtain
+$v_{t}^{\varepsilon,1}(y,\nu)=\frac{P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu)}{\varepsilon}$
+$=\frac{1}{\varepsilon}\int_{0}^{\varepsilon}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}\Big{(}U_{t}\big{(}y,\nu,X_{t}^{x+r\phi(x),\mu}\big{)}-U_{t}\big{(}y,\nu,\tilde{\theta}^{\lambda,\mu}_{t}(x+r\phi(x))\big{)}\Big{)}$
+(3.32) ×$\qquad\qquad\qquad\qquad\times\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X_{s}^{x+r\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+r\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x)\text{\rm{d}}r$
+$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}\Big{(}U_{t}\big{(}y,\nu,X_{t}^{x+\varepsilon u\phi(x),\mu}\big{)}-U_{t}\big{(}y,\nu,\tilde{\theta}^{\lambda,\mu}_{t}(x+\varepsilon u\phi(x))\big{)}\Big{)}$
+×$\qquad\qquad\qquad\qquad\times\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X_{s}^{x+\varepsilon u\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+\varepsilon u\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x)\text{\rm{d}}u,$
+where in the last step, we used the integral transform $r=\varepsilon u$.
+Similar to (3.21), noting that (B) implies $[U_{t}(y,\nu,\cdot)]_{\alpha}\leq 1$, by (2.10) in Lemma 2.3 for $p=2$ and $z=\tilde{\theta}^{\lambda,\mu}_{t}(x+\varepsilon r\phi(x))$, (3.3), (2.18) and (2.1), we find a constant $c(\mu,\phi)>0$ depending on $\phi,\mu$ such that
+(3.33) $\begin{split}&\|v_{t}^{\varepsilon,1}\|_{\infty}=\sup_{y,\nu}\frac{|P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu)|}{\varepsilon}\\ &\leq\sup_{y,\nu}\int_{0}^{1}\int_{\mathbb{R}^{d}}\bigg{|}\mathbb{E}\bigg{[}\Big{(}U_{t}\big{(}y,\nu,X_{t}^{x+\varepsilon r\phi(x),\mu}\big{)}-U_{t}\big{(}y,\nu,\tilde{\theta}^{\lambda,\mu}_{t}(x+\varepsilon r\phi(x)))\Big{)}\\ &\qquad\qquad\qquad\qquad\times\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X_{s}^{x+\varepsilon r\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+\varepsilon r\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\bigg{|}\mu(\text{\rm{d}}x)\text{\rm{d}}r\\ &\leq\frac{c(\mu,\phi)\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \ \varepsilon\in(0,1],t\in(0,T].\end{split}$
+This together with (3.29) and (3.21) implies that for a constant $c(\mu,\phi)>0$
+$h_{t,1}^{\varepsilon}(y,\nu):=\big{\{}v_{t}^{\varepsilon,1}(y,\nu)-\hat{v}_{t}^{1}(y,\nu)\big{\}}\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}$
+satisfies
+(3.34) $\lim_{\varepsilon\rightarrow 0}|h_{t,1}^{\varepsilon}(y,\nu)|=0,\ \ \sup_{\varepsilon\in(0,1]}\sup_{y,\nu}|h_{t,1}^{\varepsilon}(y,\nu)|\leq c(\mu,\phi)\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},\ \ t\in(0,T].$
+Next, we estimate $\|v_{t}^{\varepsilon,2}\|_{\infty}.$ Recall that $X_{t}^{x+\varepsilon\phi(x),\mu}$ solves (3.2) with initial value $x+\varepsilon\phi(x)$.
+For any $x\in\mathbb{R}^{d},s,t\in[0,T]$, let
+(3.35) $\begin{split}&R_{t}^{\varepsilon,x}:=\text{\rm{e}}^{\int_{0}^{t}\langle\eta_{s}^{\varepsilon,x},\text{\rm{d}}W_{s}\rangle-\frac{1}{2}\int_{0}^{t}|\eta_{s}^{\varepsilon,x}|^{2}\text{\rm{d}}s},\\ &\eta_{s}^{\varepsilon,x}:=\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})\\ &\qquad\quad\times\big{\{}b_{s}^{(1)}(X_{s}^{x+\varepsilon\phi(x),\mu},P_{s}^{\ast}\mu_{\varepsilon})-b_{s}^{(1)}(X_{s}^{x+\varepsilon\phi(x),\mu},P_{s}^{\ast}\mu)\big{\}}.\end{split}$
+By [21, Lemma 3.2], we have
+$b_{t}^{(1)}(y,P_{t}^{\ast}\mu_{\varepsilon})-b_{t}^{(1)}(y,P_{t}^{\ast}\mu)$
+(3.36) $=\int_{0}^{1}\frac{\text{\rm{d}}}{\text{\rm{d}}r}b_{t}^{(1)}(y,(1-r)P_{t}^{\ast}\mu+rP_{t}^{\ast}\mu_{\varepsilon})\text{\rm{d}}r$
+$=\int_{0}^{1}\int_{\mathbb{R}^{d}}D^{E}b_{t}^{(1)}(y,(1-r)P_{t}^{\ast}\mu+rP_{t}^{\ast}\mu_{\varepsilon})(z)(P_{t}^{\ast}\mu_{\varepsilon}-P_{t}^{\ast}\mu)(\text{\rm{d}}z)\text{\rm{d}}r.$
+Since (B) implies (A), Lemma 2.6 holds so that we find constants $c_{0},c>0$ such that
+(3.37) $\begin{split}|\eta_{s}^{\varepsilon,x}|&\leq c_{0}\mathbb{W}_{\alpha}(P_{s}^{\ast}\mu_{\varepsilon},P_{s}^{\ast}\mu)\leq c\varepsilon\|\phi\|_{L^{k}(\mu)}\frac{\alpha(s^{\frac{1}{2}})}{\sqrt{s}},\ \ s\in[0,T],\ \varepsilon\in[0,1],x\in\mathbb{R}^{d}.\end{split}$
+Then by Girsanov’s theorem, for any $x\in\mathbb{R}^{d}$,
+$W_{t}^{\varepsilon,x}:=W_{t}-\int_{0}^{t}\eta_{s}^{\varepsilon,x}\text{\rm{d}}s,\ \ s\in[0,T]$
+is a Brownian motion under $\mathbb{Q}:=R_{T}^{\varepsilon,x}\mathbb{P}$. Reformulate (3.2) with $x+\varepsilon\phi(x)$ replacing $x$ as
+$\text{\rm{d}}X_{t}^{x+\varepsilon\phi(x),\mu}$ =$=\Big{\{}b_{t}^{(0)}(X_{t}^{x+\varepsilon\phi(x),\mu})+b_{t}^{(1)}(X_{t}^{x+\varepsilon\phi(x),\mu},P_{t}^{*}\mu)\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x+\varepsilon\phi(x),\mu})\text{\rm{d}}W_{t}$
+=$=\Big{\{}b_{t}^{(0)}(X_{t}^{x+\varepsilon\phi(x),\mu})+b_{t}^{(1)}(X_{t}^{x+\varepsilon\phi(x),\mu},P_{t}^{*}\mu_{\varepsilon})\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x+\varepsilon\phi(x),\mu})\text{\rm{d}}W_{t}^{\varepsilon,x},$
+$\qquad\ X_{0}^{x+\varepsilon\phi(x),\mu}=x+\varepsilon\phi(x),\ x\in\mathbb{R}^{d}.$
+By the weak uniqueness of (3.2) with $\mu=\mu_{\varepsilon},$ we get
+$v_{t}^{\varepsilon,2}(y,\nu)$ $=\frac{P_{t}^{\mu_{\varepsilon}}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})}{\varepsilon}$
+(3.38) $=\frac{\int_{\mathbb{R}^{d}}[P_{t}^{\mu_{\varepsilon}}U_{t}(y,\nu,\cdot)(x+\varepsilon\phi(x))-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(x+\varepsilon\phi(x))]\mu(\text{\rm{d}}x)}{\varepsilon}$
+$=\frac{1}{\varepsilon}\int_{\mathbb{R}^{d}}\mathbb{E}[U_{t}(y,\nu,X_{t}^{x+\varepsilon\phi(x),\mu})(R_{t}^{\varepsilon,x}-1)]\mu(\text{\rm{d}}x),\ \ t\in[0,T].$
+By (3.37), for any $p\geq 1$ there exists a constant $c(p,\mu,\phi)>0$ such that
+(3.39) $\mathbb{E}[|R_{t}^{\varepsilon,x}-1|^{p}]\leq c(p,\mu,\phi)\varepsilon^{p}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\text{\rm{d}}s\bigg{)}^{\frac{p}{2}},\ \ t\in[0,T],\varepsilon\in[0,1],x\in\mathbb{R}^{d}.$
+Again by (2.10) in Lemma 2.3 for $p=2$ and $z=\tilde{\theta}^{\lambda,\mu}_{t}(x+\varepsilon\phi(x))$, (3.3), (3.39), (3.3), (2.18) and (2.1), we find a constant $c_{1}(\mu,\phi)>0$ such that
+(3.40) $\|v_{t}^{\varepsilon,2}\|_{\infty}$ $\leq c_{1}(\mu,\phi)\alpha(t^{\frac{1}{2}})\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in[0,T],\varepsilon\in(0,1].$
+This together with (3.33) yields that for some constant $c_{2}(\mu,\phi)>0$,
+$\|v_{t}^{\varepsilon}\|_{\infty}^{2}$ $\leq 2\|v_{t}^{\varepsilon,1}\|_{\infty}^{2}+2\|v_{t}^{\varepsilon,2}\|_{\infty}^{2}$
+$\leq c_{2}(\mu,\phi)\bigg{(}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}+\alpha(t^{\frac{1}{2}})\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\text{\rm{d}}s\bigg{)},\ \ t\in(0,T],\varepsilon\in(0,1].$
+By the definition of $\alpha$ and (2.41), we find a constant $c_{3}(\mu,\phi)>0$ such that
+(3.41) $\|v_{t}^{\varepsilon}\|_{\infty}^{2}\leq c_{3}(\mu,\phi)\frac{\alpha(t^{\frac{1}{2}})^{2}}{t},\ \ t\in(0,T],\ \ \varepsilon\in(0,1].$
+(c) Estimate on $\|v_{t}^{\varepsilon,2}-\hat{v}_{t}^{2}\|_{\infty}$. Similarly to (b), we have
+(3.42) $\begin{split}&\frac{R_{t}^{\varepsilon,x}-1}{\varepsilon}=\int_{0}^{t}R_{s}^{\varepsilon,x}\big{\langle}\varepsilon^{-1}\eta_{s}^{\varepsilon,x},\text{\rm{d}}W_{s}\big{\rangle}\\ &=\int_{0}^{t}R_{s}^{\varepsilon,x}\Big{\langle}\frac{\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})[b_{s}^{(1)}(\cdot,P_{s}^{*}\mu_{\varepsilon})-b_{s}^{(1)}(\cdot,P_{s}^{*}\mu)](X_{s}^{x+\varepsilon\phi(x),\mu})}{\varepsilon},\ \text{\rm{d}}W_{s}\Big{\rangle}\\ &=h_{t}(\varepsilon,x)+\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})v_{s}^{\varepsilon}(X_{s}^{x,\mu},P_{s}^{\ast}\mu),\ \text{\rm{d}}W_{s}\Big{\rangle},\ \ x\in\mathbb{R}^{d},\end{split}$
+where
+$h_{t}(\varepsilon,x):=\int_{0}^{t}\Big{\langle}$ $\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})R_{s}^{\varepsilon,x}\frac{[b_{s}^{(1)}(\cdot,P_{s}^{*}\mu_{\varepsilon})-b_{s}^{(1)}(\cdot,P_{s}^{*}\mu)](X_{s}^{x+\varepsilon\phi(x),\mu})}{\varepsilon}$
+$\qquad\qquad\quad-\zeta_{s}(X_{s}^{x,\mu})v_{s}^{\varepsilon}(X_{s}^{x,\mu},P_{s}^{\ast}\mu),\ \text{\rm{d}}W_{s}\Big{\rangle},\ \ x\in\mathbb{R}^{d}$
+satisfies
+(3.43) $\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|h_{t}(\varepsilon,x)|^{2}\Big{]}=0,\ \ x\in\mathbb{R}^{d}.$
+Indeed, by (3.3) and the definition of $v_{s}^{\varepsilon}$, we have
+$\frac{[b_{s}^{(1)}(\cdot,P_{s}^{\ast}\mu_{\varepsilon})-b_{s}^{(1)}(\cdot,P_{s}^{\ast}\mu)]((X_{s}^{x+\varepsilon\phi(x),\mu})}{\varepsilon}$
+$=\frac{1}{\varepsilon}\int_{0}^{1}\int_{\mathbb{R}^{d}}D^{E}b_{s}^{(1)}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})(z)(P_{s}^{\ast}\mu_{\varepsilon}-P_{s}^{\ast}\mu)(\text{\rm{d}}z)\text{\rm{d}}r$
+$=\int_{0}^{1}v_{s}^{\varepsilon}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})\text{\rm{d}}r.$
+This together with the BDG inequality implies
+$\mathbb{E}\Big{[}\sup_{t\in[0,T]}|h_{t}(\varepsilon,x)|^{2}\Big{]}$
+(3.44) $\leq 2\int_{0}^{T}\mathbb{E}\Big{|}\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})R_{s}^{\varepsilon,x}\int_{0}^{1}v_{s}^{\varepsilon}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})\text{\rm{d}}r$
+$\qquad\qquad\quad-\zeta_{s}(X_{s}^{x,\mu})v_{s}^{\varepsilon}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)\Big{|}^{2}\text{\rm{d}}s.$
+By (3.3), for any $p>1$, we can find a constant $c_{p}>0$ such that
+(3.45) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{x+\varepsilon\phi(x),\mu}-X_{t}^{x,\mu}|^{p}\Big{]}\leq c_{p}|\phi(x)|^{p}\varepsilon^{p},\ \ \varepsilon\in[0,1],\mu\in\mathscr{P}_{k}.$
+By the boundedness and continuity of $\zeta$ due to (B), $\int_{0}^{T}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}\text{\rm{d}}t<\infty$, (3.45), (3.39), (3.41), (3.3), and the dominated convergence theorem, to prove (3.43), it is sufficient to prove that for $(s,x,r)\in(0,T]\times\mathbb{R}^{d}\times[0,1]$,
+(3.46) $\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{|}v_{s}^{\varepsilon}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})-v_{s}^{\varepsilon}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)\Big{|}=0.$
+For any $(\omega,\omega^{\prime})\in\Omega\times\Omega$, let
+$U^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})=U_{s}(X_{s}^{x+\varepsilon\phi(x),\mu}(\omega^{\prime}),(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon},X_{s}^{y+\varepsilon u\phi(y),\mu}(\omega)),$
+$U^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})=U_{s}(X_{s}^{x+\varepsilon\phi(x),\mu}(\omega^{\prime}),(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon},\tilde{\theta}^{\lambda,\mu}_{s}(y+\varepsilon u\phi(y))),$
+$\tilde{U}^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})=U_{s}(X_{s}^{x,\mu}(\omega^{\prime}),P_{s}^{\ast}\mu,X_{s}^{y+\varepsilon u\phi(y),\mu}(\omega)),$
+$\tilde{U}^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})=U_{s}(X_{s}^{x,\mu}(\omega^{\prime}),P_{s}^{\ast}\mu,\tilde{\theta}^{\lambda,\mu}_{s}(y+\varepsilon u\phi(y))).$
+Since (B) implies (A), (2.29) holds such that
+(3.47) $\mathbb{W}_{k}((1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon},P_{s}^{\ast}\mu)\leq r\mathbb{W}_{k}(P_{s}^{\ast}\mu_{\varepsilon},P_{s}^{\ast}\mu)\leq cr\varepsilon\|\phi\|_{L^{k}(\mu)}.$
+By (3.3), (3.3) and Hölder’s inequality, we conclude that for any $\beta\in(1,k)$,
+$\mathbb{E}|v_{s}^{\varepsilon,1}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})-v_{s}^{\varepsilon,1}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)|$
+$\leq\int_{0}^{1}\int_{\mathbb{R}^{d}}\int_{\Omega\times\Omega}\bigg{|}\big{[}(U^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-U^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))$
+$\qquad\qquad\qquad\quad-(\tilde{U}^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-\tilde{U}^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))\big{]}$
+×$\qquad\qquad\quad\times\int_{0}^{s}\frac{1}{s}\big{\langle}\zeta_{v}(X_{v}^{y+\varepsilon u\phi(y),\mu})\nabla_{\phi(y)}X_{v}^{y+\varepsilon u\phi(y),\mu},\text{\rm{d}}W_{v}\big{\rangle}\bigg{|}\text{\rm{d}}\mathbb{P}(\omega)\text{\rm{d}}\mathbb{P}(\omega^{\prime})\mu(\text{\rm{d}}y)\text{\rm{d}}u$
+$\leq c_{0}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{\sqrt{s}}|\phi(y)|\Big{\{}\int_{\Omega\times\Omega}\bigg{|}(U^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-U^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))$
+$\qquad\qquad\qquad\quad-(\tilde{U}^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-\tilde{U}^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))\bigg{|}^{\beta}\text{\rm{d}}\mathbb{P}(\omega)\text{\rm{d}}\mathbb{P}(\omega^{\prime})\Big{\}}^{\frac{1}{\beta}}\mu(\text{\rm{d}}y)\text{\rm{d}}u.$
+By (2.9) for $\eta=\alpha(\xi)^{k-1}$ and $p=\frac{k}{k-1}$, we obtain $\|\alpha(\xi)\|_{L^{k}(\mathbb{P})}\leq\alpha(\|\xi\|_{L^{k}(\mathbb{P})})$, which together with (2.18) implies
+$\int_{\Omega\times\Omega}\bigg{|}(U^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-U^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))$
+$\qquad\qquad\qquad\quad-(\tilde{U}^{1,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime})-\tilde{U}^{2,\varepsilon}_{r}(x,y,s,u,\omega,\omega^{\prime}))\bigg{|}^{k}\text{\rm{d}}\mathbb{P}(\omega)\text{\rm{d}}\mathbb{P}(\omega^{\prime})$
+$\leq 2^{k}\mathbb{E}\alpha(|X_{s}^{y+\varepsilon u\phi(y),\mu}-\tilde{\theta}^{\lambda,\mu}_{s}(y+\varepsilon u\phi(y))|)^{k}$
+$\leq 2^{k}\alpha\left((\mathbb{E}|X_{s}^{y+\varepsilon u\phi(y),\mu}-\tilde{\theta}^{\lambda,\mu}_{s}(y+\varepsilon u\phi(y))|^{k})^{\frac{1}{k}}\right)^{k}$
+$\leq c_{k}\alpha(\sqrt{s})^{k}$
+for some constant $c_{k}>0$.
+So, it follows from the fact that $D^{E}b_{t}^{(1)}(y,\nu)(z)$ is continuous in $(y,\nu,z)\in\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{R}^{d}$ due to ($B_{2}$), (3.47), (3.45), (3.25) and the dominated convergence theorem that
+$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{|}v_{s}^{\varepsilon,1}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})\text{\rm{d}}r-v_{s}^{\varepsilon,1}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)\Big{|}=0.$
+Similarly, by (3.3) and (3.39), we have
+$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{|}v_{s}^{\varepsilon,2}(X_{s}^{x+\varepsilon\phi(x),\mu},(1-r)P_{s}^{\ast}\mu+rP_{s}^{\ast}\mu_{\varepsilon})\text{\rm{d}}r-v_{s}^{\varepsilon,2}(X_{s}^{x,\mu},P_{s}^{\ast}\mu)\Big{|}=0.$
+Therefore, (3.46) holds, which implies (3.43) as explained before (3.46).
+Moreover, by (3.43), (3.34), (2.18), (3.41), (3.3) and the argument leading to (3.21), we obtain from the dominated convergence theorem that
+(3.48) $\lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\int_{\mathbb{R}^{d}}\|\mathbb{E}[U_{t}(\cdot,\cdot,X_{t}^{x,\mu})h_{t}(\varepsilon,x)]\|_{\infty}\mu(\text{\rm{d}}x)=0,$
 and
-$\lim_{\varepsilon\downarrow 0}\Big{\|}\frac{P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)}{\varepsilon}-I_{t}^{V}(\mu,\phi)\Big{\|}_{\mathbb{B}}$
-$\leq\lim_{N\rightarrow\infty}\lim_{\varepsilon\downarrow 0}\bigg{\{}\|I_{t}^{V-V_{N}}(\mu,\phi)\|_{\mathbb{B}}+\frac{\|P_{t}^{\mu}(V-V_{N})(\mu_{\varepsilon})-P_{t}^{\mu}(V-V_{N})(\mu)\|_{\mathbb{B}}}{\varepsilon}\bigg{\}}$
-$\leq\lim_{N\rightarrow\infty}\frac{c\|\phi\|_{L^{k}(\mu)}}{\sqrt{t}}\bigg{(}\int_{\mathbb{R}^{d}}\mathbb{E}[\|(V-V_{N})(X_{t}^{x,\mu})\|_{\mathbb{B}}^{k^{*}}]\mu(\text{\rm{d}}x)\bigg{)}^{\frac{1}{k^{*}}}=0,\ \ t\in(0,T].$
-Hence,
-$h_{t,1}(\varepsilon):=\big{\{}v_{t}^{\varepsilon,1}-I_{t}^{V}(\mu,\phi)\big{\}}\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}$
-satisfies
-(3.38) $\lim_{\varepsilon\rightarrow 0}\|h_{t,1}(\varepsilon)\|_{\mathbb{B}}=0,\ \ \sup_{\varepsilon\in(0,1]}\|h_{t,1}(\varepsilon)\|_{\mathbb{B}}\leq 2c(\mu,\phi)\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},\ \ t\in[0,T].$
-(b) For $v_{t}^{\varepsilon,2}$ and $v_{t}^{\varepsilon}$. Recall that $X_{t}^{\mu_{\varepsilon},\mu}$ solves (3.2) with initial value $X_{0}^{\varepsilon}$. Let
-(3.39) $\begin{split}&R_{t}^{\varepsilon}:=\text{\rm{e}}^{\int_{0}^{t}\langle\eta_{s}^{\varepsilon},\text{\rm{d}}W_{s}\rangle-\frac{1}{2}\int_{0}^{t}|\eta_{s}^{\varepsilon}|^{2}\text{\rm{d}}s},\\ &\eta_{s}^{\varepsilon}:=\zeta_{s}(X_{s}^{\mu_{\varepsilon},\mu})\\ &\qquad\quad\times\big{\{}B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{\ast}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{\ast}\mu,P_{s}V(\mu))\big{\}},\ \ s,t\in[0,T].\end{split}$
-Since (B) implies (A), (2.24) holds so that there exists a constant $c>0$ such that
-$\sup_{x\in\mathbb{R}^{d},z\in\mathbb{B},t\in[0,T]}|B_{t}(x,P_{s}^{*}\mu_{\varepsilon},z)-B_{t}(x,P_{s}^{*}\mu,z)|\leq c\mathbb{W}_{k}(\mu_{\varepsilon},\mu)\leq c\varepsilon\|\phi\|_{L^{k}(\mu)},\ \ \varepsilon\in[0,1].$
-Hence, by (B) we find a constant $c(\mu,\phi)>0$ such that
-(3.40) $\begin{split}|\eta_{s}^{\varepsilon}|&\leq K\|P_{s}V(\mu_{\varepsilon})-P_{s}V(\mu)\|_{\mathbb{B}}+c\varepsilon\|\phi\|_{L^{k}(\mu)}\\ &\leq c(\mu,\phi)\varepsilon\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}\big{)},\ \ s\in[0,T],\ \varepsilon\in(0,1].\end{split}$
-Then by Girsanov’s theorem,
-$W_{t}^{\varepsilon}:=W_{t}-\int_{0}^{t}\eta_{s}^{\varepsilon}\text{\rm{d}}s,\ \ s\in[0,T]$
-is a Brownian motion under $\mathbb{Q}:=R_{T}^{\varepsilon}\mathbb{P}$. Reformulate (3.9) with $\gamma=\mu$ as
-$\text{\rm{d}}X_{t}^{\mu_{\varepsilon},\mu}$ =$=\Big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\mu})+B_{t}(X_{t}^{\mu_{\varepsilon},\mu},P_{t}^{*}\mu,P_{t}V(\mu))\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\mu})\text{\rm{d}}W_{t}$
-=$=\Big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\mu})+B_{t}(X_{t}^{\mu_{\varepsilon},\mu},P_{t}^{*}\mu_{\varepsilon},P_{t}V(\mu_{\varepsilon}))\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\mu})\text{\rm{d}}W_{t}^{\varepsilon},$
-$\qquad\ X_{0}^{\mu_{\varepsilon},\mu}=X_{0}^{\mu_{\varepsilon},\mu_{\varepsilon}}=X_{0}^{\mu_{\varepsilon}}=X_{0}^{\mu}+\varepsilon\phi(X_{0}^{\mu}).$
-By the weak uniqueness of (3.9) with $\gamma=\mu_{\varepsilon},$ we get
-(3.41) $v_{t}^{\varepsilon,2}:=\frac{P_{t}^{\mu_{\varepsilon}}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu_{\varepsilon})}{\varepsilon}=\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu_{\varepsilon},\mu})(R_{t}^{\varepsilon}-1)],\ \ t\in[0,T].$
-Combinig this with $[V]_{\alpha}\leq 1,$ and using Lemma 2.5 for $p=k^{\ast}$, we obtain
-(3.42) $\begin{split}&\Big{\|}v_{t}^{\varepsilon,2}-\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu})(R_{t}^{\varepsilon}-1)]\Big{\|}_{\mathbb{B}}\leq\frac{1}{\varepsilon}\mathbb{E}[\alpha(|X_{t}^{\mu}-X_{t}^{\mu_{\varepsilon},\mu}|)|R_{t}^{\varepsilon}-1|]\\ &\leq\frac{1}{\varepsilon}\Big{(}\mathbb{E}[|R_{t}^{\varepsilon}-1|^{k^{*}}]\Big{)}^{\frac{1}{k^{*}}}\alpha\Big{(}\big{(}\mathbb{E}[|X_{t}^{\mu}-X_{t}^{\mu_{\varepsilon},\mu}|^{k}]\big{)}^{\frac{1}{k}}\Big{)}.\end{split}$
-By (3.24) and (3.36), we find a constant $c_{0}(\mu,\phi)>0$ such that
-(3.43) $\mathbb{E}[\|V(X_{t}^{\mu_{\varepsilon},\mu})\|_{\mathbb{B}}^{k}]\leq c_{0}(\mu,\phi),\ \ t\in[0,T],\varepsilon\in(0,1].$
-On the other hand, by (3.40), for any $p\geq 1$ there exists a constant $c(p,\mu,\phi)>0$ such that
-(3.44) $\mathbb{E}[|R_{t}^{\varepsilon}-1|^{p}]\leq c(p,\mu,\phi)\varepsilon^{p}\bigg{(}\int_{0}^{t}\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\big{)}\text{\rm{d}}s\bigg{)}^{\frac{p}{2}},\ \ t\in[0,T],\varepsilon\in[0,1].$
-Combining this with (3.41) and (3.43), we find a constant $c_{1}(\mu,\phi)>0$ such that
-$\|v_{t}^{\varepsilon,2}\|_{\mathbb{B}}$ $\leq\big{(}\mathbb{E}[\|V(X_{t}^{\mu_{\varepsilon},\mu})\|_{\mathbb{B}}^{k}]\big{)}^{\frac{1}{k}}\frac{\big{(}\mathbb{E}[|R_{t}^{\varepsilon}-1|^{k^{*}}]\big{)}^{\frac{1}{k^{*}}}}{\varepsilon}$
-$\leq c_{1}(\mu,\phi)\bigg{(}\int_{0}^{t}\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\big{)}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}.$
-This together with (3.33) and (3.37) yields that for some constant $c_{2}(\mu,\phi)>0$,
-$\|v_{t}^{\varepsilon}\|_{\mathbb{B}}^{2}\leq 2\|v_{t}^{\varepsilon,1}\|_{\mathbb{B}}^{2}+2\|v_{t}^{\varepsilon,2}\|_{\mathbb{B}}^{2}\leq c_{2}(\mu,\phi)\bigg{(}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}+\int_{0}^{t}\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)},\ \ t\in[0,T].$
-By Grownwall’s inequality, we find a constant $c_{3}(\mu,\phi)>0$ such that
-(3.45) $\begin{split}&\|v_{t}^{\varepsilon}\|_{\mathbb{B}}^{2}\leq c_{2}(\mu,\phi)\bigg{(}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}+\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\text{\rm{e}}^{c_{2}(\mu,\phi)(t-s)}\text{\rm{d}}s\bigg{)}\\ &\leq c_{3}(\mu,\phi)\frac{\alpha(t^{\frac{1}{2}})^{2}}{t},\ \ t\in[0,T],\ \ \varepsilon\in(0,1].\end{split}$
-(c) For $|\tilde{v}_{s}^{\varepsilon}|$.
-Recall that $\Theta_{t}=u_{t}+id$ for $u_{t}=u_{t}^{\lambda_{0},\mu}$.
-Similarly to (2.16) and (2.21), by Itô’s formula for $\Theta_{t}(X_{t}^{\mu_{\varepsilon}})-\Theta_{t}(X_{t}^{\mu})$, we obtain
-(3.46) $\begin{split}&\text{\rm{d}}\tilde{v}_{t}^{\varepsilon}=\frac{1}{\varepsilon}\Big{\{}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X_{t}^{\mu_{\varepsilon}})-\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X_{t}^{\mu})\Big{\}}\text{\rm{d}}W_{t}\\ &+\frac{1}{\varepsilon}\Big{[}\lambda_{0}u_{t}(X_{t}^{\mu_{\varepsilon}})-\lambda_{0}u_{t}(X_{0}^{\mu})+\nabla_{B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu_{\varepsilon},P_{t}V(\mu_{\varepsilon}))-B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu,P_{t}V(\mu))}\Theta_{t}(X_{t}^{\mu_{\varepsilon}})\Big{]}\text{\rm{d}}t\\ &+\frac{1}{\varepsilon}[B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu,P_{t}V(\mu))-B_{t}(X_{t}^{\mu},P_{t}^{*}\mu,P_{t}V(\mu))]\text{\rm{d}}t,\\ &\tilde{v}_{0}^{\varepsilon}=\frac{\Theta_{0}(X_{0}^{\mu_{\varepsilon}})-\Theta_{0}(X_{0}^{\mu})}{\varepsilon}.\end{split}$
-Similarly to (3.30), by combining this with (B), (2.9), (3.45) and the stochastic Gronwall inequality, for any $j\geq 1$ we find constants $c(j),c_{4}(\mu,\phi)>0$ such that
-(3.47) $\begin{split}&\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c(j)|\phi(X_{0}^{\mu})|^{j}+c_{4}(\mu,\phi)\\ &\qquad+c(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}},\ \ t\in[0,T],\varepsilon\in(0,1].\end{split}$
-Letting $j=k$, taking expectation to both sides and noting that
-$c(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{\frac{k+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{k}{k+1}}$
-$\leq\tilde{c}(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\leq\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}+\frac{\tilde{c}(j)^{2}}{2}\int_{0}^{t}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}\text{\rm{d}}s,$
-for some $\tilde{c}(j)\geq c(j)$ as in (3.31), by Gronwall’s inequality we find a constant $c_{5}(\mu,\phi)>0$ such that
-(3.48) $\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}\leq c_{5}(\mu,\phi),\ \ \varepsilon\in(0,1],$
-so that (3.47) implies
-(3.49) $\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\tilde{v}_{s}^{\varepsilon}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{6}(\mu,\phi)+c(j)|\phi(X_{0}^{\mu})|^{j},\ \ \varepsilon\in(0,1]$
-for some constant $c_{6}(\mu,\phi)>0.$
-(d) For $|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}$. Now, let
-$\psi_{t}^{\varepsilon}:=\,$ $\mathbb{E}\Big{[}\big{\langle}\{D^{L}B_{t}(z,\cdot,P_{t}V(\mu))\}(P_{t}^{*}\mu)(X_{t}^{\mu}),(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}^{\varepsilon}\big{\rangle}\Big{]}|_{z=X_{t}^{\mu}}$
-+$\qquad+\big{\{}\nabla^{\mathbb{B}}_{v_{t}^{\varepsilon}}B_{t}(X_{t}^{\mu},P_{t}^{*}\mu,\cdot)\big{\}}(P_{t}V(\mu)).$
-Then (3.46) can be reformulated as
-$\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{0}^{\varepsilon}$
-$=\int_{0}^{t}\Big{\{}\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\big{[}B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}\big{]}(X_{s}^{\mu})+\nabla_{\psi_{s}^{\varepsilon}}\Theta_{s}(X_{s}^{\mu})\Big{\}}\text{\rm{d}}s$
-$\quad+\int_{0}^{t}\big{[}\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu})\big{]}\text{\rm{d}}W_{s}+\alpha_{t}^{\varepsilon},\ \ t\in[0,T],$
-where for $t\in[0,T],$
-$\alpha_{t}^{\varepsilon}:=\int_{0}^{t}\xi_{s}^{\varepsilon}\text{\rm{d}}s+\int_{0}^{t}\tilde{\xi}_{s}^{\varepsilon}\,\text{\rm{d}}W_{s},$
-$\xi_{s}^{\varepsilon}:=\frac{1}{\varepsilon}\{[B_{s}(\cdot,P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))+\lambda_{0}u_{s}](X_{s}^{\mu_{\varepsilon}})-[B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}](X_{s}^{\mu})\}$
-$\qquad\qquad+\nabla_{\frac{1}{\varepsilon}\{B_{s}(X_{s}^{\mu_{\varepsilon}},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon}},P_{s}^{*}\mu,P_{s}V(\mu))\}}u_{s}(X_{s}^{\mu_{\varepsilon}})$
-$\qquad-\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\big{[}B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}\big{]}(X_{s}^{\mu})-\nabla_{\psi_{s}^{\varepsilon}}\Theta_{s}(X_{s}^{\mu}),$
-$\tilde{\xi}_{s}^{\varepsilon}:=\frac{\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu_{\varepsilon}})-\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu})}{\varepsilon}-\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu}).$
-By the same argument leading to [29, (4.23),(4.26)] and (3.30) above, (B) together with (2.9), (3.45), (3.48), (3.49) implies
-$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\alpha_{s}(\varepsilon)|^{j}\Big{|}\mathscr{F}_{0}\Big{]}=0,\ \ j\geq 1,$
-and by Jensen’s inequality we find a constant $c>0$ such that
-$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\leq c|\tilde{v}_{0}^{\varepsilon}-\tilde{v}_{0}|^{k}+c\bigg{(}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\alpha_{s}(\varepsilon)|^{2k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{)}^{\frac{1}{2}}$
-(3.50) $\quad+c\bigg{(}\int_{0}^{t}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}+c\bigg{(}\int_{0}^{t}\Big{\{}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\Big{\}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in[0,T].$
-(e) For $\|v_{t}^{\varepsilon}-\hat{v}_{t}\|_{\mathbb{B}}$. Similarly to (d), for $N_{s}$ and $\tilde{N}_{s}$ in (3.13), we have
-(3.51) $\begin{split}&\frac{R_{t}^{\varepsilon}-1}{\varepsilon}=\int_{0}^{t}R_{s}^{\varepsilon}\big{\langle}\varepsilon^{-1}\eta_{s}^{\varepsilon},\text{\rm{d}}W_{s}\big{\rangle}\\ &=\int_{0}^{t}R_{s}^{\varepsilon}\Big{\langle}\frac{\zeta_{s}(X_{s}^{\mu_{\varepsilon},\mu})[B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu,P_{s}V(\mu))]}{\varepsilon},\ \text{\rm{d}}W_{s}\Big{\rangle}\\ &=h_{t}(\varepsilon)+\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(v^{\varepsilon})+\tilde{N}_{s}(\tilde{v}^{\varepsilon})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle},\end{split}$
-where
-$h_{t}(\varepsilon):=\int_{0}^{t}\Big{\langle}$ $\zeta_{s}(X_{s}^{\mu})\Big{\{}R_{s}^{\varepsilon}\frac{B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu,P_{s}V(\mu))}{\varepsilon}$
-$\qquad\qquad-N_{s}(v^{\varepsilon})-\tilde{N}_{s}(\tilde{v}^{\varepsilon})\Big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}$
-satisfies
-(3.52) $\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|h_{t}(\varepsilon)|^{p}\Big{|}\mathscr{F}_{0}\Big{]}=0,\ \ p\geq 1.$
-Noting that $X_{t}^{\mu}=X_{t}^{\mu,\mu}$, by (3.25) we find a constant $c(\mu,\phi)>0$ such that
-(3.53) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\mu_{\varepsilon},\mu}-X_{t}^{\mu}|^{k}\Big{]}\leq c(\mu,\phi)\varepsilon^{k},\ \ \varepsilon\in[0,1],\mu\in\mathscr{P}_{k}.$
-Combining these with (3.42), (3.45) and (3.44), we find a constant $k_{1}(\mu,\phi)>0$ such that
-(3.54) $\Big{\|}v_{t}^{\varepsilon,2}-\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu})(R_{t}^{\varepsilon}-1)]\Big{\|}_{\mathbb{B}}\leq k_{1}(\mu,\phi){{\alpha(\varepsilon)}},\ \ t\in[0,T],\ \varepsilon\in(0,1).$
-By (3.14) for $v_{t}=v_{t}(\mu,\phi),\tilde{v}_{t}=\tilde{v}_{t}(\mu,\phi)$, and by the definition of $\hat{v}_{t}$ in (3.35), we have
-$\hat{v}_{t}$ $=I_{t}^{V}(\mu,\phi)+\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(\hat{v})+\tilde{N}_{s}(\tilde{v})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}.$
-Combining this with (3.33), (3.38), (3.52), (3.51), (3.54) and the same argument to (3.27), we find a constant $k_{2}(\mu,\phi)$ and a measurable function $\tilde{h}:(0,T]\times[0,1]\rightarrow(0,\infty)$ with
-(3.55) $\sup_{\varepsilon\in(0,1],t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)\leq k_{2}(\mu,\phi),\ \ \lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)=0,$
+$\lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\int_{\mathbb{R}^{d}}\bigg{\|}\mathbb{E}\bigg{[}U_{t}(\cdot,\cdot,X_{t}^{x,\mu})$
+(3.49) ×$\qquad\qquad\quad\times\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})\big{\{}[v^{\varepsilon,1}_{s}-\hat{v}_{s}^{1}](X_{s}^{x,\mu},P_{s}^{\ast}\mu)\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\bigg{\|}_{\infty}\mu(\text{\rm{d}}x)=0.$
+Moreover, combining (3.3) with $[U_{t}(y,\nu,\cdot)]_{\alpha}\leq 1,$ and (2.9) for $p=k^{\ast}$, we obtain
+$\Big{\|}v_{t}^{\varepsilon,2}-\int_{\mathbb{R}^{d}}\frac{1}{\varepsilon}\mathbb{E}[U_{t}(\cdot,\cdot,X_{t}^{x,\mu})(R_{t}^{\varepsilon,x}-1)]\mu(\text{\rm{d}}x)\Big{\|}_{\infty}$
+$\leq\frac{1}{\varepsilon}\int_{\mathbb{R}^{d}}\mathbb{E}[\alpha(|X_{t}^{x,\mu}-X_{t}^{x+\varepsilon\phi(x),\mu}|)|R_{t}^{\varepsilon,x}-1|]\mu(\text{\rm{d}}x)$
+$\leq\frac{1}{\varepsilon}\int_{\mathbb{R}^{d}}\left\{\Big{(}\mathbb{E}[|R_{t}^{\varepsilon,x}-1|^{k^{*}}]\Big{)}^{\frac{1}{k^{*}}}\alpha\Big{(}\big{(}\mathbb{E}[|X_{t}^{x,\mu}-X_{t}^{x+\varepsilon\phi(x),\mu}|^{k}]\big{)}^{\frac{1}{k}}\Big{)}\right\}\mu(\text{\rm{d}}x).$
+This together with (3.45) and (3.39) yields that for some constant $k_{1}(\mu,\phi)>0$,
+$\Big{\|}v_{t}^{\varepsilon,2}-\int_{\mathbb{R}^{d}}\frac{1}{\varepsilon}\mathbb{E}[U_{t}(\cdot,\cdot,X_{t}^{x,\mu})(R_{t}^{\varepsilon,x}-1)]\mu(\text{\rm{d}}x)\Big{\|}_{\infty}\leq k_{1}(\mu,\phi){{\alpha(\varepsilon)}},\ \ t\in[0,T],\ \varepsilon\in(0,1].$
+Combining this with (3.30), (3.3), (3.43), (3.3), (3.42), (3.48), and the argument leading to (3.22), we find a constant $k_{2}(\mu,\phi)$ and a measurable function $\tilde{h}:(0,T]\times(0,1]\rightarrow(0,\infty)$ with
+(3.50) $\sup_{\varepsilon\in(0,1],t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)\leq k_{2}(\mu,\phi),\ \ \lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)=0$
 such that
-$\|v_{t}^{\varepsilon}-\hat{v}_{t}\|_{\mathbb{B}}$
-(3.56) $\leq\tilde{h}_{t}(\varepsilon)+\bigg{\|}\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(v^{\varepsilon}-\hat{v})+\tilde{N}_{s}(\tilde{v}^{\varepsilon}-\tilde{v})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\bigg{\|}_{\mathbb{B}}$
-$\leq\tilde{h}_{t}(\varepsilon)+k_{2}(\mu,\phi)\bigg{(}\int_{0}^{t}\Big{\{}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}+\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\big{)}^{\frac{2}{k}}\Big{\}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in(0,T].$
-(f) Proof of (3.35). By (3.49)-(3.3) and Fatou’s lemma, we see that
-$\begin{split}&\beta_{t}:=\limsup_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{]},\ \ \tilde{\beta}_{t}:=\limsup_{\varepsilon\rightarrow 0}\sup_{s\in(0,t]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}\end{split}$
-satisfy
-(3.57) $\begin{split}&\beta_{t}\leq\mathbb{E}\bigg{[}\limsup_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{]}\\ &\leq c\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\limsup_{\varepsilon\rightarrow 0}\frac{s}{\alpha(s^{\frac{1}{2}})^{2}}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}\\ &\qquad\quad+c\bigg{(}\int_{0}^{t}\Big{\{}\limsup_{\varepsilon\rightarrow 0}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\Big{\}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\\ &\leq c\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\,\tilde{\beta}_{s}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}+c\left(\int_{0}^{t}\beta_{s}^{2}\text{\rm{d}}s\right)^{\frac{1}{2}},\ \ t\in[0,T].\end{split}$
-Noting that (3.16), (3.34) and (3.45) imply that $\tilde{\beta}_{t}$ satisfies
-$\sup_{t\in(0,T]}\tilde{\beta}_{t}\leq\sup_{\varepsilon\in(0,1]}\sup_{s\in(0,T]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}=:\tilde{c}(\mu,\phi)<\infty,$
+$\left\|v_{t}^{\varepsilon,2}-\hat{v}_{t}^{2}\right\|_{\infty}$ $\leq\tilde{h}_{t}(\varepsilon)+\int_{\mathbb{R}^{d}}\bigg{\|}\mathbb{E}\bigg{[}U_{t}(\cdot,\cdot,X_{t}^{x,\mu})$
+(3.51) ×$\qquad\qquad\quad\times\int_{0}^{t}\Big{\langle}\{\zeta_{s}[v^{\varepsilon,2}_{s}-\hat{v}_{s}^{2}]\}(X_{s}^{x,\mu},P_{s}^{\ast}\mu),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\bigg{\|}_{\infty}\mu(\text{\rm{d}}x)$
+$\leq\tilde{h}_{t}(\varepsilon)+k_{2}(\mu,\phi)\bigg{(}\int_{0}^{t}\|v_{s}^{\varepsilon,2}-\hat{v}_{s}^{2}\|_{\infty}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in(0,T].$
+(d) Proof of (3.31). Let
+$\begin{split}\beta_{t}:=\limsup_{\varepsilon\rightarrow 0}\sup_{s\in(0,t]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon,2}-\hat{v}_{s}^{2}\|_{\infty}.\end{split}$
+Noting that (3.18), (3.26), (3.27), (3.21) and (3.40) imply that $\beta_{t}$ satisfies
+$\sup_{t\in(0,T]}\beta_{t}\leq\sup_{\varepsilon\in(0,1]}\sup_{s\in(0,T]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon,2}-\hat{v}_{s}^{2}\|_{\infty}=:\tilde{c}(\mu,\phi)<\infty,$
 so that by Fatou’s lemma in (3.3) we derive from (3.50) that
-$\tilde{\beta}_{t}^{2}\leq Ck_{2}(\mu,\phi)^{2}\int_{0}^{t}\Big{\{}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\tilde{\beta}_{s}^{2}+\beta_{s}^{\frac{2}{k}}\Big{\}}\text{\rm{d}}s,\ \ t\in(0,T],$
+$\beta_{t}^{2}\leq Ck_{2}(\mu,\phi)^{2}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\beta_{s}^{2}\text{\rm{d}}s,\ \ t\in(0,T],$
 where by (2.2),
 $C:=\sup_{t\in(0,T]}\frac{t}{\alpha(t^{\frac{1}{2}})^{2}}<\infty.$
-Combining this with (3.57), $\int_{0}^{T}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}\text{\rm{d}}t<\infty$, and applying Gronwall’s inequality, we prove (3.35) and hence finish the proof. ∎
+Combining this with $\int_{0}^{T}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}\text{\rm{d}}t<\infty$, and applying Gronwall’s inequality, we prove (3.31), which together with (3.34) completes the proof. ∎
 We are now ready to prove Theorem 3.3.
 Proof of Theorem 3.3.
-By (3.10) and (3.12), it suffices to prove that for any $t\in(0,T]$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$,
-(3.58) $\begin{split}\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}=\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}.\end{split}$
-Let $R_{t}^{\varepsilon}$ be in (3.39). By (3.41) for $f$ replacing $V$, we obtain
-(3.59) $\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}=\frac{1}{\varepsilon}\mathbb{E}\big{[}f(X_{t}^{\mu_{\varepsilon},\mu})(R_{t}^{\varepsilon}-1)\big{]},\ \ t\in(0,T].$
+By (3.12) and (3.13), it suffices to prove that for any $t\in(0,T]$ and $f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d})$,
+(3.52) $\begin{split}&\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}\\ &=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})N_{s}(\mu,\phi),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).\end{split}$
+Let $R_{t}^{\varepsilon,x}$ be in (3.35). By (3.3) for $f$ replacing $U_{t}(y,\nu,\cdot)$, we obtain
+(3.53) $\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}=\int_{\mathbb{R}^{d}}\frac{1}{\varepsilon}\mathbb{E}\big{[}f(X_{t}^{x+\varepsilon\phi(x),\mu})(R_{t}^{\varepsilon,x}-1)\big{]}\mu(\text{\rm{d}}x),\ \ t\in(0,T].$
 Noting that (3.45) implies
-$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\mu_{\varepsilon},\mu}-X_{t}^{\mu}|^{k}\Big{]}=0,$
-while $(B_{2})$, Lemma 3.5 and (3.39) lead to
-$\lim_{\varepsilon\rightarrow 0}\frac{R_{t}^{\varepsilon}-1}{\varepsilon}=\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}$
-in $L^{2}(\mathbb{P})$, by taking $\varepsilon\rightarrow 0$ in (3.59) we deduce (3.58) for $f\in C_{b}(\mathbb{R}^{d})$.
-By an approximation argument as in [29, Proof of (2.3)] for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, this implies (3.59) for $f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
+$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{x+\varepsilon\phi(x),\mu}-X_{t}^{x,\mu}|^{k}\Big{]}=0,$
+while (3.43), (3.42), Lemma 3.5, (3.29), (3.26), (3.17) and (3.31) lead to
+$\lim_{\varepsilon\rightarrow 0}\frac{R_{t}^{\varepsilon,x}-1}{\varepsilon}=\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})N_{s}(\mu,\phi),\ \text{\rm{d}}W_{s}\Big{\rangle}$
+in $L^{2}(\mathbb{P})$, by taking $\varepsilon\rightarrow 0$ in (3.53) and using the dominated convergence theorem, we deduce (3.52) for $f\in C_{b}(\mathbb{R}^{d})$.
+By an approximation argument as in [30, Proof of (2.3)] for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, this implies (3.52) for $f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$ By the approximation argument used in the proof of (3.4), we may further extend (3.52) to $f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}).$